REDUCE Bibliography

Anthony C. Hearn The RAND Corporation

Santa Monica, CA 90406-2138 USA

15 February 2012 Version 1.16

This document contains a list of all known references to REDUCE. It no doubt contains errors and omissions. Please report these by regular mail (preferably in BibTEX format) to the REDUCE Secretary, The RAND Corporation, P.O. Box 2138, Santa Monica CA 90406-2318, or by electronic mail to reduce@rand.org. An electronic copy of the bibliography in BibTEX format is also available from the latter address.

RAND Publication CP162 (Rev. 9/90)

Copyright \begin{matrix}c \\ ◯ \end{matrix}1990 The RAND Corporation. All rights reserved.

Title word cross-reference

#11 [WR79]. #2 [BH72, Hea72c]. #3 [Fit73]. #4 [Fit73]. #8 [Har78].

x [Köl83b]. x dx [Köl83b]. 23 ≤ n ≤ 26 [Ng89]. 3 [Eas91]. α [Adk83]. α^{2} [AB70]. BŜO_{n} < 16 > [Ng89]. CD_{N} [NMJ84a]. e^{+}e^{−} [FGK89]. \frac{1}{2} [Garxx, Gar80]. γ [DT69, WTVNF87]. γ^{5} [IK87]. ggqq̄ll̄ [GK85]. gg → c\overline{c}g [KS88]. H + F_{2} [CCEL84]. h̄^{2} [HG89]. ∫_{0}^{1} t^{−1} [Köl82]. ∫_{0}^{1} x^{ν−1}(1 − x)^{−λ} [Köl85b]. ∫_{0}^{∞} e^{−μt}t^{ν−1} [Köl83a]. ∫_{0}^{∞} x^{ν−1}(1 + βx)^{−λ} [Köl86]. ∫_{0}^{π/2} [Köl83b]. λ + N [Par71]. m = 0 [Köl81]. m = 1 [Köl81]. O + B [Par71]. \overline{ν}_{e} + μ + ν_{μ} [TH65]. π [LP79]. π_{0}π [Sch80]. P_{−1/2+iτ}^{m} (x) [Köl81]. p_{τ} [ADB^{+}84b]. → [GG78, Per78, TH65]. r^{t} [BBR85]. S [Par71]. U(5) ⊃ O(5) ⊃ O(3) [YP88, YP89]. X_{0}^{7} + X_{1}^{7} + X_{2}^{7} = 0 [Kob84]. mtdt [Köl83a]. ^{m}xdx [Köl85b, Köl86]. ^{n} [Köl83b]. ^{n−1t} [Köl82]. ^{p} [Köl83b]. ^{p}(1 − t)dt [Köl82]. _{2D}−_{13C} [NMJ84a].

-Channel [Par71]. -Clebsch [GKR85]. -correction [HG89]. -D [Eas91]. -matrix [IK87]. -Quadrupole [GKRR85a]. -ray [WTVNF87].

0 [Art95, Eas93]. 0-19-853443-4 [Eas93]. 0-7503-0277-1 [Art95].

1 [Art95]. 1-11-41 [NM89a]. 1. [Gro88a]. 1700 [GK78]. 1D [DD85]. 1SO [FGM73]. 1SO- [FGM73].

2 [Ozixx, Rod84]. 2. [Gro88a]. 286-based [YA87]. 2n [BH72, Hea72c].

3 [DMR88]. 3-Spaces [PLSV84]. 3. [Gro88b]. 3D [EH86]. 3D-Operational [EH86]. 3PO [FGM73]. 3PO- [FGM73].

4 [Eas93, IK87]. 4-dimensional [IK87]. 4. [Gro88b]. 4th [Fox71].

5. [Gro88b]. 5th [ESP88].

6. [Gro88c].

7. [Gro88c]. 74 [Hus81].

8. [Gro88c]. 86 [YA87]. 86/286 [YA87]. 86/286-based [YA87]. ’88 [ESP88]. 8th [DL86].

Åman [CFÅ84]. Abbott [ABD85, ABD86a, AD87, Abb87, Abb88, AD88, Abb89, ABD89]. Abdali [AW88]. Abe [UYSA89]. Abelian [KRxx, KR88]. Abelson [CAH90]. Aberration [GS77, Som77]. Abi-Ezzi [Abi83]. Abschätzung [Kra82]. Absolute [Cap86a, PN73b, Wan85]. Absorption [DL88a, WTVNF87]. Abstrait [Lux75]. Academic [DST88, Fit89a, GvH89, MN89b, NM89b, MN89c, RA89]. Academy [GKRR85a, GKR85, GKRR85b, Per84a, Per86a]. Accelerator [AB89, KKM81]. Accuracy [JW86, Lis84]. ACM [Gri76a, Gri77a, HN79, MHGG79, Abb89, Dew89, Fit89b, JW86, vHHGV89]. ACRITH [JW86]. Acrivos [BBA73]. Acta [Cam74, SW83, SK83]. Acting [JO84]. ACU [Sch80]. Adamchik [AM90]. Adams [Ada83]. Adaptive [DD85]. Adding [vdH86a]. Addison-Wesley [AM90, Bra90, Fit90, GS90, GMS90, Gat90b, GZ90b, GKZ90, LD90, Nor90, PB90, Tao90]. Addition [Gro88b]. Adiabatic [Bir77]. Adic [Abb89]. Adkins [Adk83, AB83, Adk85]. Administration [NS82]. Admissible [Kre88]. Adv [Ken88]. Advanced [GS78]. Aeronautics [Kam69a, Kam69b]. Aerospace [Squxx]. Aggregates [Say87a]. Aguilera-Navarro [ANGK^{+}87]. Ahmed [PA86]. AI [CAH90]. aided [SM72, DJ89]. Akademie [DJ80]. Akasaka [NM89a]. al [MMH88]. Alan [PB90]. Alberta [KKM81]. Albrecht [CvH83, vHC83]. Alderton [PSA84]. Aldins [ABDK69]. Aleksandr [EKR89]. Alekseev [AE86, AE87b, AE87a]. Alexey [GKZ90]. Alexis [Vos89]. Alfeld [Alf82]. Alfven [DL88a]. Algebra [BC90, BDM^{+}87, BCDS87, BJ87, CvH83, Cap90, Coh89, Dav88, DST88, Vos89, Eis90, EM82, EC87, FGGS87, FK87, Fit85a, Fit89a, FBC86, GL89, GSZ85, GKRR85a, GKR85, GKRR85b, GSSZ87, GKK87, GK89, GL83, GvH89, Gri75, GK78, HWH89, Hap89, Hea73, Hea80, Hea82, Kre88, Lev71, LOG85, Mac89a, McC87, MN89b, NM89b, ON90, Ogi82, Ogi89, Ray87, Sch87, Sch88b, SGM87, Sto77b, vH81, vH82, vHC83, vH89, WH89]. Algebraic [ABD86a, Abb88, Abb89, BGK86, Bro71, BH79, Cam67, CH85, CLS76, FC79, CFÅ84, Dau79, DJ80, DJRR81, Dav81, DST88, DDDD85, Dic85, ER90, Fit85b, GKZ90, GKK87, Gri74, Gri76a, Hea71b, Hea74a, Hea76a, Hea77, Hea78, Hea85, Hea86, HM75, Kan75, KMH88, KR87, Mac86a, McI85, NS82, NS85, PH81, PH83, RT87a, Sag88, Stu88, SGT90, Tao90, Tou79, WTV87, YP88, YP89, Eas93]. algébrica [RdS88]. Algébrico [dS88]. Algébriques [Duv87]. Algorithm [AD87, AM90, Bra90, FGPF88, FGP89, GM85, GM88, Gri78, Gri76c, Hea72b, IKRT89, MN89b, MN89c, Nor90, NM77, Sch85b, vH81]. Algorithmen [BC90]. Algorithms [DST88, HL73, Smi79, WD84]. Alignment [JPS87]. Alloys [Say87a]. Alternative [GP82]. Am [CBR84, FP85]. Ames [RA89]. Amirkhanov [AZZ87]. Amplitude [BHP62, SH81]. Amplitudes [AZZ87, GG78, LP79, LSS87, Par71, Sch80]. Amsterdam [BHM86, ESP88, GGRxx]. Analitik [Hus81]. Analitik-74 [Hus81]. Analysis [CH70, DR89, DG75, DT69, Eas91, Eis90, Eit73, Fle71, GL89, GP82, GM84, Lis84, LSS86, Per78, QGxx, Ren92, RT85, SM83, Sch77, Sch83b, SM72, SL74, Sto77a, TC88, vHH82, WCvH84, Wat85]. Analytic [Ada83, Adk83, BBA73, CNxx, EKR89, FH74, GK85, HT87, HW81, Loo72]. Analytical [FP85, Ger80a, GTS80, GKRR85b, LV84, LS79, LM85, MHK86, Nik87, Pic69, Rod84, CGGK86]. Analytically [Rin71, Sto77b]. Analyzing [GMS90, Hea74a, Tro89]. Anatolii [EKR89]. Andersen [NA79]. Andreas [ASW89]. Ang [JO84]. Angeles [BH84b]. Angewandte [LM73]. Angle [GBB72, GBB73, NMJ84b]. Angular [Gro88b]. Anharmonic [HT87]. Anisotropy [Say87a, Say87b]. Ann [Car73, Vor77]. Anna [HH80]. Annihilating [Ker86]. annihilation [FGK89, Kau73]. Annual [ESP88]. Anomalies [ZY75]. Anomalous [BS67, Cal72a, Car73, Par68, Van76]. Anomaly [LS80]. Anteile [Kra82]. Anthony [BH72, DR86b, FH74, GH79, Hea71a, Hea71b, Hea72a, Hea72b, Hea72c, HL73, Hea73, Hea74b, Hea74a, Hea76c, Hea78, Hea79, HN79, Hea80, HW81, Hea81, Hea82, HGB82, Hea85, Hea86, Hea87, MH85]. Antitranslator [Kry84]. Antweiler [ASW89]. Anwendung [Mac73]. any [Wan85]. ao [dS88]. Aoki [YA87].

Aperture [Ada83]. appear

[Cap89, DCM^{+}89]. Appelquist [AB70]. Appl [Kee83, Ken88, Köl86]. Applicability [ZNR78]. Application [Bir77, BC81, CNxx, CC88, CCEL84, Dau79, DG75, Ede81, Eis90, EM82, EC87, FK87, GMS90, GL89, GSZ85, GSSZ87, GL83, Hea81, Kam69b, Kat85, LM74, MS85, MIG85, Ozixx, PH81, Pic69, PR84, Rin71, Sch84, SGM87, UYSA89, WH89]. Applications [BDM^{+}87, BH79, CvH83, DR86b, GTS80, GM84, Hea76c, Lev71, LExx, Ogi82, Som85, Spi87, Squxx]. Applied [CFW87, EH87, FBC86, HvH77, HvHS81, Mac89b, SHM87, vH81]. Applying [Fit85a]. Approach [Cal72b, vH83]. Approaches [BHM86, Kad88, Per88]. Approximants [FGM73, NZ82, NZ87]. Approximation [CC88, CCF84, CCEL84, GS90, HvH77, PA86, Sch80]. Approximations [Abb89, GVxx, GV78]. Arbitrary [NM88, NM89b, Ono79, PSA84, Sas79]. Arbor [Car73]. Arbuzov [ABD86b]. Arch [GV86]. Arctangent [MM89]. Arguments [RT87b]. arise [GKZ90]. Arising [Köl84]. Arithmetic [GvH89, JW86, NM88, NM89b, Ono79, Sas79]. Arnaldo [GV86]. Arter [Art95]. Arthur [HN79, Nor78, NW83]. Articulated [Kah69]. Aso [ANK81]. Aspects [Gla82, RT87a, vH87]. assistente [RdS88]. Associated [YP88, YP89]. Ast [Say87b]. Astrogeophyics [DJ89]. Astron [DJRR81, KK85]. Astronautics [Kam69a, Kam69b]. Astronomy [DLS88]. Astrophysical [Cam68]. Astrophysics [Dau79, NS86]. Astrophysik [DJ80]. Asymptotic [AH73, CCF84, CCEL84, KS84a, NK84, Vor77]. Asymptotics [ABD86b]. ATENSOR [IK96]. Atha [Garxx]. Atherton [AH73]. Atmos [Boy78]. Atmospheric [ANK81, FBC86]. Atom [VR87, WTVNF87]. Atomic [Bro72, FC84]. Attilio [BC81]. Augustin [Lux75].

Aurenche [ADB^{+}84a, ADB^{+}84b]. Aust [Cam68]. Austral [Kea85]. Austria [WKL88]. Austriaca [Cam74]. Autin [AB89]. Automatic [Eit73, GvH85, Gat85a, GvH89, HvH84, Sto77a, vdHvHG87]. Automatically [Sch83a, Sch85a]. automation [LD90]. Autonomous [Sch85b, Sch86]. Availability [Wat85]. avalanches [Sav90]. Axial [AE86, AE87b, AE87a, LM85, LH70, LPSV82, Per86b]. Axially [BH84a]. Axis [QGxx]. Axisymmetric [BS86, Kad88, KPT71, MMH88]. Axisymmetry [Per88]. Ayse [Bil92].

B1700 [GK79]. B1700/B1800 [GK79]. B1800 [GK79]. Bäcklund [Ker86]. Baecklund [KF87]. Baekler [BH84a, BH84b, BHM86, BHHS87, BG87, BMH87, BSW88, BGHM88, BGH88, MBG87]. Bahrdt [BW90]. Baier [ADB^{+}84a, ADB^{+}84b]. Baikov [BS98]. Bail-lin [zZlH84]. Bajla [BOH78]. Baker [BBF^{+}82]. Balian [BPV78]. Bamberger [Bam88]. Band [SH72]. Band-Pass [SH72]. Bandwidth [Gre84]. Barbara [GvH85, Gat85a, Gat85c, Gat85b, Gat86]. Barber [CBR84]. Bark [BT78, CB76]. Barnes [PB90]. Barthes-Biesel [BBA73]. Barton [BH72]. based [Boc92, Fit81, HvH85, Ina80, Kam69b, Nor90, Ren92, YA87, KS81a]. Bases [BGK86, MMN88, vdH86b]. Basic [GKRR85a]. Basios [BCM^{+}95]. Bateman [BS86]. Bath [ABD89, BDS88, ABD86a]. Bau [FB89]. Bauer [BNdP^{+}87]. be [CDS88, Fit89b]. Beam [GS78]. Beams [LExx, Pes73]. Bearbeitung [LM73]. Bed [SM83]. Beginners [WH89]. Beijing [zZlH84]. Benchmark [MH85]. Beng [Ng89]. Bengt [Mal82]. Bengtsson [AB89]. Bennett [BDS88]. Benofy [BBF^{+}82]. Benson [HGB82]. Berends [BKdCW81]. Bericht [BW90]. Berkovich [BGKN89].

Berlin [Gat90a, MMN88, MN89a, Möl89, NM88, WD90]. Berman [BT63]. Bernd [EH86]. Bessel [Pie84, PA86]. Bessis [BBR85]. Bethe [FT75, Fle71]. Bethe-Salpeter [FT75, Fle71]. between [Dav82a, Gre84, HH80, HGDR84, PSA84]. Beyond [ADB^{+}84b]. BFORT [Ono79]. Bhaile [Garxx]. Bifurcation [FGP89, LP90]. big [KS81a]. big-float [KS81a]. Bilge [Bil92]. bilinear [Hie87b, Hie87c, Hie87d, Hie88, Ito88]. Billoire [BLMN78]. Bing-Lin [ZY75]. Biochemistry [BDS88]. Biology [Kee85]. BioPhys [Hor83]. Biopolymers [KM80]. Biro [BZZ86, BNdP^{+}87]. Birrell [Bir77]. Biswas [BCTC75]. Bittencourt [BC90]. Blankenbecler [GBB72, GBB73]. Bledsoe [GLB75]. Bocko [Boc92]. Boege [BGK86]. Bogdanova [BH88]. Book [Art95, Eas93]. Boole [DT84]. Boos [ABD86b]. Bordoni [BC81]. Borel [LS80]. Boson [ANGK^{+}87, BT63, BCTC75, BS67, FT75, GKRR85a, GKR85, GKRR85b, Per78]. Bosons [Cam68]. Bottom [GvH89]. Bottom-Up [GvH89]. Bound [WTVNF87]. Boundary [BS87, MMR87, MS85, RA89, Kea85]. Boundary-Value [Kea85]. Bounding [AZZ87]. Bowyer [BDM^{+}87]. Boyd [Boy78, FBC86]. Brackx [BS87, BCDS87, BCRS89]. Bradford [ABD85, ABD86a, ABD89, Bra90, BHPS86, BD88]. Branch [Wat79]. Branched [Hor83]. Brazil [dS88]. Breaking [LP79]. Bremsstrahlung [BKdCW81, BGJ72, Dar86, DMMS86, MM73, PN73b, PN73a, Tsa74]. Breuer [vH81]. Brian [EC87]. Broadhurst [Bro85]. Brodsky [ABDK69, AB70, BHP62, BS67, Bro69, Bro70, Bro71, Bro72, BGJ72, BR72, BRS73, GBB72, GBB73]. Broughan [Bro82]. Brown [AB83, BH79]. Bruce [GP82]. Brussels [ESP88]. Bryan-Jones [BJ87]. Bubble [LS79, LS81, LS80].

Buchberger [Bra90, CvH83, MN89c, vHC83, MN89b]. Buchberger’s [GM85, GM88]. Buckling [LH70, RT85]. Budapest [LPSV82]. Bui [KPB85]. Built [FF82b]. Bulgaria [GKRR85a, GKR85, GKRR85b]. Bulgarian [GKRR85a, GKR85, GKRR85b]. Bull [Kea85]. Bulletin [ABD85, Abi83, ASW89, BH72, BC81, Cal74, CH85, Cap86b, Cap86a, Dav82a, Dav82b, Fit73, FF82a, Gat85a, Har78, Har79b, Hea72b, Hea72c, HvH83, IKRT89, KS81a, Kre88, KO83, Kry84, KR85a, KR85b, eS85, Loo72, Mac89a, MHGG80, MF83, MH85, Maz85, McI85, MG89, MM89, Nam86, Nor78, NW83, Nor80, PH81, PH83, Rod84, RT88b, Sch82a, Sch88a, Sch83b, SvHH81, Ste82, GH82, vdH86a, vdHHvH87, vH80, vHH82, WR79]. Bundeswehr [EH86]. Burroughs [GK79].

C [Mac89b]. CA [Kau73, Rin71]. CAGe [Vos89]. Cairo [Wanxx]. Calcul [Duv87, Fit87, Tou87]. Calculate [KF87]. calculating [AM90, BGK86]. Calculation [Cal72b, Car75, DR85, Dul87, Ede81, GKRR85a, GKR85, GKRR85b, Hea71a, IKRT89, MM73, Nik87, PRH90, vdH86b, Gro88b]. Calculations [Bir77, CB76, FC79, DMR88, Ger80a, GTS80, GK85, IK87, Rod84, Sei91, ST89b]. Cálculo [dS87, dS88]. Calculus [Hap89, Kad96, Sch82a, SHM87, ST89a]. California [BH84b, HM75, KKM81, SH72]. Calmet [BC90, Cal72a, Cal72b, Cal74, CvH83, vHC83]. CALSYF [Cap89]. CAMAL [FC79]. Cambridge [Mac86b, McC84b]. Campbell [BCTC75, Cam67, Cam68, CH70, CCH70, Cam74, CFW87]. Can [CDS88, Fit89b]. Canada [KKM81]. Canonical [Kar85, Sht75]. Caprasse [Cap84, CH85, Cap86b, Cap86a, CDS88, Cap89, Cap90, DCM^{+}89]. Carlson [Car80]. Carroll [Car73, Car75]. Case [BBA73, Hea82, WR79]. Cassing [BNdP^{+}87]. Catastrophe [WD84]. Causmaecher [BKdCW81]. CBPF [dS88]. Cejchan [CNxx]. CEN [GV78]. CEN-Saclay [BLMN78]. Center [Ede81, FGPF88, Hus81, Par71]. Central [LPSV82, Per84a, Per86a, GL83, GZ84]. Centre [DT81, dSS85]. CERN [GG78, Köl84, Köl85b, Köl84]. Certain [GV86, Köl85a]. Chaffy-Camus [CC88]. Chain [IC90, Ito90, Kee85, YP88, YP89]. Chains [Kah69]. Champs [Cap89]. Chang [WCvH84]. Channel [MMK81, Par71]. Channels [PLR85]. Characteristics [DR85]. Characterization [Per84a]. Charge [AZZ87]. Charged [BH84a, BGHM88, BGJ72, Cam68, CCH70, Tsa74]. Charged-Pion [CCH70]. Charmed [KS88]. Charmonium [CS77, Gro88c]. Chaudhuri [BCTC75]. Cheaper [DP85]. Checkerboard [DM85]. Chekanov [BCM^{+}95]. Chem [CCEL84, HT87]. Chemical [IY77, Ina80, JPS87, MMN89, WTV87, WTVNF87]. Chemistry [Ogi82]. chemostat [Kee83]. Chen [Pic69, MMH88]. Chenin [Fit87]. Cherry [TC88]. Chew [Ger80b]. Chew-Low [Ger80b]. Chikyu [IC90]. China [zZlH84]. Chinnick [CGGK86]. Chris [HWH89]. Cim [MBG87]. Cimento [Gar80, HG84]. Circ [BCDS87]. Circuit [LOG85]. circular [KG75, LH70]. Ciência [RdS88]. Claire [DDDD85, Dic85]. Clarification [Abi83]. Clark [CCH70]. Class [BSW88, BGH88, Bro85, HvH77, RT87b]. Classes [Rin71]. Classical [GV86, GVxx, GV78, Hie84a, McC84b, WTVNF87, Gro88a]. Classification [GSZ85, GZ90a]. Clebsch [GKR85]. Cliath [Garxx]. Clifford [BCDS87]. Cline [CAH90]. Closed [BBR85, Köl82, WTV87]. Closed-Form [BBR85]. CNRS [Per78, Fit87]. Coalescing [CCF84]. Code [FP85, Gat85a, Gat86, GvH89, HvHS81, KK89, LOG85, SvHH81, vH83, vHHGV89, WCvH84]. coded [GK78, vdHvHG87]. codimension [Kee83]. Coding [Nik87]. Coefficient [KT85, NK84]. Coefficients [GKR85, GP74, HT87, Tou79, vH80]. coexistence [Kee83, Kee85]. Cohen [CLS76, CB76, CS77, FC79, CFÅ84, Coh89]. Cohomology [Ng89]. Coincident [RT87b]. Colagrossi [BC81]. Cold [DLS88]. collection [Tha89a, Tha89b]. Collective [HG84]. College [Bir77, Mac86a, Tha89a, Tha89b, Wanxx]. Colliding [Pes73]. Collinear [CCEL84]. Collins [CvH83, vHC83]. Collision [LD87, MS84, Bro82]. Collisions [BNdP^{+}87, GGRxx]. COLOR [KR88]. Colorado [FBC86]. com [dS87]. Combinatorial [RT87a]. Combined [Som77]. Comm [BH79, Cal72b, DR85, EFK85, FGK89, FK86, GvPV79, Ger80a, GK80, Hap89, Ito85, IK85, Ito88, KK89, Köl81, KS84a, KS84b, Lis84, NK84, SKxx, ST89b, YP88, YP89]. Comment [BH72]. Comments [Mac89a, QGxx]. Common [Gri78]. Commun [BH88, KR88, PRH90, Stu88, SGT90, zZlH84]. Communications [Art95, BS98, BCM^{+}95, Bil92, Boc92, Eas91, Eas93, Har97, IK96, Ito94, Kad96, Ren92, Sch82b, Sch86, Sei91]. commutative [Rod84]. commutator [Sei91]. Comp [AD88, Alf82, AH73, BBA73, BS86, BH88, BH79, Cal72b, CH70, DR85, dS89, EM82, EFK85, EH87, FGK89, FK86, Fle71, FH74, GvPV79, Ger80a, GSZ85, GZ90a, GK80, Hap89, Hea76c, Ito85, IK85, Ito88, KK89, Köl81, Köl82, Köl83b, Köl83a, KS84a, KS84b, Köl86, KR88, Lie75, Lis84, MHK86, NZ82, NSJ85, NK84, PRH90, Pie84, PA86, RN86, SKxx, Sto75, Stu88, SGT90, ST89b, YP88, YP89]. Compactifications [DR86a]. Compactified [DR85]. Company [KK89, Köl81, NM89a, ST89b]. Comparative [Smi79]. Comparison [GZ84, Gre84, Hus81, KPB85, NW83]. competing [Kee85]. Compilation [KRR87]. Compiler [GH79, HGB82]. Compiling [Fit89a]. Complete [Cap86a, KKP^{+}82, Per84b, Per84c]. Completion [Nor90]. Complex [Hie88, KR82, LV84]. Complex-Energy [LV84]. Composite [GBB72]. Comprehended [Sht77]. Compressed [PLR85]. Compression [HvH85, LH70]. Compton [BHP62]. Computacão [RdS88]. Computation [AH73, AB89, BGK86, BS87, BCRS89, BHPS86, Bro71, BH79, Cam67, Cap90, DST88, Dew89, Dic85, DJ89, DR86b, Eis90, EC87, Fit85b, FH87, FH74, FGPF88, FGP89, GM88, GKK87, Hea76c, Hea76b, Hea79, Hea81, Hea85, HI89, Ken88, Köl85a, LV84, LP90, Mac87, Mar85, MS85, MMN88, NM89a, NZ87, Ng89, RT87b, RT89, Sag88, Sav90, Sch85b, ST89a, UYSA89]. Computational [DL86, Gla82, vH80, WD90]. Computations [Dau79, DJ80, DJRR81, Gra81, Hea77, KRxx, SKxx]. Compute [Gri78]. Computer [Art95, BS98, BBA73, BCM^{+}95, Bil92, BC90, Boc92, BDM^{+}87, BJ87, Cal74, CvH83, CH70, Cap90, Coh89, DMR88, Dav81, Dav88, DST88, Vos89, Eas91, Eas93, EKR89, EM82, FGGS87, FK87, Fit85a, Fit89a, Fle71, FB89, GL89, Gat85c, Gat85b, GZ90b, Ger80a, GSZ85, GKRR85a, GKR85, GKRR85b, GSSZ87, GKK87, GK89, GZ90a, GL83, GvH89, Gri75, GK78, GK79, HWH89, Har79a, Har97, Hea71b, Hea72a, Hea73, Hea76c, Hea78, HW81, IK96, Ito94, Kad96, Kah69, Kre88, Lev71, LOG85, McC87, MN89b, Mur85, NM89b, Nor80, Ogi82, Ogi89, Ray87, Ren92, Rin71, SHM87, Sch82b, Sch86, Sch87, Sch88b, Sei91, SGM87, SM72, SL74, Sto77a, Sto75, Sto77b, vdH86b, vHvH88, vH81, vH82, vHC83, vH88, vH89, Wanxx, WH89, DJ89]. Computer-aided [SM72, DJ89]. Computeralgebrasysteme [WKL88]. Computerized [EH87, NA79, Sch77]. Computers [ER90, GTS80, GMS86, Gri78, Kot86, Nik87, NA79, Ogi82, Ogi89, RT85, YA87]. Computing [Cam74, CAH90, CLS76, CFÅ84, DDDD85, Eas93, ER90, GQ86, GZ84, Hus81, KPB85, Kau73, KS81b, Köl81, KRR88, KR88, Mac86a, MMN89, NS85, QG84, Sch85a]. Concavity [Kea85]. Concentration [KM80]. Concepts [GKRR85a, GBB72]. condition [Hie87b, Hie87c, Hie87d, Hie88]. Conditions [AZZ87, GZ90b, GKZ90]. Cone [Dau83]. Conference [DL86, DL88b, ESP88]. Confined [MHK86]. Conformally [LPSV82, Per84b, PLSV84, Per84c, Per86b]. Conical [Köl81]. Connecting [WTV87]. Connection [HH80]. Connor [CCF84, CCEL84]. Consequences [Cam68]. conserved [Ito94, IK85, Ker83, KE86a, KE86b]. Conserving [PFF83]. Consistency [GMS90]. Consistently [BSW88]. Constales [BCDS87, BCRS89]. Constant [HGDR84, Tou79, WD77]. construct [Hie85]. Constructing [Sch80]. Construction [Her83, PR84, QGxx, QG84]. constructive [Ren92]. Contact [EFK85]. Context [HvHS81, Smi79]. Continued [WTVNF87]. Continuity [ST89a]. Continuous [GP82]. Continuum [HH80]. Contribution [ABDK69, BS67, EH86, Fox71, LS79, LS80, LS81, Par68, SKxx]. Contributions [Adk85]. Control [DG75, HvH84, Kah69, UYSA89, WD77]. Controlled [DL88b, HvH85]. Conversational [NS82]. conversion [Vos89]. Convexity [Ken89b]. Conwell [CBR84]. Cooper [FC84]. Cooper-Type [FC84]. Coordinate [Car80]. Coordinates [BS86, Eas91, Han87, Hap89, Per86a]. Copeland [FC84, FC85]. Core [BBF^{+}82]. Cornell [KC72, KC73]. Corporation [Hea87]. Correction [Adk83, HG89]. Corrections [AB70, Cal72a, Cam67, HKY69, Kau73, KC72, SH74, TSM86, Van76]. Correlation [NK84]. cos [Köl83b]. Cosmological [Bir77, CDS88, Dau83]. Cosmologies [dRD88]. COSMOS [NS82]. Cotton [FBC86]. Couch [EC87]. Coulomb [BBR85, BH88]. Counting [GGRxx]. Couple [vH89]. Coupled [BMH87, BSW88, GZ90a]. Couples [JO84]. Coupling [HGDR84]. Coupling-Constant [HGDR84]. Covariant [ABD86b, Gar80, Geo68, RT87b, SH81]. Cowan [CG79]. CP [MF83]. CP/M [MF83]. CPT [Per78]. Creating [Ker86]. Creep [LH70]. Creeping [BBA73]. Critical [Nor90]. Critical-Pair [Nor90]. Critical-Pair/Completion [Nor90]. Cross [Cun75, FGK89, FC84, FC85, KG75, Kot86, KS88, MM73, PN73b, Pes73, TH65]. Cross-Section [Kot86, TH65]. Cross-Sections [MM73]. Crossed [Fox71]. Cryst [SW83, SK83]. Crystal [LV84]. CSAV [CNxx]. CTS [KR87]. Cultura [RdS88]. Cumulation [HvH84]. Cung [Cun75]. Current [HGB82, ZY75]. Currents [Ker83]. Curtis [CCF84, CCEL84]. Curvature [Gro88b]. Curve [Kob84]. curved [WW86]. Curves [GV86]. curvilinear [Eas91, Hap89]. Cvitanovic [KC72, KC73]. Cycle [KT85]. Cycles [LP90]. Cyclotomic [BD88]. Cyfronet [Ozixx]. Cylindrical [LH70, McC83, Rao85, ZRS87]. Cylindrically [BSW88, McC82].

D [Eas91, Gar84]. Dan [Fla86].

Dangelmayr [WD84]. dans [Cap89].

Darbaidze

[Dar86, DMMS86, DMR88, DR89, MDMS87]. Darstellung [LM73]. Data [DP85, Gri76b, HM75, Köl84, Köl85b, PH83]. Data-Structures [Gri76b]. Dautcourt [Dau79, DJ80, DJRR81, Dau83]. Davenport [ABD85, ABD86a, AD87, AD88, ABD89, BDS88, BDM^{+}87, BD88, Dav81, Dav82a, Dav82b, DP85, Dav88, DST88, ND79]. David [BH72, Har87, HWH89, Hap89, Har97, Sto75, Sto77b]. Davis [SH72]. Davydychev [ABD86b]. Debugging [KR85a, KR87]. Decay [AB83, Adk85, SH74]. Decays [BLMN78, LMSZ83, TSM86]. Decomposition [BCRS89]. Decoupler [NMJ84b]. Deep [BGJ72, Garxx]. Deep-Inelastic [BGJ72]. Deferred [Gri77b]. Definite [Köl85a]. Definition [Gri76b]. Deflection [Som77]. Deflective [GS77]. Deformation [Kag88, Rao85]. Deformations [ZRS87]. Delanghe [BCDS87]. Delbrückstreuung [Kra73]. delivery [Fit90]. Della [DT81, DT84, Fit89a, GvH89, MN89b, NM89b]. Della-Dora [DDDD85, MN89c]. Demaret [CDS88, DCM^{+}89, dRD88]. Demichev [DR85, DR86a]. Denominator [Gri78]. densities [Ito94, IK85]. Density [KS84a]. Department [BH84b, Cal72a, CLS76, CB76, DLS88, Fla86, Gat85c, Gat85b, HvH77, HvHS81, MG88, NS82, Ono79, Sht77, vdH86b, vHvH88, vH81, vH88, vH89]. dependent [KE86a, WTV87, GVxx]. Dept [Bir77, Boy78, FBC86, Kah69, Kam69a, Kam69b, Ken89a, Ken89b, NMJ84a, Rao85, Squxx, Sto75]. D’Equations [Tou87, DT81]. deren [Mac73]. Derivation [DJ89, Han87]. Derivatives [HvH83, RT87b]. descents [CFW87]. Described [WTVNF87]. Description [Cap84, Cap86b]. Design [CAH90, Gat85b, UYSA89]. DESIR [Tou87]. Determination [BS86, BHP62, NMJ84b, SH81]. Determining [BT63, EFK85, FK87, KF87, Sch82b, Sch83a, Sch85a, Sch85b, Sch86]. Deuflhard [WD90]. Deuterium [NMJ84b]. Deuteron [Geo68]. Development [Hea82]. Dewar [Dew89]. Dhar [DM85]. Diagonal [HT87]. Diagram [LS79, LS80, LS81]. Diagrams [Bro85, Cal72b, CH70, FG84, GvPV79, KRxx, KR88, Stu88, SGT90]. Diameters [Hor83]. Dicrescenzo [DDDD85, Dic85]. Dielectric [CBR84]. Difference [DT84, GS90, GL89, LD90, Lis84, MS85]. Different [DR86a]. Differential [Abb87, BGKN89, Bil92, Cun75, DR89, EFK85, FK87, GK89, Har97, Her83, IK85, KF87, Mac88, MHK86, Rin71, RT88b, RT89, Sch82b, Sch83a, Sch85a, Sch85b, Sch86, Sch88b, TS76, Tou79, TH65, Wat76]. Differentielles [DT81, Tou87]. Diffusion [Gre84, Ken88, Ken89a]. Digital [GTS80, Rin71]. Dilute [BT78]. Dimensional [DR86a, DGH^{+}86, Hie83b, IK87, KR82, Pic69, Sch82c, zZlH84, CDS88, DCM^{+}89, IK87]. Dimensionally [GvPV79]. Dimensions [Bro85, GDRH85]. DIMREG [IK87]. Dirac [BSW88, BBR85, Gar84, IKRT89, Ker83]. Dirac- [Gar84]. Direct [BS86, Hie87a]. Discrete [Ito90]. Discrete-Time [Ito90]. Disintegration [Geo68]. Disk [vH80]. Dispersion [Feu84, SW83]. Dissipative [SS84]. Distribution [DR89, KM80, KS84a, KS84b, NK84, SW83]. Diver [DLS88, DL86, DL88b, DL88a]. Divergence [EH87, GP74]. Divergences [Cal74, Van79]. Diverses [Duv87]. Division [Hea79, Köl84, Köl85b]. do [Dav82b]. Doing [HI89]. Domain [BHPS86, DG75]. Dominated [McC81]. Dominic [Ede81]. Dominique [DDDD85, Duv87]. Dora [DT81, DT84, Fit89a, GvH89, MN89b, NM89b]. Dorfi [DD85]. Dorizzi [DGH^{+}86, GDRH85, HGDR84]. Douglas [ZY75]. Douir [ADB^{+}84a, ADB^{+}84b]. Drawings [TC88]. Drska [LD90, LD87]. Drury [DD85]. Dual [KM84, WTV87, WTVNF87]. Duality [Hie83a, HGDR84]. Dubna [AZZ87, BOH78, BGKN89, Ger80b, GMS86, KF87, KRR87, KRR88, MS85, Nik87, vH87, VR87, ZNR78]. Dubowsky [DG75, WD77]. Dudley [DJ89]. Due [Kam78]. Dufner [ABDK69, DT69]. Dulyan [Dul87]. d’un [Lux75]. Duncan [DR86b]. d’Une [Cap84]. Dunham [HT87]. During [CGGK86]. Dust [ND79]. Dutch [vH88]. Duval [DDDD85, Duv87]. Dynamic [Ada83, BGHM88, DG75, KR85a, WD77]. Dynamic-Debugging [KR85a]. Dynamical [LExx, Mac86b]. Dynamics [BH84b, FBC86, GQ86, Har87, IC90, Ito90, QG84, WTVNF87]. Dyson [KC73].

e- [Cun75, GGRxx, GG78, Per78, TH65]. Earles [Ear70]. Earth [Kam78, Kam69b]. Earth-Moon [Kam69b]. Eastwood [Eas91, Eas93]. ed [CFÅ84]. EDAM [CAH90]. Edelen [Ede81, Ede82]. Edge [CCEL84]. Editing [KO83]. Editions [Fit87]. Edneral [AE86, AE87b, AE87a, EKR89]. EDS [Har97]. Efetivo [dS87]. Effect [PLR85]. Effective [BD88, Coh89, dSS85]. Effects [Boy78, Gar84, Hor83, JPS87, Kam78, KMS86, SW83]. efficiency [Dav82a]. Efficient [Gri78, Gri76c, Gri77b, Gri77a, NSJ85, SKxx, WD90]. Egypt [Wanxx]. Eigenparameter [Sha87]. Eiichi [LOG85, MG89]. Eikelder [KE86a, KE86b]. Eikonal [SH81]. Einstein [Sch88a]. Eire [Garxx]. Eisenberger [Eis90]. Eißfeller [EH86]. Eitelbach [Eit73]. Elastic [Eis90, Garxx, Gar84, Kag88, Say87a, Say87b]. Electric [Som77, VR87]. Electric-Magnetic [Som77]. Electrodynamic [AB70, Bro70]. Electrodynamics [Bro69, Bro71, Bro72, BR72, BRS73, FH74]. electromagnetic [Hie85, Van79]. Electron [ABDK69, Cal72a, Cam67, Car73, Car75, Ear70, Garxx, Gar80, GS78, HKY69, KC72, LS79, MM73, PN73b, Par68, Gar84]. Electron-Electron [PN73b]. Electron-Electron-Bremsstrahlung [MM73]. Electron-muon [Garxx]. Electron-Nucleon [Garxx]. Electron-Positron [Cam67, HKY69]. Electron-Production [Ear70]. Electrostatic [PSA84]. Electrovac [McC82]. Elektron [PN73a]. Elektron-Elektron-Bremsstrahlung [PN73a]. Element [FH87, WCvH84]. Elementarprozess [PN73a]. Elementarteilchen [Mac73]. Elementarteilchen-Physik [Mac73]. Elementary [DP85, Eit73, PN73b]. Elements [BH88, GKRR85a, GKR85, GKRR85b, HT87, YP89]. eler [eS85]. Eleutério [EM82]. Eleven [DR86a]. Eleven-Dimensional [DR86a]. Eliseev [EFK85]. Elishakoff [EH87, EC87, LExx]. ellipsoidischen [Kra82]. Elliptic [AH73, Kea85]. Elliptical [Kot86]. Embedding [KPB85]. Energy [Bro70, DR86a, Vos89, Geo68, Ger80a, GKRR85a, GKR85, GKRR85b, Han87, Hea81, LV84, Lie75, PSA84, QGxx, QG84, ZY75]. Energy-Momentum [ZY75]. Eng [Ede82, Rao85]. Engineering [CAH90, EH87, MMN89, NS82, RA89, Squxx]. England [BDS88, HWH89]. Enlarging [BHPS86]. Entwicklung [Kra82]. Environment [Dew89, Ede81]. Ephemeris [Kam78]. EQSHELL [Boc92]. Equation [Abb87, CS77, DT84, Fle71, FT75, Gat90b, Ker83, LSS86, Mac88, NS86, RN86, RT88b, Sch82c, Sha87, Sto75]. Equations [AH73, BG87, BGH88, BBA73, BDS88, BGKN89, Boc92, BGK86, Bro82, DR89, DJ89, EFK85, FK87, GKZ90, Ger80b, GSZ85, GK89, Gri74, Her83, Hie87b, Hie87c, Hie87d, Hie88, Ito94, IK85, KM84, KMH88, KF87, LD87, Loo72, MHK86, McC83, McC84a, MBG87, MMK81, PLSV84, McC84b, Ren92, Rin71, RT89, Sch82b, Sch83a, Sch84, Sch85a, Sch85b, Sch86, Sch88b, Sto77b, TS76, Tou79, WTV87, Wat76]. Equatorial [Boy78]. Equilibria [MMR87]. Equilibrium [BS86, Boc92]. Equivalence [EM82]. Ernst [Per86a]. error [Bam88, HvH84, Sto77a]. ESPRIT [ESP88]. Esteban [ER90]. Estimate [Par68]. Estimates [Ada83]. Etude [Lux75, Tou87]. Euclidean [BS87, Ken89a]. European [DL88b]. Europhysics [DL86]. Evaluating [Ito85]. Evaluation [Adk83, AE87a, AB89, Bro85, Cal72a, CFW87, Gri77b, Har77, Har79b, Hea86, Köl85a, LD87, LSS87, MM89]. Evaluations [GvPV79]. Evelyne [Tou79, Tou87]. Even [Maz85]. Evolution [AH73, GZ90b, GKZ90, GSZ85, GSSZ87, Ito94, Sch84]. evolutions [Ren92]. Exact [BH84b, BHHS87, BG87, BH88, McC87, McC84b, WTVNF87]. Examples [BGK86, FBC86]. EXCALC [CDS88, dRD88, SHM87]. Exchange [BCTC75, FT75]. excitable [Kee89]. Excited [FC85]. Exclusive [DMR88, DR89]. Executing [KO83, WTVNF87]. Execution [PH83]. Existence [Cam68, Ken89b, Wat79]. Expansion [GP74, Gri78, Gri76a, Gri76c, Gri77b, Gri77a, HvH85, Pie84, Smi79]. Expansions [GKK87, GM84, KS84a, NK84, PB90, Vor77, WTV87]. Experience [GH79]. Experiment [HKY69]. Experimental [CGGK86, NMJ84a, ZNR78]. Experiments [AW88, Bro70, Sch87, vdH86b]. Explanation [DR89, Mar85]. Explicit [CFW87, Köl85a, Mal82, PSA84, WD84]. Exploration [AD87]. Exponential [Kar85]. Expression [Gri77b, HvH85, LS79, vHH82, WTVNF87]. Expressions [BBR85, GvH89, HT87, Hea86, Kar85, Köl82, RT87a, RT88a]. Extended [HL73, KO83]. Extending [GDRH85]. Extension [Cap84, Cap86b]. extensions [Spi87, GH82]. Extensive [SKxx]. Exterior [BGHM88, Har97, Sch82a, SHM87]. External [Dul87, KMS86]. Extrapolation [KPB85].

F [Eas93, CC88]. Fabrication [GS78]. Face [Wanxx]. Facilities [HvHS81, HvH83, MG88, vdHHvH87, vHvH88]. Facility [Gat85a, Gat86]. Factor [Car75, SK83, vH81]. Factorisation [ABD85, Abb88]. Factorization [AD87, HvH85, MG89]. Factors [DT69]. Facts [BHM86]. Falck [FGK89]. Families [Her83, Hie83a, RT89]. Fancies [BHM86]. Farrelly [CCF84]. Fast [Dav82a, GM85, IKRT89]. Fazio [FC84]. Federbush [Ker86, KE86a, KE86b]. Fedorova [EFK85, FGGS87, FK87, KF87]. Feldmar [FK86]. Fermi [KKM81]. Fermion [ANGK^{+}87, BBF^{+}82]. Fernuniversität [GM85]. Feuillebois [Feu84]. Feynman [KC73, AE87a, Bro85, Cal72b, Cal74, CH70, Dul87, GvPV79, KRxx, KR88, Stu88, SGT90]. Feynman-Dyson [KC73]. Fibres [BT78]. FIDE [LD90]. Field [BH84a, BG87, BMH87, BGHM88, BGH88, DLS88, Dul87, Gro88c, IK87, KMS86, McC83, McC84a, MBG87, McC84b, ST89b, VR87]. Fields [Abb88, BSW88, BW90, BZZ86, DDDD85, Ede81, Ede82, Kad88, KM84, KR82, Lie75, LPSV82, Per84b, Per84c, Per86b, ZNR78]. filaments [Kee89, Kee90]. Filled [Rao85, ZRS87]. Filters [SH72, SM72, SL74]. finding [Ito94, IK85, Mal82, McC84b]. Fine [Ken89b, Tal84]. Finite [FH87, GP82, Lis84, Sto75, WCvH84, LD90]. First [KK85, Kau73, Köl81, LSS86, Pie84, SS84, Sch85b, Sch86]. First-Order [KK85]. Fitch [FC79, Fit90, Fit73, Fit81, Fit83, Fit85b, Fit85a, Fit87, FH87, Fit89b, Fit89a, GvH89, MF83, MN89b, NM89b, MN89c]. Five [AB83, FGK89]. five-parton [FGK89]. Fixed [Alf82]. Flat [LPSV82, Per84b, PLSV84, Per84c, Per86b]. Flatau [FBC86]. Flath [Fla86]. Fleischer [Fle71, FGM73, FT75]. Fletcher [PFF83]. Flip [IC90, Ito90]. float [KS81a]. Flow [BT78, JO84, Kot86, Rao85, Sav90, Sch77, ZRS87]. Flows [BBA73]. Fluid [BT78, BBA73, Feu84, FBC86, Kee90, LD87, MS84, Rao85, Sch77, WW86, ZRS87]. Fluid-Filled [Rao85]. Fluids [MMR87, Ede82, Kot86]. Fluorescence [JPS87]. Flutter [EH87]. Fock [LOG85]. Focusing [Som77]. Focusing-Deflection [Som77]. Fogelholm [FF82a, FF82b]. Fokker [Bro82, Hie85]. Fontannaz [ADB^{+}84a, ADB^{+}84b]. food [Kee85]. Force [GBB73]. Forces [JO84, PSA84]. Form [BBR85, DT69, Ito85, Kar85, Ren92, Sht75, Tou79, WD84, BS98]. Formal [HvH83]. Formalism [Car73, ER90]. Formel [Duv87, Fit87, Tou87]. Formelles [DT81, Tou87]. Formula [Kat85, Wat76, Wat79]. formulae [CFW87, PSA84, Tao90]. Formulas [Som77]. Formulation [Gar80, GS77]. Formule [vH88]. Fortes [ANGK^{+}87, BBF^{+}82]. Fortran [Ina80, KK89, Ono79, Sha87, vdHvHG87]. Fortran-Based [Ina80]. FORTRAN-Coded [vdHvHG87]. Fortschritte [Sch80]. Foster [FB89]. Foundation [Eis90]. Foundations [McC87]. Four [CCF84]. Fourier [DDDD85, Dic85]. Fourth [FH74, SH72]. Fourth-Order [SH72]. Fox [Fox71, FH74]. Fraction [WTVNF87]. Fractionally [BGJ72]. Frame [Bro72, BR72, BRS73]. France [DDDD85, Dic85]. Franceschetti [FP85]. Franquelo [FGPF88]. Franz [FF82b]. Fredholm [Loo72]. Free [Car80, MMR87, GS77]. Freire [FGPF88, FGP89]. French [Cap89]. Freund [DR86a]. Freund-Rubin [DR86a]. Frick [CFÅ84, FF82a, FF82b]. Friedrich [BH84b, Ren92, WH89]. Fritz [BT78, Sch83b, Sch83a, Sch85b]. Frobenius [DT84]. Frobenius-Boole [DT84]. Fröman [CFW87]. Fry [PFF83]. Fujimoto [FG84]. Function [Alf82, Cap86a, Cap90, Hea74b, KS81b, KR82, KS84a, MM89, Pie84]. Functionals [KE86a]. functions [AM90, ABD86b, BH72, Dav81, GKK87, Hea72c, HvH77, Köl81, Köl84, LV84, Maz85, Möl89, PA86, Wat85]. Further [Cal72a]. Fusion [DL88b, MDMS87, Wat85]. Fuzio [FC85]. Fys [Van76].

g [Car75, LS79, LS81]. g-2 [LS79, LS81]. g-Factor [Car75]. G. [GH82]. Gaemers [GGRxx, GG78]. Gamero [FGPF88, FGP89]. Gamma [DR89, Hea71a, IKRT89]. Gamma-Distribution [DR89]. Gamma-Matrices [IKRT89]. Gammel [FGM73]. Ganzha [GS90, GMS90, GL89]. Garavaglia [FG84, Garxx, Gar80, Gar84]. Garcia [GV86]. Garrad [GQ86, QGxx, QG84]. Gastmans [GGRxx, GvPV79]. Gatermann [Gat90b, Gat90a]. Gates [GvH85, Gat85a, Gat85c, Gat85b, Gat86, vHHGV89]. Gauge [AE86, AE87b, AE87a, ABD86b, BH84a, BH84b, BHM86, BHHS87, BG87, BMH87, BSW88, BGHM88, BGH88, BKdCW81, KR82, KRxx, KR88, LM85, McC83, McC84a, McC87, MBG87, MMH88, McC84b, SHM87]. GCD [ADB^{+}84b, Hea72b]. GCDs [Hea79]. Gebauer [BGK86, GM85, GM88]. Gen [dRD88, HG89, SS84]. General [CLS76, FC79, CFÅ84, Dau79, DJ80, DJRR81, Ger80b, IY77, Mac89a, Ozixx, Per84b, Per86b, McC84b, SHM87]. generality [Dav82a]. Generalized [FK86, NZ87, SW83]. Generate [AH73, DP85]. Generated [BBA73]. generation [Boc92, GvH85, Gat85a, Gat86, GZ90b, LOG85, Mac89b, vdHvHG87, vH87, WCvH84]. Gent [Vos89, Tha89a, Tha89b]. Gentran [Gat85a, Gat85c, Gat85b]. geodätischen [Kra82]. Geodesy [EH86]. Geodetic [RT87b]. Geometric [Her83]. Geometrical [DR85]. Geometry [Dau83, Tha89a, Tha89b]. Geophys [KK85, Kag88]. Geophysical [ANK81, FBC86, Say87b]. George [Geo68]. Gerald [Mag81]. Gerdt [BGKN89, FGGS87, GZ90b, GKZ90, Ger80a, GTS80, Ger80b, GSZ85, GKRR85a, GKR85, GKRR85b, GMS86, GSSZ87, GKK87, GK89, GZ90a]. Gerez [SGM87]. Gerhard [WR79]. German [WKL88]. Germany [DD85]. Gervois [GP74]. Gibson [CGGK86]. Gihodo [NM89a]. Gini [KT85]. GITA [BCM^{+}95]. Giuseppe [Maz85]. Gladd [Gla82]. Gladkih [GL83, GZ84]. Glasgow [DLS88]. Global [Ger80b]. Gluon [AE87b, Dar86, DMMS86, KKM81, MDMS87]. Gluonic [ABD86b, Dul87]. Go [WTV87, WTVNF87]. Goldman [GvH89, vdHvHG87]. Golley [GP82]. Góngora-T [SGT90]. Good [GLB75]. Gordan [GKR85]. Gordon [Hie87d]. Goto [GS77, GS78, LOG85, MIG85, MG88, MG89]. Gould [GM84]. Gounaris [GG78]. Govorun [FGGS87]. Grad [Bro82]. Gragert [Gra81]. Grammaticos [DGH^{+}86, GVxx, GV78, GDRH85, HGDR84, HG89]. Grant [DG75, Kea85]. granular [Sav90]. Graph [Fox71, GK85]. Graphical [eS85]. Graphics [Mar88]. Graphs [Cal74, Dul87, PRH90]. Graudenz [FGK89]. Grav [BSW88, BGH88, dRD88, RT87b]. Gravitation [FC79, CFÅ84, McC87, Per84b, Per86b, SHM87]. Gravitational [BGHM88, Van79, ZNR78]. Gravity [BH84b, BHM86, BMH87, BSW88, Van76]. Graz [LP79]. Greatest [Gri78]. Green [ABD86b]. Greenland [Gre84]. Green’s [LV84]. Greiner [HG84]. Grenoble [DDDD85]. Grids [DD85]. Griffiths [CGGK86]. Grimm [GK80, KG75]. Griss [CG79, Gri78, Gri74, Gri76a, Gri75, Gri76b, Gri76c, Gri77b, Gri77a, GK78, GH79, GK79, HGB82, MHGG79, MHGG80]. Gröbner [MMN88]. Groebner [BGK86, vdH86b]. Ground [ANGK^{+}87]. Ground-State [ANGK^{+}87]. Group [Cap90, Vos89, DLS88, KRxx, KR88, LS79, LS80, QGxx, QG84, Woo89]. Group-Theoretic [KR88]. Group-Theoretical [KRxx]. Groups [YP88, YP89]. Grow [vH81]. Grozin [Gro88a, Gro88b, Gro88c]. Gruppentheoretische [Gat90a]. Guardiola [ANGK^{+}87]. Guerses [BG87, BGHM88, BGH88, MBG87]. Guide [DJRR81, HWH89]. Guilherme [BC90]. Gunion [BGJ72, GBB72, GBB73, GK85]. Günter [EH86].

H-C [LM85]. Hacinliyan [GH82].

Hadamard’s [Sch83b]. Hadinger [HT87]. Hadrochemistry [BZZ86]. Hadrons [ADB^{+}84a, Per78]. Hagen [GM85]. Haifa [Sht77]. Haim [FB89]. Half [Kag88]. Half-Space [Kag88]. Hall [FH87]. Halphen [GK89]. Hamada [BS86]. Hamiltonian [GKRR85a, GKR85, GKRR85b, GDRH85, Hie83b, HGDR84, Ren92]. Hamiltonians [BCM^{+}95, GVxx, GV78, Hie83a, Hie84b, Hie85]. Handling [Köl84, Köl85b]. Handy [Han87]. Hans [CH85]. Hans-Guenther [KO83]. Hao [zZlH84]. Haprer [Hap89]. Hard [BBF^{+}82, GP74, WW86]. Hard-Core [BBF^{+}82]. Harmonic [BCRS89, DJ89, Ken89b, KM84, Woo89]. Harmonically [WTVNF87]. Harold [CAH90, Dav81]. Harper [Har87, HWH89]. Harrington [Har79a, Har77, Har78, Har79b]. Harris [CAH90]. Hartley [Har97]. Hasenfratz [HH80]. Hashing [CG79, Gri77b]. Hauptaufgabe [Kra82]. Hayd [MHK86]. Head [MS84]. Head-On [MS84]. Hearn [BOH78, BH72, BHPS86, BHP62, BH79, CH70, FH74, GH79, HKY69, Hea71a, Hea71b, Hea72a, Hea72b, Hea72c, HL73, Hea73, Hea74b, Hea74a, Hea76c, Hea76a, Hea76b, Hea77, Hea78, Hea79, HN79, Hea80, HW81, Hea81, Hea82, HGB82, Hea85, Hea86, Hea87, MHGG79, MHGG80, MH85, TH65]. Heat [SM83]. Heating [DL88b]. Heavy [BLMN78, KKM81]. Hecht [BHHS87]. Heerwaarden [vHvH88, vHHGV89]. Hehl [BH84a, BH84b, BHM86, BHHS87, BMH87, BGHM88, BGH88, MMH88, SHM87, WH89]. Heidelberg [DD85, GM85]. Heiles [Hie83a]. Hein [EH86]. Heinrich [PRH90]. Heintzmann [SH81]. Heinz [Kre88]. Helicity [LSS87, Par71]. Helsinki [Hus81]. Henon [Hie83a]. Henon-Heiles [Hie83a]. Herbert [MN89a, MN89b, NM88, MN89c]. Herman [BT78]. Hermann [Her83]. Hess [HG84]. Hessians [vdHvHG87]. heterogeneous [WD90]. Hettich [HvH77]. HEUGCD [DP85]. Heuvel [vdH86a, vdH86b, vdHvHG87, vdHHvH87]. Hicks [PH81, PH83]. Hidden [Sha87]. Hideharu [NM89a]. Hidestune [Kob84]. Hierarchies [KE86a, KE86b]. Hietarinta [DGH^{+}86, GDRH85, Hie83b, Hie83a, Hie84a, Hie84b, HGDR84, Hie85, Hie87a, Hie87b, Hie87c, Hie87d, Hie88, HG89]. High [DR86a, Geo68, Ger80a, Hea81, JW86, Lie75, CDS88, Dav82b]. High-Accuracy [JW86]. High-dimensional [CDS88]. High-Energy [Hea81]. high-level [Dav82b]. Higher [NK84, SM72, SL74]. Hiroshi [KT85]. Hirota [HI89]. Hirota’s [Hie87b, Hie87c, Hie87d, Hie88, Ito88]. HLISP [Kan75]. Hodgkinson [HWH89]. Hoff [LH70]. Hogan [KMH88]. Hogreve [BH88]. Holland [KK89, Köl81, ST89b]. Hollkamp [EH87]. Holt [SH74]. Homogeneous [DT84, dRD88, Kag88, MG89]. Homsey [AH73]. Hopf [FGP89]. Horizontal [QGxx]. Horn [CCH70]. Horowitz [HM75]. Horwitz [Hor83]. Huey [SH72]. Hulshof [HvHS81, HvH83, HvH84, HvH85, SvHH81, vdHHvH87, vHH82, vHHGV89]. Hulzen [CvH83, GvH85, GvH89, HvH77, HvHS81, HvH83, HvH84, HvH85, SvHH81, SvH82, vdHvHG87, vdHHvH87, vHvH88, vH80, vH81, vHH82, vH82, vHC83, vH83, vH87, vH88, vHHGV89, vH89, WCvH84]. Hümeyra [Bil92]. Hungarian [Per84a, Per86a]. Hungary [LPSV82]. Husberg [Hus81]. Hydrogen [CGGK86, FC84, FC85, Tal84, VR87]. Hydrogen-Like [Tal84]. Hypergeometric [AM90]. Hyperons [TSM86]. Hypersphere [NSJ85].

I. [Hie87b]. Ia [EKR89]. iAPX [YA87]. ideal [Kee90]. Identities [LM85, McI85]. Identity [ZY75]. Idesawa [IY77]. IEEE [Ada83, Gri78, GK78]. Ignition [CGGK86]. IHEP [AE86, AE87a, ABD86b]. ihrer [Kra82]. II [BCDS87, LPSV82, SGT90, GVxx, Hus81, Per86b, SH81, Wat76]. II. [Hie87c]. III [Per84b]. III. [Hie87d]. Illinois [KKM81]. Ilyin [IK96, IK87, IKRT89]. Image [Ken89b]. Impact [MMN89]. Implantation [Lux75]. Implementation [FF82a, FP85, FF82b, Gat85b, GK79, Kan75, MN89a, NM88, NM89b, Sch82a]. Implementing [Fit83, NM77]. Improved [GK85, Hea72b, Hea85, Maz85, PFF83, Per84a, WD77]. Improvement [MG88]. Improvements [Nam86]. Inada [Ina80, MIG85]. independent [KE86a, Mac89b, TSM86, GV78]. Indicating [Wan85]. Induced [JPS87]. Industrial [BC81]. Inelastic [BGJ72, Garxx]. Inexact [Alf82]. Infinite [Bro72, BR72, BRS73, KE86a, KE86b]. Infinite-Momentum [BRS73]. Infinitesimal [Ker83, Pic69]. Informatik [LM73]. Information [HW81, Ina80, Kan75, MG88]. Informationstechnik [Gat90a, MMN88, MN89a, Möl89, NM88, WD90]. Infra [AE87b]. Infra-Red [AE87b]. Infrared [AE86, ABD86b, LM85]. Inge [FF82a]. Inhomogenieties [Feu84]. Initial [Dau83, DD85]. Inner [Adk85]. Inner-Vertex [Adk85]. Inst [Garxx, Vor77].

Instability [EH87, EC87, Gla82]. Installation [GM88]. Institut [BC90, DDDD85, Dic85, LP79, TS76]. Institute [Art95, CNxx, CLS76, CB76, CS77, DR86a, FF82b, GKRR85a, GKR85, GKRR85b, GL83, GZ84, Gro88a, Gro88b, Gro88c, IY77, Ina80, Kob84, KRxx, LPSV82, Per84a, Per86a, RT88a, Sht77, Wat76, Wat85, FGGS87]. Instituut [Van76]. Int [Ede82, Mur85, SM83]. Integer [Ono79]. Integrability [GZ90b, GKZ90, GSSZ87, Hie84a, HG89]. integrable [DGH^{+}86, GSZ85, GZ90a, GDRH85, Hie83b, Hie83a, Hie84b, HGDR84, Hie85]. Integral [Köl83b, Köl83a, Köl85b, Köl86, Loo72, Sha87, Sto77b, Tro89]. Integrals [AM90, AE87a, AB89, BS98, BBR85, CFW87, CCF84, FH74, Köl85a, LD87, LSS86, NS86, SS84, Sch85b, Sch86, Stu88, SGT90]. Integrated [Dew89]. Integrating [BC90]. Integration [Abb87, Bil92, CNxx, Dav81, Fit81, Har79a, HW81, KO83, NSJ85, Nor90, NM77, ND79]. Intensities [Tal84]. Intensity [SW83]. Interacting [GKRR85a, GKR85, GKRR85b]. Interaction [BH88, PLR85]. Interactions [BCTC75, Van79]. Interactive [GLB75, KO83, KR85b, eS85, NS85, ST89a]. Interchange [GBB73]. Interest [Bir77]. Interface [eS85, Mar88, vH87, ASW89]. Intermediate [BT63, Cam68]. Internal [Han87, MS84]. International [dSS85]. Interpolation [Möl89, Tao90]. Interpretations [CGGK86]. Interval [RT87b]. Introducão [dS88]. Introduction [HI89]. invariant [Hie87a, Par71]. invariants [Hie84b]. Inversion [Squxx]. Investigate [CDS88]. Investigating [GSSZ87, ZNR78]. Investigation [MS85, NMJ84a, Pic69]. Investigations [AZZ87]. Involution [Sch87]. Involving [Köl84, Köl85a]. ion [PFF83]. ion-Ure [PFF83]. IRENA [Dew89]. Isaac [EH87, EC87, LExx]. ISBN [Art95, Eas93]. Ising [DM85, IC90, Ito90, KMS86]. Isomerization [KM80]. Isothermal [CGGK86]. Isovector [Ede81, Ede82]. Israel [Sht77]. Iteration [Alf82, MIG85, MMK81]. Iterations [Cap90]. Iterative [WD84]. Ito [HI89, Ito94, Ito85, IK85, Ito88, IC90, Ito90]. Itzhak [LExx]. IV [Per84c]. IV. [Hie88]. Izd-vo [EKR89].

J-M [DM85]. J.I.N.R [AZZ87, BOH78, BGKN89, Ger80b, GMS86, KF87, KRR87, KRR88, MS85, Nik87, VR87, ZNR78]. J.I.N.R. [vH87]. J=0 [FT75]. Jacobians [vdHvHG87]. Jacques [BC90, Cal72a, Cal72b, Cal74]. Jaffe [BGJ72]. James [Dav81, Dav82a, Dav82b, DP85, DJ89, Eas91, Kee83, Kee85, Kee89, Kee90]. Jane [BJ87]. Janeiro [dS88]. Janet [Sch84]. Jann [DJ80, DJRR81]. Jansen [JW86]. Janssen [JPS87]. Japan [HW81, KMS86, MG88, Sch82c, NM89a]. Jaycor [Gla82]. Jean [DDDD85]. Jed [Mar78, MF83, MH85, Mar85]. Jeffrey [JO84]. Jessner [NSJ85]. Jet [Per78]. Jets [KKM81]. JINR [Dar86, DMR88, DR89]. Joel [Coh89]. John [Boy78, Fit73, FBC86, FH74, MF83, NMJ84b, NMJ84a, Woo89]. Joint [FGGS87]. Joseph [EH87, GP82]. Journ [AH73, BBA73, BS86, Cam68, CH70, DG75, Ede82, EM82, Feu84, Fle71, FH74, GQ86, GM84, Gre84, HT87]. Journal [AE87b, Har79a, Köl86, Wat79, WW86, WD77]. József [Kad96]. Jr. [Mag81, WTV87, WTVNF87]. JRSS [Mac89b]. Julian [DP85, PB90].

K-Expansions [Vor77]. Kachina [Sch77]. Kadlecsik [Kad96, Kad88]. Kaeppeler [MHK86]. Kagaku [ON90]. Kagan [KK85, Kag88]. Kahn [Kah69]. Kako [IK85]. Kaldor [SK83]. Kaltofen [FGP89, GL89, GK89, Tro89]. Kamal [KKM81]. Kamel [Kam69a, Kam69b, Kam78]. Kanada [Kan75, KS81a]. Kaneko [KK89]. Kaps [KPB85]. Karin [Gat90b, Gat90a, KO83]. Karlsruhe [BC90]. Karr [Kar85]. Kato [ANK81]. Katsura [Kat85]. Kauffman [Kau73]. Kawabata [KK89]. Kazasov [Kaz87]. KdV [GZ90a, Hie87b]. KdV-Like [GZ90a]. KdV-type [Hie87b]. Keady [Kea85]. Keener [Kee83, Kee85, Kee89, Kee90]. Keio [NS82]. Keller [ANGK^{+}87]. Kendall [Ken88, Ken89a, Ken89b]. Kenichi [UYSA89]. Kenji [Nam86]. Kenneth [FB89]. Kepler [NS86]. Kerner [KG75, SKxx]. Kernphysik [DD85]. Kerr [BGHM88, MBG87, Per84a]. Kerr-Like [MBG87]. Kersten [Ker83, KM84, Ker86, KE86a, KE86b]. Kessler [GK78, GK79]. keV [PN73a, PN73b]. Kevin [McI85]. Key [CG79, Hea85]. Khutornoy [GKZ90]. Kia [LOG85]. Killalea [KM80]. Killing [KM84]. Kim [KKP^{+}82]. Kind [Köl81]. Kindai [ON90]. Kinematic [Kah69]. Kinetic [Han87]. Kinetics [MMN89]. King [Bir77]. Kinky [BMH87]. Kinoshita [ABDK69, KC72, KC73]. Kitatani [KMS86]. Kiyokazu [ON90]. Kiyoshi [Ono79]. Kleiss [BKdCW81]. Knopoff [KK85]. Knotted [Kee89, Kee90]. Knowledge [BC90]. Kobayashi [Kob84, KMH88]. Kodaira [KKM81, KT85]. Kognitive [BC90]. Koh [KKP^{+}82]. Koinzidenzmessung [PN73a]. Kölbig [FK86, KS81b, Köl81, Köl82, KR82, Köl83b, Köl83a, KS84a, KS84b, Köl84, Köl85b, Köl85a, Köl86]. Konishi [NK84]. Konrad-Zuse-Zentrum [Gat90a, MMN88, MN89a, Möl89, NM88, WD90]. konstruktion [Gat90a]. Kordylewski [CGGK86]. Kornyak [EFK85, FK87, KF87].

Korteweg [Sch82c]. Kostov [GKRR85a, GKR85, GKRR85b, GKK87, GK89]. Kostova [BGKN89, GKK87]. Kotorynski [Kot86]. Krack [Kra82]. Krakow [Ozixx]. Kramer [FGK89]. Kraus [Kra73]. Kredel [BGK86, Kre88]. Kruse [KO83]. Kryukov [EKR89, IK96, IK87, IKRT89, KRxx, Kry84, KR85a, KR85b, KRR87, KR87, KRR88, KR88, KS88]. kubaturformeln [Gat90a]. Kühnelt [GK80]. Kunszt [GK85]. Kuo [HKY69]. Kuppers [KPT71]. Kutzler [WKL88]. Kyoto [Wat76].

Lab [Kan75, KC72, KC73]. Laboratoire [Dic85]. Laboratories [SL74]. Laboratory [HWH89, Ina80, KKM81]. Lacaze [BLMN78]. Laganà [CCEL84]. Lagrangians [CDS88]. Laing [DLS88, DL86, DL88b, DL88a]. Lamb [AB70, Fox71]. Lambin [LV84]. Laminar [Kot86]. Lanczos [WTV87, WTVNF87]. Landau [KS84a, KS84b]. Lang [LP79]. Language [BOH78, Dav82b, EKR89, Gro88a, Kan75, Kry84, Mur85, Sto77a]. Languages [GL83]. Laplace [Kaz87, Nor80, RN86]. Laplacian [BS87]. Large [ADB^{+}84a, ADB^{+}84b, Gri74, GBB72, GBB73, MMN89, MN89b, MN89c, ZRS87]. Laser [Gre84, JPS87]. Laser-Induced [JPS87]. Lathrop [SL74]. Latitudinal [Boy78]. Lattice [HH80, KR82]. Lattices [Ada83]. Laursen [LS79, LS80, LS81]. Lawrence [WR79]. Lax [Ito85]. Leading [ADB^{+}84b]. Lecture [Dav81]. Lectures [WH89]. Lee [LM85]. Lehigh [Ede81]. Lens [GS77]. Lenstra’s [AD87]. Lepage [LMSZ83]. Lepton [LS80]. Leptons [Tsa74]. Leringe [CLS76]. LET [Sht77]. Lett [ABDK69, ADB^{+}84a, ADB^{+}84b, BH84a, BG87, BGHM88, BKdCW81, BT63, Bro85, BR72, Cun75, DMMS86, DGH^{+}86, Ear70, FC84, GDRH85, GBB72, GK85, HH80, Hie83b, Hie84b, HGDR84, MM73, MDMS87, McC84a, MMH88, PN73b, PLSV84, SH81, WTV87, WTVNF87]. Letters [AB70, BZZ86, FG84]. Leuven [Van79, Van76]. level [Dav82b]. Levi [LH70, Lev71]. Lewis [Nor80]. Libration [Kam69b]. Lichtenberger [WKL88]. Lie [EFK85, Kam69b, Ker86, KF87, Sch82b, Sch88b]. Lie-Baecklund [KF87]. Liebermann [Lie75]. Liège [Cap84]. Life [KT85]. Light [Dau83]. Like [McC84a, MBG87, Tal84, GZ90a]. Limit [Har77, Har79b, LP90]. Limits [ZNR78]. Linear [BGKN89, CS77, DT84, GSZ85, Gri74, Gri76a, Kam69a, ON90, Rin71, TS76, Tou79, UYSA89, Wat76, CDS88]. Linearized [BG87]. Linkages [Pic69]. LINZ [WKL88]. Liquid [ZRS87]. Liquid-Filled [ZRS87]. Liska [GL89, LD90, Lis84, LD87]. LISP [FF82a, FF82b, GH79, GK79, HGB82, Ina80, KS81a, LM73, Lux75, MHGG79, MHGG80, MH85, MN89a, Nam86, NM88, NM89b, SGM87]. LISP- [NM89b]. LISP-Arbitrary [NM88]. LISP-based [KS81a]. LISP-Programme [LM73]. Lithography [GS78]. Liverpool [HWH89]. Llano [ANGK^{+}87, BBF^{+}82]. Lloyd [LP90]. ln [Köl85b, Köl86]. Local [FK87, GS90, PLR85]. Loe [LOG85]. log [Köl82, Köl83b, Köl83a]. Logarithmic [GP74]. Logarithms [Köl85a]. Logiciel [Tou87]. Lond [CGGK86]. London [Bir77, Fit89a, GvH89, GLB75, LM74, Mac86a, MN89b, NM89b, Woo89]. loop [BS98, dSS85, Kah69, Stu88, SGT90]. Loops [Bro85, ST89b]. Loos [CvH83, HL73, Loo72, vHC83]. Lorentz [SH81]. Lorentz-Covariant [SH81]. Lösung [Kra82]. Lottati [LExx]. Louis [Rao85]. Louw [LSS86]. Lovas [GL83]. Lovelock [DCM^{+}89]. Lovelock-type [DCM^{+}89]. Low [Bro70, CCH70, JO84, SH72, Ger80b]. Low-Pass [SH72]. Low-Reynolds-Number [JO84]. Low-T [CCH70]. LSI [GS78]. Ltd [NM89a, QGxx, QG84]. Luciana [BC81]. Luegger [LM73]. Lukács [LPSV82, PLSV84]. Lúkaszuk [LSS87]. Luni [Kam78]. Luni-Solar [Kam78]. Luongo [PLR85]. Lux [Lux75].

M [Eas93, MF83]. m.b.v [vH88]. M.H.D. [KPT71]. MacCallum [Eas93, Mac86b, Mac86a, Mac87, Mac88, Mac89a]. MacDonald [Art95]. Machine [Hea80, Lux75, Per88]. Mack [Mac73, MM73]. Mackenzie [LMSZ83]. Maclaren [Mac89b]. Macromolecules [KM80]. MACSYMA [Fit73, Tha89a, Tha89b]. Madison [Wat85]. Magnetic [ABDK69, BT63, BS67, Cal72a, Car73, DLS88, KC72, Lie75, NMJ84b, NMJ84a, Par68, Som77, Van76]. Magneto [MHK86]. Magneto-Plasma [MHK86]. Magnetofluid [DL88a]. Magnetohydrodynamic [BS86]. Magnetohydrodynamics [SH81]. Maguire [Mag81]. Maillard [DM85]. Majorana [Gar84]. Majorana-neutrino-mass [Gar84]. Makoto [MM89, NS82, NS85]. Malcolm [Mac89a]. Malm [Mal82]. Man [Per88]. Manifolds [FGPF88]. Manipulatie [vH88]. Manipulating [Kat85]. Manipulation [DG75, FK86, FB89, Hea74a, Hea78, Kan75, NA79, RT85, Ste82, Sto77a, Tao90, Wan85, Wat85, WD77, YP89]. Manipulations [Wat76, Wat79]. Mantle [Say87b]. Manual [Gat85c, Hea87, SvHH81]. Map [KM84]. Mappings [zZlH84]. Maps [Ken89b]. Margaritis [TSM86]. Marichev [AM90]. Mark [LS80]. Markovski [BCM^{+}95]. Marseille [Per78]. Marti [Mar78, MHGG79, MHGG80, MF83, MH85, Mar85, Mar88]. Martin [CG79, Gri78, Gri75, Gri76b, Gri76c, Gri77b, GK78, GH79, GK79, Sag88]. Martini [KM84]. Mary [Mac86a]. Masaaki [HI89, Ito94, Ito88]. Mass [BHHS87, Car73, Car75, SM83, Gar84].

Mass-Operator [Car75]. Massive [Van79]. Massless [BMH87]. Mat [BCDS87, HG89]. Matching [CG79, Hea71b, McI85]. matemãtico [RdS88]. Materials [Say87a]. Math [AD88, Alf82, BG87, Bir77, Car80, GV86, Her83, Hie84a, Hie85, Hie87b, Hie87c, Hie87d, Hie88, JO84, Kea85, Kee83, Kee85, Ker83, Ker86, KE86a, KE86b, Köl82, Köl83b, Köl83a, Köl86, KKP^{+}82, LM85, MMK81, NZ82, Pie84, SS84, Woo89]. Mathematical [ABD89, BDS88, DR86b, GM84, LM74, Wat76, LS79]. Mathematics [CFW87, Ede81, FBC86, Fla86, GTS80, HvH77, HvHS81, Ste82, vH81, vH89, Wat79]. Mathématique [Cap84]. Mathématiques [Dic85]. Maths [LS81, Mur85]. Matrices [AW88, Hea71a, Par71, IKRT89]. Matrix [BH88, Cap86b, FGM73, GKRR85a, GKR85, GKRR85b, HT87, IK87, Squxx, YP89]. Matsui [NM89a]. Matsumoto [MM89]. Matter [ANGK^{+}87, BCTC75]. Matveev [DMMS86, MDMS87]. Maurer [MHK86]. Max [TS76]. Max-Plack-Institut [DD85]. Max-Planck-Institut [KPT71]. Mays [GM84]. Mazepa [MS85]. Mazzarella [Maz85]. McClung [NMJ84b, NMJ84a]. McCoy [KM80, SM83]. McCrea [BGHM88, McC81, McC82, McC83, McC84a, McC87, MBG87, MMH88, McC84b, SHM87]. McIsaac [McI85]. MD [GK78]. Mean [BZZ86]. Means [BT63, GS90, GM84]. Measurement [DG75, Ear70, PN73b, WD77]. Measurements [CGGK86]. Mech [BT78, Feu84, Kee90, Rao85, ZRS87]. Mechanical [Squxx]. Mechanics [BBA73, BC81, CLS76, CB76, Ede82, Eit73, EH87, Kat85, Lev71, MS84, NM89a, NA79]. media [Kee89]. Melenk [LM73, MMN88, MMN89, MN89a, MN89b, NM88, NM89b, MN89c]. Meleshko [GMS90]. Membrane [Rao85, ZRS87]. Mendes [EM82]. Menna [MMR87]. Menzel [FGM73]. Merebashvili [DMMS86, DMR88, MDMS87]. Meshcheryakov [GMS86]. META [Mar78]. META/REDUCE [Mar78]. Metamorphosis [HGDR84]. Method [BGH88, CFW87, DT84, DDDD85, FH87, Kam69a, LV84, LD90, MHK86, NMJ84b, SH81, ZNR78]. Methods [Boy78, CCF84, EH87, Hie87a, KPB85, Lis84, Nik87, SvH82, Tro89]. Método [dS87]. Metric [BH84a, BGHM88, Per84a]. Metrics [Sch88a]. MHD [KG75]. MI [Car73]. Miano [MMR87]. Michael [GM85, GM88, Kar85, Möl89, WD90]. Michail [GS90]. Michigan [Boy78, FBC86]. Micro [GK78, GK79]. Micro-coded [GK78]. Micro-programmed [GK79]. Microprocessor [Fit83]. Microscopic [BNdP^{+}87]. Microwave [SvH82]. Mielke [BHM86, BMH87, MMH88]. Milgram [LM85]. Military [Wanxx]. Mills [KM84]. Milne [BDM^{+}87]. Minato-Ku [NM89a]. Minima [FC84]. Minimum [Kah69]. Minor [Gri78, Gri76a, Gri76c, Gri77b, Gri77a, Smi79]. Mirie [MS84]. Mitra [SM72]. Mitsuyuki [ON90]. Mittelbreitenformeln [Kra82]. Mitter [MM73]. Mixed [WD77]. Miyashita [KMS86]. mKdV [Hie87c]. mKdV-type [Hie87c]. MO [Rao85]. Mod [Tsa74]. Mode [Abi83, CH85, Hea74a, PH81, PH83]. Model [BH84b, BMH87, BCTC75, DM85, dSS85, GKRR85a, GKR85, GKRR85b, Hea76a, Kee85, Ker86, KE86a, KE86b, KT85, TSM86, ZY75]. Model-Independent [TSM86]. Modele [Lux75]. Modelling [BDM^{+}87]. Modelos [dS87]. Models [CDS88, CS77, Gre84]. modern [Ogi89]. Modes [HG84]. modular [Hea72b, Hea79]. Moiré [IY77]. Mol [Han87, GS77]. Molecular [PSA84]. Molecules [HG84, PSA84]. Möller [GM85, GM88, MMN88, MMN89, Möl89]. Moment [BT63, BS67, Cal72a, Car73, KK85, KC72, Par68, WTV87]. Moments [ABDK69, Bro82, KM80, SK83, Van76, WTV87]. Momentum [ADB^{+}84a, Bro72, BR72, BRS73, Gro88b, ZY75]. monde [Cap89]. Monogenic [BCRS89]. Monopole [KKP^{+}82]. Moon [Kam69b]. Moore [NM77]. Morel [BLMN78]. Morio [AM90, Bra90, Fit90, GS90, GMS90, Gat90b, GZ90b, GKZ90, LD90, NS82, NS85, Nor90, PB90, Tao90]. Moritsugu [KMH88, MIG85, MG88, MG89, MM89]. Morten [LS80]. Moscow [EKR89, KRxx, KS88, RT88a]. Mosel [BNdP^{+}87]. Moshe [Eis90]. Moskovskogo [EKR89]. Motion [Kam69b, Pic69, WTVNF87]. Moussiaux [DCM^{+}89]. Movable [Wat79]. Moving [GS77, KM80]. mu [Cun75]. mu- [Cun75]. Mulder [SGM87]. Müller [MMK81]. Müller-Kirsten [MMK81]. Multi [GP82, IC90]. Multi-Span [GP82]. Multi-Spin-Flip [IC90]. Multidimensional [DR85, dRD88]. Multiloop [ST89b]. Multiphoton [LMSZ83]. Multiple [Mac89b]. Multiplication [ABD89]. Multipole [FC84, RN86]. Multivariable [UYSA89]. Multivariate [Möl89, vH83]. München [EH86]. Muon [ABDK69, BS67, Ear70, Garxx, Gar80]. Murzin [Mur85]. Musser [HM75, LM74]. Muta [KKM81]. Mystery [DR89].

N [Art95, KR82, LS79, LS81, zZlH84, GDRH85, LS80]. n-Bubble [LS79, LS81, LS80]. Nachr [DJRR81]. Nadrchal [CNxx]. Nagata [AM90, Bra90, Fit90, GS90, GMS90, Gat90b, GZ90b, GKZ90, LD90, NS82, NS85, Nor90, PB90, Tao90]. Nagatomo [ON90]. Nahm [LSS86]. Nakamura [NM89a]. Nakashima [NMJ84b, NMJ84a]. Nakel [PN73b, PN73a]. Namba [Nam86]. Naoto [NK84]. National [DT81, Fla86, KKM81]. Navelet [BLMN78]. Near [Kah69, Kam69b, McC81].

Near-Minimum-Time [Kah69]. Nearly [JO84]. Necessary [GZ90b, GKZ90]. Nechaevsky [BGKN89]. Nedyalkov [ZNR78]. Neil [eS85]. Neilson [RN86]. Neito-Frausto [WTVNF87]. Németh [NZ82, NZ87]. Netform [HvHS81, SvHH81]. Netherlands [ESP88, Gat85c, Gat85b, HvH77, HvHS81, vdH86b, vHvH88, vH81, vH88, vH89]. Neun [MMN88, MMN89, MN89a, MN89b, NM88, NM89b, MN89c]. Neurons [Hor83]. Neutrino [BCTC75, GGRxx, Gar84, Gar84]. neutrino-electron [Gar84]. Neutsch [NSJ85, NS86]. Newman [BGHM88, ER90]. Newton [MIG85]. Ng [Ng89]. Nichtnumerische [Mac73]. Nielsen [Mac89a]. Nihon [Kob84]. Niita [BNdP^{+}87]. NIKHEF-H [GGRxx]. Niki [NK84]. Nikityuk [Nik87]. Nikolai [GKZ90]. NN* [DT69]. Nobuyasu [IC90, Ito90]. Nobuyuki [Ina80]. Nombres [Duv87]. Non [GV86, GSZ85, Hea72b, Hea79, Kam69a, KG75, KRxx, LH70, Wat79, CDS88, KR88, Rod84]. Non-Abelian [KRxx, KR88]. Non-circular [KG75]. Non-Classical [GV86]. non-commutative [Rod84]. Non-existence [Wat79]. Non-Linear [GSZ85, Kam69a, CDS88]. Non-modular [Hea72b, Hea79]. Non-Symmetric [LH70]. Nonconservative [EH87, EC87]. Nondiagonal [dRD88]. Nonlinear [Ada83, DG75, FP85, GZ90b, GKZ90, GSSZ87, Gro88a, Gro88c, Her83, Ito94, Ker83, MHK86, Ren92, Rin71, RA89, Sch84, WD77, Gro88b]. Nonmetricity [BHM86]. Nonoyama [ANK81]. Noor [NA79]. Noritaka [LOG85]. Norm [PFF83]. Norm-Conserving [PFF83]. Normal [McC81, Ren92, WD84]. Normal-Dominated [McC81]. normalization [BCM^{+}95]. Norman [HN79, Nor90, NM77, Nor78, ND79, NW83]. North [KK89, Köl81, ST89b].

North-Holland [BHM86, ESP88]. Norton [Nor80]. Note [EM82, MG89, MM89, NSJ85, Nor80, PA86, Sch88a]. Notes [Dav81]. Notices [GLB75, HN79, Mar78, MHGG79]. Notorious [vH82]. November [ESP88]. Novosibirsk [Gro88a, Gro88b, Gro88c]. nubar [Cun75]. Nucl [Ada83, Cam67]. Nuclear [AE87b, BMH87, BNdP^{+}87, FT75, GKRR85a, GKR85, GKRR85b, GVxx, GV78, Gro88a, Gro88b, Gro88c, HG84, KC72, KC73, KRxx, RT88a, Say87a, FGGS87]. Nucleon [BHP62, Fle71, FGM73, FT75, Garxx]. Nucleon-Nucleon [Fle71, FGM73, FT75]. Nucleus [BNdP^{+}87]. Nucleus-Nucleus [BNdP^{+}87]. Number [ABD86a, Abb88, DDDD85, JO84, KE86b]. Numbers [Abb89, Bro85, Dic85, Mac89b]. Numeric [SvH82, vH87]. Numerical [ANK81, Dew89, Gat86, Mac86b, MM89, PFF83, vH89]. Nuovo [Gar80, HG84, MBG87]. NUT [BH84a, McC84a]. NUT-Like [McC84a].

O [dS87, Adk83]. Objective [GS77]. Observer [Dau83]. Obtained [SH72]. O’C [DD85]. Occurring [AB89]. Occurs [EH87]. Oceanic [Boy78, FBC86]. Ochiai [ON90]. Odd [Maz85, vH89]. ODEs [KPB85]. off [Dav82a]. Ogilvie [Ogi82, Ogi89]. Ohlsen [KO83]. Ohsawa [LOG85]. Oklahoma [LS79, LS80]. One [EH87, FT75, HN79, Stu88, SGT90, zZlH84]. One-Boson [FT75]. One-Dimensional [zZlH84]. one-loop [Stu88, SGT90]. One-Pass [HN79]. Onishi [JO84]. Ono [Ono79]. Open [CGGK86, Kah69]. Operational [EH86]. Operations [Ito88, MN89b, MN89c]. Operator [AE86, AE87b, Car73, Car75, GKRR85a, Ito88, Tro89]. Operators [CH85, Fla86, GS90, Han87]. Opt [CBR84, FP85]. Optik [GS77, Som77]. Optimal [Hea86]. Optimisation [PH81].

Optimization

[HvHS81, vH83, vHHGV89, WCvH84].

Optimized [GvH85]. Optimizer [SvHH81]. O’Raifeartaigh [dS87]. Orbit [AB89]. Order [ABDK69, AB70, ADB^{+}84a, ADB^{+}84b, BGKN89, Cal72a, Fox71, FH74, KK85, Kau73, KC72, NK84, Par68, SH72, SL74]. Ordinary [GK89, Hie85, Mac88, Sch82b, Sch83a, Sch85b, Sch86, Wat76]. Ortho [Kau73]. Ortho-Positronium [Kau73]. orthogonal [Eas91, Hap89, RT89]. Orthogonalization [BOH78]. Orthonormalization [YP89]. Orthopositronium [Adk83, Adk85, SH74]. ORTHOVEC [Eas91]. Osami [UYSA89]. Oscillating [CCF84]. Oscillation [AZZ87]. Oscillations [Kam69a]. Oscillator [BPV78, Gro88a, Gro88c, HT87]. Oscillatory [CGGK86, Kee83, Kee85]. Ososkov [BOH78]. Outer [Har87]. Outputs [Sht77]. Overall [PLR85]. Oxford [Eas93]. Oxidation [CGGK86]. OZI [BLMN78]. Ozieblo [Ozixx].

p [Abb89]. p-Adic [Abb89]. P.D.E.’s [GMS90]. Pacheco [YP88, YP89]. Package [ABD86a, Bam88, BH88, Cap86b, Cap86a, DJ80, Har97, HvH85, IK87, Ito94, Kad96, KS84b, LD90, Ren92, Sas79, SHM87, Sch82b, Sch83b, Sch86, Sei91, vHH82, vHHGV89]. Packed [SM83]. Packed-Bed [SM83]. Packet [RT88a]. Padé [FGM73, NZ82, NZ87]. Padget [BDM^{+}87, BHPS86, DP85, PB90]. pages [Art95]. Painlevé [Wat79, Ren92]. Pair [BT63, Ito85, Pes73, PSA84, Tsa74]. Pair/Completion [Nor90]. Pairs [Ear70]. Palermo [BCDS87]. Pankau [PN73b, PN73a]. Paoli [BNdP^{+}87]. Papadopoulos [DCM^{+}89]. Paper [AD88, Mac89a]. paperback [Art95]. para [dS87]. parallel [Fit89b, MN89b, MN89c]. Parallelism [Fit89a, GvH89, MN89b, NM89b, Wan85]. parallelization [Bra90]. Parameter [EH87]. Parameters [HH80, JPS87, Wan85, WTV87]. Parametric [KC73]. Paraquarkonium [KKM81]. Parisi [BPV78]. Park [KKP^{+}82]. Parker [JPS87]. Parsons [BHP62, Par68, Par71]. Part [Boy78, BHP62, Bro82, Tha89a, Tha89b, MS84, TSM86]. Partial [Fle71, FGM73, FC85, Her83, IK85, Sch82b, Sch85a]. Particle [AZZ87, Cam74, Par71, Van76]. Particles [BW90, Kad88, KR82, KS88]. Partition [KR82]. Parton [GBB72, FGK89]. Partons [BGJ72]. Pasadena [Kau73]. Pass [HN79, SH72]. passing [Hie87b, Hie87c, Hie87d, Hie88]. Past [Dau83]. Patrick [ST89a]. Pattern [CG79, Hea71b, KRR87, McI85]. Pattnaik [PFF83]. Paul [JW86, NW83, WCvH84]. Paulo [GV86]. PC [HI89]. pDE [LD90]. Pearce [PH81, PH83]. Pearson [LP90]. Pedersen [Mac89a]. Peltier [BBF^{+}82]. Penrose [ER90]. performance [Mac89a]. Period [zZlH84]. Period-n-Tupling [zZlH84]. Periodicity [CGGK86]. Perjes [Per88, LPSV82, Per84b, PLSV84, Per84a, Per84c, Per86a, Per86b, Per88]. Perlt [PRH90]. Perrottet [Per78]. Personal [FB89, Hea80, NS85, YA87]. Perturbation [ANGK^{+}87, CB76, Kam69a, Kam69b]. Perturbations [CS77]. Perturbed [Kam69b]. Pesic [Pes73]. Peter [BH84b, Gra81, HH80, JW86, WD90]. Petrolito [GP82]. Petrov [McC81]. Pfirsch [KPT71]. Ph [DCM^{+}89]. Phase [FG84, Gre84, Sch77]. Phenomena [KMS86]. Phenomenological [DT69]. Photon [ABDK69, ADB^{+}84b, BNdP^{+}87, Pes73]. Photon-Photon [ABDK69]. Photons [AB83, BCTC75, Geo68]. Photoproduction [ADB^{+}84a, CCH70]. Phys [Adk83, AB83, Adk85, ANGK^{+}87, ABDK69, AB70, AH73, ADB^{+}84a, ADB^{+}84b, BH84a, BHHS87, BG87, BMH87, BGHM88, BBF^{+}82, BBA73, BS86, BKdCW81, BT63, BBR85, BCTC75, BH88, Bro85, BHP62, BS67, BGJ72, BR72, BRS73, Bro82, BH79, Cal72b, Cam67, Cam68, CH70, CCH70, Car80, Car75, CCF84, CCEL84, Cun75, DMMS86, Dau83, DCM^{+}89, MMR87, DR85, DM85, DGH^{+}86, DT69, Ear70, EM82, EFK85, FGK89, FC84, FK86, Fle71, FGM73, Fox71, FH74, FC85, GvPV79, Geo68, Ger80a, GTS80, GP74, GDRH85, Gre84, GK80, GBB72, GK85, HT87, Han87, Hap89, HH80, HKY69, Her83, Hie83b, Hie83a, Hie84a, Hie84b, HGDR84, Hie85, Hie87b, Hie87c, Hie87d, Hie88, HG89, Ito85, IK85, Ito88, JPS87, JO84, KK89, Ker83, KM84, Ker86, KE86a, KE86b, KMS86, Köl81, KR82, KS84a, KS84b, KKP^{+}82, KR88, LV84, LS81, LM85, LMSZ83, Lie75, Lis84, LSS86, LSS87, MM73, MDMS87, MHK86, McC81, McC82, McC83, McC84a, MMH88, MMK81, NSJ85, PN73b, Par68, PFF83, PLSV84, PRH90, PA86, PSA84, RN86, SH81, Sch82c]. Phys [SS84, SKxx, SH74, Stu88, SGT90, ST89b, Tal84, TH65, Tsa74, WTV87, WTVNF87, YP88, YP89, ZY75, zZlH84]. Physica [Cam74, IC90, Kee89, Sch84]. Physical [FGGS87, Gar84, Gro88a, Gro88b, Gro88c, IY77, Ina80, Ogi82, TSM86, Wan85]. physicists [Art95]. Physics [AE87b, Art95, BH84b, BS98, BCM^{+}95, Bil92, BZZ86, BNdP^{+}87, Boc92, Bro72, Cal72a, Cam74, CNxx, CS77, DR86a, DLS88, DL86, dSS85, Eas91, Eas93, ER90, FT75, FF82b, FG84, Ger80a, GTS80, GMS86, Gro88a, Gro88b, Gro88c, Har97, Hea72a, Hea81, Hie87a, IK96, Ito94, Ito90, Kad96, Köl84, KRxx, LPSV82, Ogi89, Ono79, Per84a, Per86a, Ren92, RT88a, Sch82b, Sch86, Sei91, Sht77, WW86, GL83, GZ84]. Physik [Kad88, LP79, Mac73, PN73a, Sch80]. Physique [Cap84]. Pictiaw [Pic69].

Piessens [Pie84, PA86]. Pignataro [PLR85]. Pim [vdH86a, vdH86b]. Pinto [FP85]. Pion [CCH70]. Pions [Geo68]. Piotr [FBC86]. Pipe [Kot86]. Planar [TC88]. Planck [TS76, Bro82, Hie85]. Plane [BT78, LV84, McC83]. Plasma [Bro82, MMR87, DLS88, DL88b, LD87, MHK86]. Plasmaphysik [KPT71, TS76]. Plasmas [KPT71]. Plastino [BBF^{+}82]. Plates [Eis90, GP82, LExx]. Podgórak [PR84]. Poincare [Vor77, McC84a, McC87, BH84a, BHHS87, BG87, BMH87, BGHM88, BGH88, McC83, MBG87, MMH88, McC84b, SHM87]. Point [Alf82, EFK85]. Points [CCF84, DT81, GV86, Kam69b, Kob84, PA86, Wat79]. Poiseuille [BT78]. Poland [Ozixx, PR84]. Polarization [AE86, AE87b, GG78]. Polarized [Garxx, Gar80]. Polycrystalline [Say87a]. polymer [WD90]. Polynomial [AD87, BCM^{+}95, Gat90b, GZ90b, GKZ90, Gri78, Hea72b, HL73, Hea74b, Hea79, Hie83b, MN89b, NZ82, MN89c, RN86, RT88b, Sch85b, vH83]. Polynomial-Nonlinear [GZ90b, GKZ90]. Polynomials [Abb88, ABD89, BOH78, Bil92, BCRS89, BD88, MG89, Sht75]. Pomeau [GP74]. Ponce [FGPF88, FGP89]. Poon [KPB85]. Popovic [ANGK^{+}87]. Porod [LP79]. Portable [GH79, HGB82, MN89a]. Positive [KS81b, Pie84]. Positron [Cam67, HKY69]. Positronium [AB83, Kau73, LMSZ83]. Postbuckling [PLR85, RT85]. Potencial [dS87]. Potential [CS77, dSS85, Hie83b, NA79]. potentials [DGH^{+}86, GVxx, GV78]. pour [Lux75]. Power [FK86, Kea85, PB90]. Powers [Köl85a]. Practical [CCF84]. Practice [GH79]. Pragmatic [vH83]. Prague [CNxx]. Precision [Bro70, NM88, NM89b, Ono79, Sas79]. Preconditioning [MG89]. predators [Kee85]. Predictions [CGGK86].

Preliminary [Hus81]. Preprocessor [KK89]. Presence [Sht75]. Presentation [Smi79]. Press [Abb89, DST88, Dew89, Eas93, Fit89b, Fit89a, GvH89, GK78, MN89b, NM89b, MN89c, RA89, Sch87, vHHGV89]. pressure [WW86]. Pretty [HvH83, vHvH88, vdHHvH87]. Pretty-Print [vdHHvH87]. Prettyprinter [HN79]. Price [Art95, PSA84]. Print [vHvH88, vdHHvH87]. Printing [HvH83]. Prob [Ken88]. Probability [Ken89b]. Problem [BH72, CB76, Dau83, Har78, Hea72c, JW86, MS85, Nor78, vH80, WR79]. Problem-Solving [JW86]. Problems [BS87, DD85, Ede82, Eit73, Fit73, Fit85b, GK80, Gro88a, Gro88b, Gro88c, Hea71b, Hea72a, Kea85, Köl84, MMN89, RA89, vH80, VR87, Wanxx]. Proc [CGGK86, DL86, DL88b, Woo89]. Procedures [FK86, ST89b]. proceedings [ESP88]. Process [BKdCW81, KS88, PN73b]. Processes [DMR88, DR89, Lie75]. Processing [HW81]. Processor [MN89a, NM88, NM89b]. Produced [KK89]. Producing [Sht77, vH80]. Production [ADB^{+}84b, BT63, BNdP^{+}87, BCTC75, Ear70, FGK89, KS88, Per78, Pes73, Tsa74]. Products [Hea71a]. Proeyan [Van79]. Proeyen [GvPV79, Van76]. Professional [BJ87]. Prog [BHHS87]. Program [BOH78, BCM^{+}95, Bil92, Boc92, CNxx, DJ80, DR85, Eas91, EKR89, EFK85, GLB75, Hap89, Har77, Har79b, Hea73, Hea74a, IK96, Ito85, IK85, Ito88, Köl81, KS84b, KF87, KR88, Lis84, LM74, Mag81, Mal82, Sch77, vH87, WR79, YP88]. Programme [Cap84, LM73]. programmed [GK79]. Programming [KRR88]. Programs [Ede81, FB89, GvH85, KR85a, KR87, PH81, PH83, SL74, Tha89a, Tha89b]. Progress [Ito90, NA79]. Projection [IY77]. Projection-Type [IY77]. Prolongation [Gra81]. Prompt [ADB^{+}84b]. Proof [LPSV82, Per86b]. Propagation [FP85]. Properties [Her83, Mur85, ST89a]. Property [AD88, zZlH84]. Proposal [MG88, GH82]. Proton [BNdP^{+}87]. Proton-Nucleus [BNdP^{+}87]. prover [ST89a]. Provisional [FF82a]. Prüfung [Kra73]. Pseudopotentials [PFF83]. Pseudorandom [Mac89b]. Pseudoscalars [BLMN78]. Publications [Wat76]. Publishing [Art95, KK89, Köl81, ST89b]. Puiseux [GKK87]. Pulse [NMJ84b]. putting [ESP88].

QCD [ADB^{+}84a, Dar86, DMMS86, Dul87, HH80, MDMS87]. QD [Sch83b]. QED [FG84, PRH90]. Qingsheng [Tao90]. Quadratic [McC83, McC84a, McC84b]. Quadratic-Poincaré [McC84a]. Quadrature [Tao90]. Quadrupole [Van76, GKRR85a]. Quadtree [AW88]. Quantenelektrodynamik [Kra73]. Quantities [KE86b, WTV87]. Quantum [BSW88, BGH88, Bro69, Bro70, Bro71, Bro72, BR72, BRS73, FH74, Hie84a, HG89, IK87, LS79, LS80, RT87b, Sag88, ST89b, Van76, Vor77, Gro88c]. Quartic [BPV78, DGH^{+}86]. Quarton [GQ86, QGxx, QG84]. Quasi [MHK86]. Quasi-Analytical [MHK86]. Queen [Mac86a]. Quentin [Mag81]. questions [Duv87, Nik87]. Quirais [dS87].

Rabbe [FF82a]. Radial [BBR85, HT87]. Radiative [BLMN78, Cam67, FC85, HKY69, Kau73, KC72, SH74, TSM86, Gro88c]. Radiative-Recombination [FC85]. Ralph [DR86b]. Ramani [DGH^{+}86, GDRH85, HGDR84]. Ramos [ER90]. RAND [Hea87]. Random [Gre84]. Ranft [PRH90]. Range [EH87]. Rank [IK85]. Rao [Rao85, ZRS87]. Rapid [CG79]. Rate [Adk83, AB83, Adk85, SH74].

Rational [Hea74b, HvH77, Kar85, Möl89]. ray [WTVNF87]. Raychev [GKRR85a, GKR85, GKRR85b]. Rayna [Ray87, WR79]. reactions [WD90]. Real [BHP62, Ono79, Sas79]. Realisation [Tou87]. realization [AM90, KO83]. Realizations [DMR88]. Really [KO83]. Recalculation [Fox71]. Recherche [DT81]. Rechnerunterstützte [Kra82]. Recognition [Cal74, GvH89]. Recombination [FC85]. Reconstruction [Möl89]. Recovery [Abb89]. Recurrence [HT87, PR84, RT88b]. Recursive [Gri77a, Smi79]. Red [AE87b]. Reduce [WD84, Abi83, AM90, ASW89, BS98, BOH78, Bam88, BCM^{+}95, Bil92, Boc92, BH88, BC81, BCDS87, BHPS86, Cal72b, Cap84, CH85, Cap86b, Cap86a, Cap89, CNxx, CC88, CB76, DMR88, Dau79, DJ80, DJRR81, Dav82a, DR85, Vos89, dS89, dS88, Dul87, Eas91, Ede81, EKR89, EFK85, FK86, Fit90, Fit73, Fit83, Fit85b, Fit89b, FBC86, GS90, GMS90, GL89, Gat85a, Gat85c, Gat85b, Gat86, GK80, Gri74, Gri75, Gri76b, GK78, GK79, Gro88a, Hap89, Har79a, Har77, Har78, Har79b, Har97, Hea72c, Hea73, Hea76a, Hea82, Hea87, HI89, HvH83, HvH85, Hus81, IK96, Ina80, Ito94, Ito85, IK85, Ito88, Kad96, Kan75, KK89, Kaz87, KF87, KO83, KRxx, KR85a, KR85b, KRR87, KR87, KRR88, eS85, LD90, LD87, LP90, Mac88, Mag81, Mal82, Mar78, MF83, MH85, Mar88, Maz85, MMN88, MG88, NM89a, Ng89, Nor78, NW83, ON90, Ozixx, PB90, PH81, PH83, PR84, Ray87, McC84b, Ren92, Rod84, RT89, Sas79, Sch82a, SHM87, Sch82b, Sch83b, Sch86, Sei91, Sha87, Som85, Spi87, ST89b, Tha89a, Tha89b]. REDUCE [GH82, vdH86a, vdH86b, vdHHvH87, vHvH88, vHH82, vH88, vHHGV89, VR87, Wan85, WR79, WH89, YA87, Eas93, Art95]. REDUCE-2 [Rod84, BOH78, CNxx, Dul87, Ede81, Hus81]. REDUCE-3 [DMR88]. REDUCE-based [Boc92]. REDUCE-Package [SHM87].

REDUCE. [Gro88a, Gro88b, Gro88c]. REDUCE/1700 [GK78]. Reducible [BGKN89]. Reduction [Stu88, SGT90]. Reentrant [KMS86]. REFAL [Mur85]. Refined [LExx]. Region [AE86, AE87b, BGJ72, LP79]. Regularization [Bir77, IK87, LM85]. Regularized [GvPV79]. Reguliers [DT81]. Rel [dRD88]. Related [CS77, vdH86b]. Relation [Bro70, RT88b]. Relations [HT87, PR84, Sht75]. Relationship [Cam74]. Relative [Tal84]. relatives [Duv87]. Relativistic [BZZ86, Gla82, Som77]. Relativity [Car80, CLS76, FC79, CFÅ84, Dau79, DJ80, DJRR81, Mac86b, Mac86a, Mac87, Mac89a, Ozixx, Per84b, PLSV84, Per86b, McC84b, SHM87]. Remark [ABD85, AD88, ABD89, MMH88]. Remarks [Fla86]. Renard [GGRxx]. Renato [dS89, dS88, RdS88]. Rend [BCDS87]. Renner [Ren92]. Renormalizable [BCTC75]. Renormalization [BR72, BRS73, Cap90]. Report [MHGG79, MHGG80, Sch82a]. Reports [Hie87a]. Representation [AW88, BC90]. Representations [DR86b, Hea74b]. Requirements [Sch80]. Res [Kag88]. Research [FGGS87, GKRR85a, GKR85, GKRR85b, Gla82, GL83, GZ84, IY77, Ina80, Kob84, LS79, LS80, LPSV82, Per84a, Per86a, RT88a, Wat76, FGGS87]. Resonance [AZZ87, BHP62, DL88a, NMJ84b, NMJ84a]. Resonant [CBR84, WTVNF87]. Results [GKRR85b, KMS86]. Reusch [RN86]. Rev [Adk83, AB83, Adk85, ANGK^{+}87, ABDK69, AB70, BBF^{+}82, BT63, BBR85, BCTC75, BHP62, BS67, BGJ72, BRS73, CCH70, Car75, DCM^{+}89, DT69, Ear70, FC84, FGM73, Fox71, FC85, Geo68, GP74, HKY69, Hie83a, Hie84b, HGDR84, LV84, LMSZ83, LSS87, Par68, PFF83, SH74, TH65, Tsa74, ZY75]. Review [Art95, Eas93, Gar84, Sch88b, TSM86].

Reynolds [JO84]. Ricci [Kad96]. Richard [CG79]. Richardson [KPB85]. Riemannian [DR85]. Riemer [DJRR81]. Rigorous [KMS86]. RIKEN [Som85]. Rink [Rin71]. Rio [dS88]. Riquier [Sch84]. Riquier-Janet [Sch84]. RISC [WKL88]. Risch [Abb87, NM77]. Rizzi [PLR85, RT85]. RLISP [Kry84]. Robert [GK78, GK79]. Robin [Stu88, SGT90]. Rodioniov [IKRT89]. Rodionov [DR85, DR86a, EKR89, KRxx, KR85a, KR85b, KRR87, KR87, KRR88, KR88, Rod84, RT87b, RT87a, RT88a]. Rogers [RA89]. Role [KS88, Mar85]. Romanowska [PR84]. Ronveaux [BCRS89, RT88b, RT89]. Rop [dRD88]. Roque [RdS88]. Rosenbrock [KPB85]. Roskies [BR72, BRS73, DR86b]. Rostovtsev [BCM^{+}95, DMR88, DR89, KRR87, KRR88, VR87, ZNR78]. Rotating [DLS88, vH80]. Rotation [Han87]. Rotationsellipsoid [Kra82]. Rotator [Gro88c]. Roussev [GKRR85a, GKRR85b]. Routine [BW90]. Routines [JW86]. Roux [BBR85]. Royal [CGGK86, CLS76, CB76]. RTENSOR [RT88a]. Rubin [DR86a]. Rubinacci [MMR87]. Rüdiger [GM85, GM88, HL73, Loo72]. Ruffini [BHM86]. Rühl [KR82]. Rule [BLMN78, Mag81]. Rules [KC73, KRR88]. run [Fit89b]. Rushworth [CBR84]. Russell [Bra90]. Russian [GZ84, KS88]. Ryogo [HI89].

Saclay [GV78]. Sadanori [NK84]. Saddle [CCF84]. Saez [SM83]. Sage [Sag88]. Saitama [IY77]. Saito [UYSA89]. Salim [Abi83]. Salpeter [Fle71, FT75]. Sample [NK84]. Samuel [LS79, LS80, LS81]. Sandia [SL74]. Santos [dSS85, dS89, dS87, dS88, RdS88]. Sasaki [KS81a, Sas79]. Satellite [Kam78]. Satellites [Har87]. Saturation [JPS87]. Saturn [Har87]. Sauro [BDS88]. Savage [Sav90]. Sayers [Say87a, Say87b]. Scalar [BMH87, BCTC75, Stu88, SGT90, FG84]. Scalar-Boson-Exchange [BCTC75]. Scaling [BGJ72, zZlH84]. Scattering [ABDK69, Cam67, FGM73, FT75, FP85, Garxx, Gar80, Gar84, GBB72, GBB73, HKY69]. Scheme [Sch80]. Schemes [GL89, Sch83b, vH83]. Schiff [ADB^{+}84a, ADB^{+}84b]. Schmuck [Sch77]. School [ABD89, BDS88]. Schorr [KS84a, KS84b]. Schrödinger [CS77, MMK81]. Schruefer [CDS88]. Schrüfer [BHPS86, NSJ85, NS86, SH81, Sch82a, SHM87, Sch88a]. Schwarz [DGH^{+}86, KS81b, LSS86, Sch80, Sch82c, Sch82b, Sch83b, Sch83a, Sch84, SS84, Sch85a, Sch85b, Sch86, Sch87, Sch88b]. Sci [Ada83, Ede82]. Science [Boy78, Dav81, FBC86, FB89, Gat85c, Gat85b, Hea76c, HI89, Ina80, Kah69, Kan75, Kob84, MMN89, MG88, NS86, RA89, Sch87, Sto75, vdH86b, vHvH88, vH88, vH89]. Sciences [ABD89, BDS88, GKRR85a, GKR85, GKRR85b, Per84a, Per86a, Wat76]. Scientific [DR86b, Hea76c, Hea85, Kat85, KT85, LOG85, Mar85, MIG85, Per88, RT88a, Som85]. Scientifique [DT81]. Scientist [Dav88]. scroll [Kee89]. search [Hie83b, Hie87a, Hie87b, Hie87c, Hie87d, Hie88]. Sebestyén [LPSV82, PLSV84]. Second [ADB^{+}84a, BGKN89, Hie87a]. Section [Cun75, FGK89, Kot86, KS88, PN73b, Pes73, TH65]. Sections [FC84, FC85, KG75, MM73]. Sedimentation [Feu84]. Seiler [Sei91]. Seismic [KK85]. Seitz [BSW88]. Self [BSW88, KM84, MHK86]. Self-Confined [MHK86]. Self-Consistently [BSW88]. Self-Dual [KM84]. Sellar [DLS88]. Semi [GVxx, GV78]. Semi-Classical [GVxx, GV78]. Semidiurnal [ANK81]. Semileptonic [TSM86]. Semilinear [Kea85]. Sensitivities [SH72]. Sensitivity [SM72, SL74]. Separation [LM85].

Sequences [Mac89b, zZlH84]. Sequencing [Mag81]. Ser [BCDS87]. Serdyukova [MS85]. series [CFW87, MMR87, FK86, GM84, KS81b, Mac89b, PB90, Pie84, Sch83b, TS76]. Serras [BS87, BCDS87, BCRS89]. Set [KKP^{+}82, WTV87]. Setsuya [KK89]. Settles [ND79]. sha [HI89, ON90]. Shabat [GSSZ87]. Shablygin [Sha87]. Shape [Ken89a]. Shapes [Ken88]. Shaskov [GS90]. Shear [Boy78, Rao85, ZRS87]. Shelest [GMS90]. shells [Boc92, LH70, LExx]. Shibayama [NS82, NS85]. Shift [AB70, Fox71]. Shigetoshi [Kat85]. Shirafuji [BHHS87]. Shirikov [FGGS87]. Shirkov [GTS80, GMS86]. Shmueli [SW83, SK83]. Short [Sht77]. Shouichi [NM89a]. Shtokhamer [Sht75, Sht77]. Shuichi [MG89, MM89]. Shunro [Wat76, Wat79]. Shuppan [NM89a]. Shuzo [ST89a]. Shvachka [GSZ85]. Siam [Sch88b, Kee83]. Side [Sht75]. Siemienczuk [LSS87]. Signal [Gre84]. Signature [Per78]. Significance [Wan85]. Sigplan [GLB75, HN79, Mar78, MHGG79]. SIGSAM [ASW89, BH72, Cal74, Fit73, Hea72b, Hea72c, Loo72, Sch88a, ABD85, Abi83, BC81, CH85, Cap86b, Cap86a, Dav82a, Dav82b, FF82a, Gat85a, Har78, Har79b, HvH83, IKRT89, KS81a, Kre88, KO83, Kry84, KR85a, KR85b, eS85, Mac89a, MHGG80, MF83, MH85, Maz85, McI85, MG89, MM89, Nam86, Nor78, NW83, Nor80, PH81, PH83, Rod84, RT88b, Sch82a, Sch83b, SvHH81, Ste82, GH82, vdH86a, vdHHvH87, vH80, vHH82, WR79]. Silver [GK78]. Similar [Loo72]. Similarity [Sch77]. Simple [DD85, NMJ84b, NS86, vdHHvH87]. Simplification [Cap86a, Hea76a, IK96, Maz85, RT87a]. Simulation [ANK81, LOG85, WD77]. sin [Köl83b]. Sine [Hie87d]. Sine-Gordon-type [Hie87d]. Singapore [Fla86, Hea85, Kat85, KT85, LOG85, Mar85, MIG85, Per88, Sch87, Som85]. Single [BKdCW81, GP82, Ito90, MMK81]. Single- [MMK81]. Single-Span [GP82]. Single-Spin-Flip [Ito90]. Singularities [EM82, LM85]. Singularity [McC81]. Singuliers [DT81]. Siret [DST88]. Sistema [dS88]. Six [Wat79, WH89]. Sixth [ABDK69, Cal72a, KC72, Par68]. SLAC [Bro70, Bro72, GBB73]. Slavnov [KS88]. Slavyanov [CS77]. Slender [BT78]. Slepchenko [DMMS86, MDMS87]. slow [KS81a, Rao85]. Small [Ken89b, KO83]. Smit [HvHS81, Smi79, SvHH81, SvH82, SGM87]. Smoluchowski [WTVNF87]. Smooth [CS77]. Soc [CGGK86, CBR84, FP85, KK85, Kea85, Say87b, Woo89]. Society [GK78, HW81]. Soderstrand [SH72, SM72, SL74]. Sofia [GKRR85a, GKR85, GKRR85b]. Software [Fit81, GH79, JW86]. Soiffer [eS85]. solar [Vos89, Kam78]. Solid [BDM^{+}87, CNxx]. Solids [Ede82]. Solitary [MS84]. soliton [Hie87b, Hie87c, Hie87d, Hie88]. Solution [BH84a, BDS88, EH87, FP85, Gat90b, Ger80b, Gri74, Gri76a, Har78, Hea71b, Hea72a, Hea72c, KKP^{+}82, McC84a, MBG87, MMH88, MMN89, Nor78, Per84b, PLSV84, Per84c, RT88b, SKxx, Sto75, Tou79, Wanxx]. Solutions [AH73, BH84b, BHHS87, BG87, BSW88, BBA73, DT81, MMR87, DR86a, Her83, Kea85, Mal82, McC83, McC87, McC84b, RN86, TS76, Tou87]. Solve [BGH88, Rin71]. Solved [BS87, WR79]. Solver [Mac88, Sha87]. Solving [Abb87, BGK86, Fit85b, GKZ90, Gro88a, Gro88b, Gro88c, JW86, KMH88, LD90, MHK86, NS86, Sto77b, Wat76]. Soma [GS77, GS78, Som77, Som85]. Some [BDS88, BGK86, DMR88, FH74, GKR85, Her83, HvH83, JW86, Kea85, KMS86, Köl84, MS85, Nam86, Nik87, QGxx, RT89, Sch87, Spi87, Squxx, vdH86b, vdHHvH87]. son [Lux75]. Sophus [Sch88b]. Sound [GQ86].

Source [BGHM88, KK89]. Sources

[Kag88]. Southern [HM75]. Sov [GTS80]. Space [BS87, DR85, DR86a, Kag88, KC73, LPSV82, NS86, Per84b, Per84c, Per86b, Wan85, DCM^{+}89]. Space-Times [DCM^{+}89]. Spaces [PLSV84]. Spacetime [BH84b, McC81, McC82]. Spacetimes [Mac86b]. Span [GP82]. SPARC [MN89a]. Sparse [ABD89, Gri78, Gri74, Gri76a, Gri76c, Gri77b]. Spatial [KM80, Pic69]. Spatially [dRD88, DLS88]. Special [Köl84, Rin71]. Specifications [HM75]. Spectra [CBR84, KG75, NMJ84a]. Spectral [DJ89]. Spectrum [WTVNF87]. Sphere [CBR84, WW86]. Spheres [GP74, JO84, Woo89]. Spherical [DJ89]. Spin [BHHS87, CB76, Garxx, Gar80, IC90, Ito90, KMS86, Van76, GVxx, GV78]. Spin- [Gar80]. Spin-1 [Van76]. Spin-dependent [GVxx]. Spin-independent [GV78]. Spin-Zero [Gar80]. Spinning [BGHM88]. Spinor [PRH90]. Spiridonova [Spi87]. Spring [GK78]. Springer-Verlag [AD87, Abb87, AW88, BDM^{+}87, BS87, BD88, CvH83, CDS88, CC88, CG79, DP85, DT84, Dul87, FGGS87, FK87, Fit83, FH87, FGPF88, FGP89, GL89, GSSZ87, GKK87, GK89, HvH84, HvH85, IK87, Kar85, Kaz87, KMH88, KR87, LD87, Mac87, Mac88, Mal82, Mar88, NZ87, ND79, Ray87, RT87a, Sas79, Sch83a, Sha87, Smi79, SvH82, SGM87, Spi87, TC88, Tou79, Tro89, vdHvHG87, vHC83, vH83, WCvH84, WH89, YA87]. Square [BBF^{+}82]. Square-Well [BBF^{+}82]. Squire [Squxx]. Srivastava [dSS85]. St [Rao85]. Stability [BT78, GL89, Kam69b, KPT71, Lis84, MS85, PFF83, QGxx]. Standard [FF82a, MHGG79, MHGG80, MN89a, FF82b, GK79, Nam86]. Stanford [Kah69, KKM81, Kam69a, Kam69b, Pes73, Rin71]. Stanly [Ste82]. State [ANGK^{+}87, CNxx, Vos89, FBC86, KS88, LS79, LS80, RT88a, Tha89a, Tha89b]. Statements [Sht77, vdH86a]. States [FC85, Vor77, YP88, YP89]. Static [GP82, Kag88, McC81, McC83]. Stationary [LPSV82, McC82, MMN89, Per84b, PLSV84, Per84c, Per86b, Vor77]. Statist.-Simula [NK84]. Statistical [KK85, Kat85, WW86]. Statistician [BJ87]. Statisticians [BJ87]. Statistics [Ken89a, Ken89b, Mac89b]. Status [Bro69, HGB82, Sch82a]. Stauffer [WH89]. Steady [Kot86]. Steeb [LSS86, SS84]. steepest [CFW87]. Steinberg [Ste82]. Steinhauser [BS98]. Stellar [BCTC75]. Steps [Kau73]. Steuerwald [SKxx, TS76]. Steven [Har79a, Har77, Har79b]. Stiff [KPB85]. Stockholm [CS77, FF82b]. Stolte [JPS87]. Stone [PSA84]. Storage [SM83]. Storer [BS86]. Stoutemyer [Sto77a, Sto75, Sto77b]. Street [GK78]. Streng [LMSZ83]. Strip [GP82]. Strong [Lie75, ZNR78]. Stroscio [SH74]. Strotmann [ASW89]. Structural [Lev71, NM89a, NA79]. Structure [AE86, AE87b, Ger80b, GvH89, Hea77, Hea85, LS79, SK83, Tal84, Wat85]. Structured [SGM87]. Structures [DR86b, Gri76b, HM75, NA79, PH83, PLR85, RT85]. Stuart [Sav90, Stu88, SGT90]. Studiengang [EH86]. Studies [CFW87, KC72, KC73, SW83]. Study [dRD88, GS90, Hea82, WR79]. Suaya [BRS73]. Subcentric [SW83]. subprocess [GK85]. Subresultants [Gri78]. Subroutine [TS76]. Subroutines [DJRR81]. Subscripts [HvH83]. Substitution [KRR88]. Suite [GK78]. Sullivan [BS67]. Sun [Kam69b]. Sun-Perturbed [Kam69b]. Sundblad [CLS76]. SUPERCALC [Sei91]. Supercampos [dS87]. Supercomputer [MMN88]. Superfields [dSS85]. Supergravity [GK80]. Supersymmetric [DMMS86]. Supersymmetry [Dar86, dS89, GK80]. Suppes [ST89a]. surface [WW86]. Surguladze [ST89b]. Surprising [AD88]. Susceptibility [DM85].

Suspension [BT78]. SUSY [MDMS87]. Sutera [ZRS87]. Suzuki [KMS86]. Svinolupov [GSSZ87]. Swallowtail [CCF84, CCEL84]. Symb [dS89, GSZ85, GZ90a]. Symbol [Ste82]. Symbolic [Abi83, AH73, AB89, BGK86, BS87, BCRS89, BH79, CH70, Cam74, Cap90, CAH90, Dew89, DG75, DJ89, DR86b, Eis90, EH87, EC87, Fit85b, FH87, FBC86, FB89, FP85, FGPF88, FGP89, GQ86, GvPV79, Gat90b, GM88, GZ84, Gra81, Har79a, Har77, Har79b, Hea72a, Hea76c, Hea76b, Hea81, HI89, Ken88, Köl85a, Lev71, LP90, LM74, Mac87, Mar85, MS85, MMN89, MIG85, NM89a, Ng89, NA79, ND79, QG84, RT85, RT89, Sag88, Sav90, Sch85b, SvH82, Squxx, SKxx, Sto77a, Sto75, ST89a, Tao90, TC88, Tro89, UYSA89, vH87, Wanxx, Wat85, WD77]. Symbolic-Numeric [vH87]. SYMCD [Ito94]. Symmetric [BH84a, BSW88, LH70, McC82, McC83, RN86]. Symmetries [EFK85, FK87, Ito94, Ker83, KF87, Sch82c, Sch82b, Sch83a, SS84, Sch85a, Sch87, Sch88b]. symmetrischen [Gat90a]. symmetry [Gat90b, LP79, LPSV82, Per86b, PSA84]. Synchronous [Kam78]. Syntactic [Mur85]. Synthesis [HM75]. System [AM90, CGGK86, Dul87, EH87, EC87, Fit90, GS90, GMS90, GL89, GLB75, GS77, Gri75, GK78, Har79a, Hea82, Ina80, KS81a, Kat85, KO83, KR85a, KR87, KRR88, LM74, Mar78, MS85, NS82, NS85, Ono79, Ray87, Rod84, SGM87, Som77, Spi87, ST89b, Tou79, VR87]. Systeme [BC90]. Systems [BGK86, Coh89, DST88, DG75, GMS90, Gat90b, GZ90b, GKZ90, GSSZ87, GZ90a, GZ84, GDRH85, Gri74, Gri76a, HWH89, Har97, Hie83b, HGDR84, Ito94, KMS86, KMH88, Kre88, Mac89a, SM83, SS84, Sch85b, Sch86, Sch87, UYSA89, vH82, vHC83, WD77]. Syuichi [UYSA89]. Szego [TSM86]. Szymanowski [LSS87].

T [CCH70, KE86a]. t-dependent [KE86a]. t-independent [KE86a]. Taank [BCTC75]. Tables [Nor80]. Takahashi [ST89a, ZY75]. Takaji [UYSA89]. Takashi [Som85]. Tallents [Tal84]. Tandem [Sha87]. Tao [Tao90]. Taranov [IKRT89, RT87b, RT87a, RT88a]. Tarasov [GTS80]. Targets [Garxx, Gar80]. Tasso [KPT71, TS76]. Tateaki [KS81a, Sas79]. Tatone [RT85]. Taub [BH84a]. Taub-NUT [BH84a]. Taylor [QGxx, QG84, vH80]. Technical [Wanxx]. Technion [Sht75]. Technion-Israel [Sht77]. Technique [GP82, LS80, PRH90, WD77]. Techniques [Bro71, GK85, KPB85, PH81]. Technischer [BW90]. Technology [CLS76, CB76, ESP88, Gat85c, Gat85b, HvH77, HvHS81, Hus81, Kob84, Sht77, SvH82, vdH86b, vH81, vH88, vH89, Wat85]. Teich [Garxx]. Telegraph [Gre84]. Ten [KE86a, KE86b, DCM^{+}89]. Ten-dimensional [DCM^{+}89]. Tensor [AE86, AE87b, Fla86, Gro88b, IK96, KK85, ZY75]. tensoric [RT88a]. Tergimen [HT87]. termorderings [Kre88]. terms [HG89, Sto75, WD77]. Terry [CAH90]. Test [BGJ72]. Tests [BD88, GBB72, JW86]. Tetrahedron [LV84]. Tetsuhiko [IC90]. TEX-REDUCE-Interface [ASW89]. Texas [Par71]. Thas [Tha89a, Tha89b]. their [Abb89, Hor83, LExx, MS84]. Theodore [Garxx, Gar84]. Theor [BHHS87, LP79, Van76, zZlH84]. Theorem [Sch83b, ST89a]. Theorem-prover [ST89a]. Theorems [CCH70]. Theoretic [KR88]. Theoretical [BH84b, CNxx, dSS85, GMS86, Hea72a, Ito90, KRxx, LS79, LS80, NMJ84a]. Theories [BSW88, BGH88, BKdCW81, KR82, KRxx, KR88, LExx, Cap89]. Theorique [Tou87, Cap84]. Theory [ANGK^{+}87, AB89, BH84a, BHHS87, BG87, BGHM88, BNdP^{+}87, Boy78, Bro70, BR72, BRS73, Vos89, Geo68, GK89, Gra81, GBB72, IY77, IK87, Kam69a, Kam69b, Mac87, McC87, MMH88, Nik87, Ozixx, Par71, Pic69, SHM87, Sch84, ST89b, WTV87, WTVNF87]. Thermal [SM83]. Thin-Walled [PLR85]. Third [Wanxx]. Thiry [RT88b, RT89]. Three [BS98, DGH^{+}86, Hie87b, Hie87c, Hie87d, Hie88, Loo72, LPSV82, Per84b, Per84c, Per86b, Pic69]. three- [Hie87b, Hie87c, Hie87d, Hie88]. Three-Dimensional [Pic69]. Three-loop [BS98]. Three-Space [LPSV82, Per84b, Per84c, Per86b]. Tides [ANK81]. Time [DG75, Hor83, Ito90, Kah69, WTV87, WD77]. Time-Dependent [WTV87]. Time-Varying [WD77]. Times [PH83, DCM^{+}89]. Tinoco [BT78]. Tjon [FT75]. Tkachov [ST89b]. Todd [TC88]. Tokamaks [KG75]. Tokyo [HI89, Kan75, MG88, NM89a, ON90, Ono79, Wat79]. Tombal [DCM^{+}89]. TOMS [Gri76a, Gri77a, Sto77a, Sto77b, JW86]. Tool [GQ86, QG84]. Tools [KRR88]. Top [FF82b]. Topography [IY77]. Torodially [RN86]. Torques [PSA84]. Torres-Vega [WTV87, WTVNF87]. Torsion [BH84a, BH84b, BHM86, BMH87, BGHM88]. Toshiaki [KK89]. Toshima [KT85]. Tóth [TSM86]. Touching [JO84]. Tournier [DST88, DT81, DT84, Tou79, Tou87]. Trace [ZY75]. Traces [Hea71a, IKRT89, Lie75]. Tracking [BW90]. trade [Dav82a]. trade-off [Dav82a]. Trans [Ada83]. Transactions [Gri78]. transcendental [Hie84b]. Transcendents [Wat79]. Transfer [SM83]. Transform [LS80, Nor80]. Transformation [Mag81, Par71, SH72, WTV87, WTVNF87]. Transformations [Kaz87, Ker86]. Transforms [Kam69b]. Transient [SM83, Sch77]. Transitions [Gro88c]. Translation [WR79]. Translator [Mar78]. Transport [GP74]. Transverse [ADB^{+}84a]. Treatment [Loo72, Sag88, WD90]. Tree [GK85]. Triads [Ken88]. Trial [Hea79].

Triangular [Dul87, Kam69b]. Triaxiality [Kam78]. Trigonometric [SK83]. Trotter [Tro89]. Tsai [BT63, DT69, TH65, Tsa74]. Tübingen [WD84]. Tupling [zZlH84]. Turbine [GQ86, QG84]. Turbines [QGxx]. Turbulent [MHK86]. Turning [PA86]. Tutorial [BJ87]. Twente [Gat85c, Gat85b, HvH77, HvHS81, vdH86b, vHvH88, vH81, vH88, vH89]. Twisted [Kot86]. Two [BOH78, dSS85, Hie83b, JO84, Kee83, KR82, MMK81, Pes73, Sch77, Sch82c, SL74, Woo89]. Two-Channel [MMK81]. two-dimensional [Hie83b, KR82]. Two-loop [dSS85]. Two-Photon [Pes73]. type [AM90, DCM^{+}89, DR86a, FC84, GK89, Hie83a, Hie87b, Hie87c, Hie87d, IY77, McC81, NZ82]. Tze [Ng89].

U [GKR85, KR82]. Ücoluk [GH82]. Ultraviolet [LM85]. Umeno [UYSA89]. Umfangreicher [LM73]. un-ta [EKR89]. Undergoing [KM80]. Undulator [BW90]. Unequal [Hor83]. unfolding [Kee83]. Uniform [CCF84, CCEL84, IK85]. Uniformly [PLR85]. Uniqueness [Ken89b]. Unitary [Woo89]. Univ [Boy78, LP79, LS79, LS80, Sto75]. Univariate [PB90, WD84]. Universitärer [EH86]. Universität [BC90, EH86, GM85, WD84]. Université [Cap84, DDDD85]. Universiteit [Van79]. University [ABD89, BH84b, BDS88, Cal72a, CS77, Vos89, DLS88, Eas93, Ede81, FBC86, Fla86, FF82b, Gat85c, Gat85b, GK79, HWH89, HvH77, HM75, HvHS81, Hus81, Kah69, KKM81, Kam69a, Kam69b, Kan75, Ken89a, Ken89b, Kob84, KS88, Mac86a, MG88, NS82, Ono79, Par71, Pes73, Rao85, McC84b, RT88a, SH72, Squxx, Tha89a, Tha89b, vdH86b, vHvH88, vH81, vH88, vH89, Wat76, Wat85]. Unphysical [LP79]. Unusual [DR86b]. upon [Ren92]. Upper [Say87b].

Upperbounds [DP85]. Ure [PFF83]. USA [Car73, GK78, Kau73, Rao85, Rin71]. Usage [KRxx]. Use [AH73, CH85, CLS76, Coh89, Vos89, ESP88, GZ84, Gri76b, HM75, LV84, Nor80, McC84b, Sht77, Tro89, VR87]. Used [CDS88, Kre88]. User [DJRR81, Fit81, Gat85c, Hea87, vH82]. User-based [Fit81]. Using [CB76, FC79, dRD88, dSS85, dS89, Dul87, GKRR85a, GKR85, GKRR85b, GK80, Gri78, Gri74, Gri77b, Hea79, LD87, Mag81, MMN88, NM89a, Ng89, Nik87, ON90, Sch77, Sto77a, Sto77b, Wan85]. USP [GTS80]. USSR [KRxx, Gro88a, Gro88b, Gro88c, vH87]. Utah [Cal72a, GK79, Sto75, Nam86]. Utilisation [Cap84, Fit87].

vacuum [BS98, Kad88, LPSV82, McC81, McC83, Per84b, PLSV84, Per84c, Per86b]. Valentini [LPSV82, PLSV84]. validation [Bam88]. Value [BS87, Cap86a, Dau83, DD85, RA89, Kea85]. Values [Alf82]. Variable [Eis90]. Variables [BOH78, Rod84]. Variant [GM85]. Varying [Hor83, WD77]. VAX [FF82a]. Vaxima [NW83]. Vector [Cam68, Eas91, GKRR85a, GKR85, GKRR85b, Hap89, NM88, RN86, NM89b]. Vector33 [Hap89]. Vectorization [GvH89]. Vectorparticles [Van79]. Verbaeten [GvPV79]. Verfahren [Mac73]. Verification [GLB75, LM74, Sch77, Wat79]. Vermessungswes [Kra82]. Vermessungswesen [EH86]. version [Eas91, Gat85c, Gat85b, Hea87]. versus [Hie84a]. Vertex [Adk83, Adk85]. Vertical [Feu84]. Via [Gri76a]. Viana [GV86]. Vibration [GQ86, Han87]. Vibration-Rotation [Han87]. Vibrations [Sag88]. Vicinity [AZZ87]. Victor [GS90]. Viewed [vH82]. Vigneron [LV84]. Viktor [EKR89]. Vinitsky [BCM^{+}95, VR87]. Violating [BLMN78]. Virginia [Squxx].

Virial [GP74]. Viscous [Rao85, ZRS87]. Vladimar [GKZ90]. Voisinage [DT81]. Volker [ASW89, WH89]. Voltages [Hor83]. Volterra [MMR87]. voor [Van76]. Voros [BPV78, GVxx, GV78, Vor77]. vortex [Kee90]. Vos [Vos89]. Vries [Sch82c].

W [BS67, GG78, Per78, Pes73, TH65]. W- [GG78, Per78, TH65]. W-Boson [BS67]. W-Pair [Pes73]. Waals [ANGK^{+}87]. Wako-Shi [IY77]. Waldir [RdS88]. Walles [CFW87]. Wallis [BDM^{+}87]. Wan-zhen [zZlH84]. Wanas [Wanxx, Wan85]. Wang [NW83, WCvH84, AD88]. want [Dav82b]. Ward [LM85, ZY75]. Ward-Takahashi [ZY75]. Warren [CAH90]. Warwick [Ken89a, Ken89b]. Washington [Rao85]. Wassam [WTV87, WTVNF87]. Watanabe [AM90, Bra90, Fit90, GS90, GMS90, Gat90b, GZ90b, GKZ90, HW81, LD90, Nor90, PB90, Tao90, Wat76, Wat79, Wat85]. Water [Gro88b]. Watt [FGP89, GL89, GK89, Tro89]. Wave [Fle71, Kee89, Say87b, SH81]. Waves [Boy78, DLS88, FGM73, Gro88b, MS84]. Weak [BCTC75, Gro88c]. Wealth [KT85]. Weber [WR79]. Wehner [WW86]. Weidner [JW86]. Weierstrass [GV86, Kob84]. Weight [KRxx, KR88]. Well [BBF^{+}82]. Werner [ASW89, Sei91]. Wess [dS87, dSS85]. West [DD85, Squxx]. Which [EH87, GKZ90]. Whistler [Gla82]. Wiggler [BW90]. William [FBC86]. Wilson [SW83]. Wind [GQ86, QGxx, QG84]. Winfried [MN89a, MN89b, NM88, MN89c]. Winkelmann [ASW89, BSW88, WH89]. Winkler [WKL88]. Wisconsin [Wat85]. Wise [AW88]. Wissenschaften [DJ80]. Witham [WD77]. within [AZZ87, RT87b]. Wm [eS85]. Wolfer [WW86]. Wood [Woo89]. Woodrow [QGxx, QG84]. Wooff [HWH89]. Work [Rod84, RT88a]. World [Dav88, Hea85, Kat85, KT85, LOG85, Mar85, MIG85, Per88, Sch87, Som85]. Wright [Eas93, WD84]. Writing [Mar78]. Wroclaw [PR84]. Wronskian [Ito88]. Wu [BKdCW81]. Wulkow [WD90]. Wüstefeld [BW90].

x [CC88]. xii [Art95]. XV [DL88b].

Ya [Dar86, DMR88, DR89, DR85, DR86a, IKRT89, KRxx, KR85a, KR85b, KRR87, KR87, KRR88, KR88, MDMS87, Rod84, RT87b, RT87a, RT88a]. Yamamoto [YA87]. Yamashita [UYSA89]. Yang [KM84]. Yannouleas [YP88, YP89]. Yasumasa [KS81a]. Yatagai [IY77]. Yennie [HKY69]. Yoichi [Wat85]. York [Abb89, Dew89, Fit89b, FGP89, GL89, GK89, Tro89, vHHGV89]. Yoshida [LSS86]. Young [ZY75]. Yu [CS77, GS90, GZ90b, GKZ90, GSZ85, GSSZ87, GZ90a, IKRT89, RT87b, RT87a, RT88a].

Zabolitzky [WH89]. Zacrep [ZY75]. Zahalak [ZRS87]. Zeitschrift [Kad88]. Zeng [zZlH84]. Zentralinstitut [DJ80]. Zernas [LMSZ83]. Zero [Garxx, Gar80, Pie84]. Zeros [KR82]. Zharkov [GZ90b, GKZ90, GSZ85, GSSZ87, GZ90a]. Zhidkova [ZNR78]. Zhydkov [AZZ87]. Zhydkova [AZZ87]. Zimanyi [BZZ86, GZ84, NZ82, NZ87]. Zirconium [Say87a]. zum [PN73a]. Zumino [dS87, dSS85]. zur [Kra82]. zweiten [Kra82]. ZZ [GG78].

References

Appelquist:70

[AB70] T. W. Appelquist and S. J.

Brodsky. The order α^{2} electrodynamic corrections to the Lamb shift. Phys. Rev. Letters, 24:562–565, 1970.

Adkins:83a

[AB83] G. S. Adkins and F. R. Brown.

Rate for positronium decay to five photons. Phys. Rev. A, 28: 1164–1165, 1983.

Autin:89

[AB89] B. Autin and J. Bengtsson. Symbolic evaluation of integrals occurring in accelerator orbit theory. J. Symbolic Computation, 7 (2):183–187, February 1989.

Abbott:87a

[Abb87] J. A. Abbott. Integration: Solving the Risch differential equation. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 465– 467. Springer-Verlag, 1987.

Abbott:88

[Abb88] J. A. Abbott. Factorisation of Polynomials over Algebraic Number Fields. PhD thesis, Univ. of Bath, England, 1988.

Abbott:89

[Abb89] J. A. Abbott. Recovery of algebraic numbers from their p-adic approximations. In Proc. of IS-SAC ’89, pages 112–120. ACM Press, New York, 1989.

Abbott:85

[ABD85] J. A. Abbott, R. J. Bradford, and J. H. Davenport. A remark on factorisation. SIGSAM Bulletin, 19(2):31–33, May 1985.

Abbott:86

[ABD86a] J. A. Abbott, R. J. Bradford, and J. H. Davenport. The Bath algebraic number package. In Proc. of SYMSAC ’86, pages 250–253, 1986.

Arbuzov:86

[ABD86b] B. A. Arbuzov, E. E. Boos, and A. I. Davydychev. Infrared asymptotics of gluonic Green functions in covariant gauge. Preprint 86-123, IHEP, 1986.

Abbott:89a

[ABD89] J. A. Abbott, R. J. Bradford, and J. H. Davenport. A remark on the multiplication of sparse polynomials. Technical Report TR 89-21, School of Mathematical Sciences, University of Bath, 1989.

Aldins:69

[ABDK69] J. Aldins, S. J. Brodsky, A. J. Dufner, and T. Kinoshita. Photon-photon scattering contribution to the sixth order magnetic moments of the muon and electron. Phys. Rev. Lett., 23: 441–443, 1969.

Abiezzi:83

[Abi83] Salim S. Abi-Ezzi. Clarification to the symbolic mode in REDUCE. SIGSAM Bulletin, 17(3 and 4):43–47, August and November 1983.

Abbott:87

[AD87] J. A. Abbott and J. H. Davenport. Polynomial factorization: An exploration of Lenstra’s algorithm. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 391– 402. Springer-Verlag, 1987.

Abbott:88a

[AD88] J. A. Abbott and J. H. Davenport. A remark on a paper by Wang: Another surprising property of 42. Math. Comp., 51:837– 839, 1988.

Adams:83

[Ada83] K. J. Adams. Analytic estimates for the dynamic aperture of nonlinear lattices. IEEE Trans. Nucl. Sci., NS-30:2436– 2438, 1983.

Aurenche:84

[ADB^{+}84a] P. Aurenche, A. Douir, R. Baier, M. Fontannaz, and D. Schiff. Photoproduction of hadrons at large transverse momentum in second order QCD. Phys. Lett., 135B:164–168, 1984.

Aurenche:84a

[ADB^{+}84b] P. Aurenche, A. Douir, R. Baier, M. Fontannaz, and D. Schiff. Prompt photon production at large p_{τ} in GCD beyond the leading order. Phys. Lett., 140B: 87–92, 1984.

Adkins:83

[Adk83] G. S. Adkins. Analytic evaluation of an O(α ) vertex correction to the rate of orthopositronium. Phys. Rev. A, 27:530–532, 1983.

Adkins:85

[Adk85] G. S. Adkins. Inner-vertex contributions to the decay rate of orthopositronium. Phys. Rev. A, 31:1250–1252, 1985.

Alekseev:86

[AE86] A. I. Alekseev and V. F. Edneral. Tensor structure of axial gauge polarization operator in the infrared region. Preprint 86-46, IHEP, 1986.

Alekseev:87a

[AE87a] A. I. Alekseev and V. F. Edneral. On evaluation of Feynman integrals in axial gauge. Preprint 87-118, IHEP, 1987.

Alekseev:87

[AE87b] A. I. Alekseev and V. F. Edneral. Tensor structure of gluon polarization operator in the axial gauge for infra-red region. Journal of Nuclear Physics, pages 1105–1114, 1987.

Atherton:73

[AH73] R. W. Atherton and G. M. Homsey. Use of symbolic computation to generate evolution equations and asymptotic solutions to elliptic equations. Journ. Comp. Phys., 1:45–59, 1973.

Alfeld:82

[Alf82] P. Alfeld. Fixed point iteration with inexact function values. Math. Comp., 38:87–98, 1982.

Adamchik90

[AM90] V. S. Adamchik and O. I.

Marichev. The algorithm for calculating integrals of hypergeometric type functions and its realization in REDUCE system. In S. Watanabe and Morio Nagata, editors, Proceedings of the International Symposium on Symbolic and Algebraic Computation, pages 212–224. ACM, Addison-Wesley, 1990.

Aguilera-Navarro:87

[ANGK^{+}87] V. C. Aguilera-Navarro, R. Guardiola, C. Keller, M. de Llano, M. Popovic, and M. Fortes. Van der Waals perturbation theory for fermion and boson groundstate matter. Phys. Rev. A, 35: 563–584, 1987.

Aso:81

[ANK81] T. Aso, T. Nonoyama, and S. Kato. Numerical simulation of semidiurnal atmospheric tides. J. Geophysical R., 86(11):388– 400, 1981.

Arter:1995:RPM

[Art95] W. Arter. Book review: RE-DUCE for physicists: By N. MacDonald. Institute of Physics Publishing, 1994. xii + 167 pages. Price £25 (paperback). ISBN 0-7503-0277-1. Computer Physics Communications, 85 (2):323, February 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0010465595900578.

Antweiler:89

[ASW89] Werner Antweiler, Andreas Strotmann, and Volker Winkelmann. A TEX-REDUCE-Interface. SIGSAM Bulletin, 23: 26–33, February 1989.

Abdali:88

[AW88] S. K. Abdali and D. S. Wise. Experiments with quadtree representation of matrices. In Proc. of ISSAC ’88, volume 358, pages 96–108. Springer-Verlag, 1988.

Amirkhanov:87

[AZZ87] I. V. Amirkhanov, E. P. Zhydkov, and I. E. Zhydkova. The conditions of bounding of the oscillation amplitudes of charge particle within the resonance vicinity investigations. Technical Report P11-87-452, J.I.N.R., Dubna, 1987.

Bamberger:1988:EVP

[Bam88] L. Bamberger. An error validation package for REDUCE. [ESP88], pages 467–474 (vol. 1). ISBN 0-444-87145-4. LCCN QA75.5.E84 1988. Two volumes. Publication no. EUR 11852 of the Commission of the European Communities, Directorate-General Telecommunications, Information Industries, and Innovation, Luxembourg.

Barthes-Biesel:73

[BBA73] D. Barthes-Biesel and A. Acrivos.

On computer generated analytic solutions to the equations of fluid mechanics, the case of creeping flows. Journ. Comp. Phys., 3:403–411, 1973.

Baker:81

[BBF^{+}82] G. A. Baker, L. P. Benofy, M. Fortes, M. de Llano, S. M. Peltier, and A. Plastino. Hardcore square-well fermion. Phys. Rev. A, 26:3575–3588, 1982.

Bessis:85

[BBR85] N. Bessis, G. Bessis, and D. Roux. Closed-form expressions for the Dirac–Coulomb radial r^{t} integrals. Phys. Rev. A, 32:2044–2050, 1985.

Bordoni:81

[BC81] Luciana Bordoni and Attilio Colagrossi. An application of RE-DUCE to industrial mechanics. SIGSAM Bulletin, 15(2):8–12, May 1981.

Bittencourt:90

[BC90] Guilherme Bittencourt and Jacques Calmet. Integrating computer algebra and knowledge representation. Preprint, Universität Karlsruhe Institut für Algorithmen und Kognitive Systeme, 1990.

Brackx:87a

[BCDS87] F. Brackx, D. Constales, R. Delanghe, and H. Serras. Clifford algebra with REDUCE. Rend.

Circ. Mat. Palermo, Ser. II, 16: 11–19, 1987.

Basios:1995:GRP

[BCM^{+}95] V. Basios, N. A. Chekanov, B. L. Markovski, V. A. Rostovtsev, and S. I. Vinitsky. GITA: a REDUCE program for the normalization of polynomial Hamiltonians. Computer Physics Communications, 90(2– 3):355–368, October 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/001046559500080Y.

Brackx:89

[BCRS89] F. Brackx, D. Constales, A. Ronveaux, and H. Serras. On the harmonic and monogenic decomposition of polynomials. J. Symbolic Computation, 8(3): 297–304, September 1989.

Biswas:75

[BCTC75] S. N. Biswas, S. R. Chaudhuri, K. S. Taank, and J. A. Campbell. Neutrino production in stellar matter by photons in a renormalizable scalar-boson-exchange model of weak interactions. Phys. Rev. D, 8:2523– 2525, 1975.

Bradford:88

[BD88] R. J. Bradford and J. H. Davenport. Effective tests for cyclotomic polynomials. In Proc. of ISSAC ’88, volume 358, pages 244–251. Springer-Verlag, 1988.

Bowyer:87

[BDM^{+}87] A. Bowyer, J. H. Davenport, P. S. Milne, J. A. Padget, and A. F. Wallis. Applications of computer algebra in solid modelling. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 244– 245. Springer-Verlag, 1987.

Bennett

[BDS88] J. P. Bennett, J. H. Davenport, and H. M. Sauro. Solution of some equations in biochemistry. Technical Report 88-12, School of Mathematical Sciences, University of Bath, England, 1988.

Baekler:87a

[BG87] P. Baekler and M. Guerses. Exact solutions of the Poincaré gauge theory from its linearized field equations. Lett. Math. Phys., 14:185–191, 1987.

Baekler:88b

[BGH88] P. Baekler, M. Guerses, and F. W. Hehl. A new method to solve the field equations of Poincaré gauge theories. Class. Quantum Grav., 1988.

Baekler:88a

[BGHM88] P. Baekler, M. Guerses, F. W.

Hehl, and J. D. McCrea. The exterior gravitational field of a charged spinning source in the Poincaré gauge theory: A Kerr– Newman metric with dynamic torsion. Phys. Lett., A128:245– 250, 1988.

Brodsky:72a

[BGJ72] S. J. Brodsky, J. F. Gunion, and R. L. Jaffe. Test for fractionally charged partons from deepinelastic bremsstrahlung in the scaling region. Phys. Rev. D, 6: 2487–2494, 1972.

Boege:86

[BGK86] W. Boege, R. Gebauer, and H. Kredel. Some examples for solving systems of algebraic equations by calculating Groebner bases. J. Symbolic Computation, 2(1):83–98, March 1986.

Berkovich:89

[BGKN89] L. M. Berkovich, V. P. Gerdt, Z. T. Kostova, and M. L. Nechaevsky. Second order reducible linear differential equations. Preprint E5-89-141, J.I.N.R., Dubna, 1989.

Barton:72

[BH72] David Barton and Anthony C. Hearn. Comment on problem #2 - the Y(2n) functions. SIG-SAM Bulletin, 15, 1972.

Brown:79

[BH79] W. S. Brown and A. C. Hearn. Applications of symbolic algebraic computation. Comp. Phys. Comm., 17:207–215, 1979.

Baekler:84

[BH84a] P. Baekler and F. W. Hehl. A charged Taub-NUT metric with torsion: A new axially symmetric solution of the Poincaré gauge field theory. Phys. Lett., 100A:277–316, 1984.

Baekler:84a

[BH84b] Peter Baekler and Friedrich W.

Hehl. On the dynamics of the torsion of spacetime: Exact solutions in a gauge theoretical model of gravity. Technical Report UCLA/84/TEP/19, Department of Physics, University of California, Los Angeles, December 1984.

Bogdanova:88

[BH88] N. Bogdanova and H. Hogreve.

A REDUCE package for exact Coulomb interaction matrix elements. Comp. Phys. Commun., 48(2):319–326, February 1988.

Baekler:87

[BHHS87] P. Baekler, R. Hecht, F. W.

Hehl, and T. Shirafuji. Mass and spin of exact solutions of the Poincaré gauge theory. Prog. Theor. Phys., 78:16–21, 1987.

Baekler:86

[BHM86] P. Baekler, F. W. Hehl, and E. W. Mielke. Nonmetricity and torsion: Facts and fancies in gauge approaches to gravity. In R. Ruffini, editor, Proc. 4th Marcel Grossmann Meeting on General Relativity, ed., pages 277–316. North-Holland, Amsterdam, 1986.

Brodsky:62

[BHP62] S. J. Brodsky, A. C. Hearn, and R. G. Parsons. Determination of the real part of the Compton amplitude at a nucleon resonance. Phys. Rev., 187:1899– 1904, 1962.

Bradford:86

[BHPS86] R. J. Bradford, A. C. Hearn, J. A. Padget, and E. Schrüfer. Enlarging the REDUCE domain of computation. In Proc. of SYMSAC ’86, pages 100–106, 1986.

Bilge:1992:RPI

[Bil92] Ayşe Hümeyra Bilge. A RE-DUCE program for the integration of differential polynomials. Computer Physics Communications, 71(3):263–268, September 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/001046559290013O.

Birrell:77

[Bir77] N. D. Birrell. The application of adiabatic regularization to calculations of cosmological interest. Technical report, Dept. Math, King’s College, London, 1977.

Bryan-Jones:87

[BJ87] Jane Bryan-Jones. A tutorial in computer algebra for statisticians. The Professional Statistician, 6(6):TBD, December 1987.

Berends:81

[BKdCW81] A. Berends, R. Kleiss, P. de Causmaecher, and T. T. Wu. Single bremsstrahlung process in gauge theories. Phys. Lett., 103B:124– 128, 1981.

Billoire:78

[BLMN78] A. Billoire, R. Lacaze, A. Morel, and H. Navelet. The OZI rule violating radiative decays of the heavy pseudoscalars. Report DpH-T 43/78, CEN-Saclay, 1978.

Baekler:87b

[BMH87] P. Baekler, E. W. Mielke, and F. W. Hehl. Kinky torsion in a Poincaré gauge model of gravity coupled to a massless scalar field. Nuclear Phys., B288:800– 812, 1987.

Biro:87

[BNdP^{+}87] T. S. Biro, K. Niita, A. L.

de Paoli, W. Bauer, W. Cassing, and U. Mosel. Microscopic theory of photon production in proton-nucleus and nucleusnucleus collisions. Nuclear Physics, 475A:579–597, December 1987.

Bocko:1992:ERB

[Boc92] J. Bocko. EQSHELL — a REDUCE-based program for generation of equations of equilibrium for shells. Computer Physics Communications, 69(1): 215–222, February 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/001046559290141K.

Bajla:78

[BOH78] I. Bajla, G. A. Ososkov, and A. C. Hearn. The orthogonalization program of polynomials in two variables in REDUCE-2 language. Report P10-11944, J.I.N.R., Dubna, 1978.

Boyd:78

[Boy78] John P. Boyd. The effects of latitudinal shear on equatorial waves, part I: Theory and methods. Preprint, Dept. of Atmos. and Oceanic Science, Univ. of Michigan, January 1978.

Balian:78

[BPV78] R. Balian, G. Parisi, and A. Voros. Quartic oscillator. In Proc. of the Colloq. on Mathematical Problems in Feynman Path Integrals, Marseille, May 1978.

Brodsky:72b

[BR72] S. J. Brodsky and R. Roskies. Quantum electrodynamics and renormalization theory in the infinite momentum frame. Phys. Lett., 41B:517–520, 1972.

Bradford90

[Bra90] Russell Bradford. A parallelization of the buchberger algorithm. In S. Watanabe and Morio Nagata, editors, Proceedings of the International Symposium on Symbolic and Algebraic Computation, page 296. ACM, Addison-Wesley, 1990.

Brodsky:69

[Bro69]  S. J. Brodsky.  Status of

quantum electrodynamics. In Proc. International Symposium on Electron and Photon Interactions at High Energies, Liverpool, England, 1969.

Brodsky:70

[Bro70] S. J. Brodsky. Quantum electrodynamic theory: Its relation to precision low energy experiments. Report SLAC-PUB-795, SLAC, August 1970.

Brodsky:71

[Bro71] S. J. Brodsky. Algebraic computation techniques in quantum electrodynamics. In Proc. 2nd Computing Methods in Theoretical Physics, Marseilles, volume II, pages IV–1 to IV–27, 1971.

Brodsky:72

[Bro72] S. J. Brodsky. Atomic physics and quantum electrodynamics in the infinite momentum frame. Report SLAC-PUB-1118, SLAC, August 1972.

Broughan:82

[Bro82] K. A. Broughan. Grad–Fokker– Planck plasma equations. part 1. Collision moments. J. Plasma Phys., 27:437–452, 1982.

Broadhurst:85

[Bro85] D. J. Broadhurst. Evaluation of a class of Feynman diagrams for all numbers of loops and dimensions. Phys. Lett. B, 164:356– 360, 1985.

Brodsky:73

[BRS73] S. J. Brodsky, R. Roskies, and R. Suaya. Quantum electrodynamics and renormalization theory in the infinite-momentum frame. Phys. Rev. D, 8:4574– 4594, 1973.

Brodsky:67

[BS67] S. J. Brodsky and J. D. Sullivan. W-boson contribution to the anomalous magnetic moment of the muon. Phys. Rev., 156:1644–1647, 1967.

Bateman:86

[BS86] G. Bateman and R. G. Storer. Direct determination of axisymmetric magnetohydrodynamic equilibrium in Hamada coordinates. Journ. Comp. Phys., 64: 161–176, 1986.

Brackx:87

[BS87] F. Brackx and H. Serras. Boundary value problems for the Laplacian in Euclidean space solved by symbolic computation. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 208– 215. Springer-Verlag, 1987.

Baikov:1998:TLV

[BS98] P. A. Baikov and M. Steinhauser. Three-loop vacuum integrals in FORM and RE-DUCE. Computer Physics Communications, 115(2–3):161– 169, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465598001258.

Baekler:88

[BSW88] P. Baekler, M. Seitz, and V. Winkelmann. Cylindrically symmetric solutions of selfconsistently coupled Dirac fields in gauge theories of gravity. Class. Quantum Grav., 5:479– 490, 1988.

Berman:63

[BT63] S. M. Berman and Y. S. Tsai.

Intermediate boson pair production as a means for determining its magnetic moment. Phys. Rev. Lett., 11:483–487, 1963.

Bark:78

[BT78] Fritz H. Bark and Herman Tinoco. Stability of plane Poiseuille flow of a dilute suspension of slender fibres. J. Fluid Mech., 87:321–333, 1978.

Bahrdt:90

[BW90] J. Bahrdt and G. Wüstefeld. A new tracking routine for particles in undulator and wiggler fields. Report BESSY TB Nr. 158, Technischer Bericht, October 1990.

Biro:86

[BZZ86] T. S. Biro, J. Zimanyi, and M. Zimanyi. Hadrochemistry in relativistic mean fields. Physics Letters, 167B(3):271– 276, February 1986.

Cline:90

[CAH90] Terry Cline, Harold Abelson, and Warren Harris. Symbolic computing in engineering design. AI EDAM, February 1990.

Calmet:72

[Cal72a] Jacques Calmet. Further evaluation of the sixth order corrections to the anomalous magnetic moment of the electron.

Technical report, Department of Physics, University of Utah, 1972.

Calmet:72a

[Cal72b] Jacques Calmet. A REDUCE approach to the calculation of Feynman diagrams. Comp. Phys. Comm., 4:199–204, 1972.

Calmet:74

[Cal74] Jacques Calmet. Computer recognition of divergences in Feynman graphs. SIGSAM Bulletin, 8(3):74–75, August 1974.

Campbell:67

[Cam67] J. A. Campbell. Algebraic computation of radiative corrections for electron-positron scattering. Nucl. Phys., B1:283–300, 1967.

Campbell:68

[Cam68] J. A. Campbell. Astrophysical consequences of the existence of charged intermediate vector bosons. Aust. Journ. of Phys., 21:139–148, 1968.

Campbell:74

[Cam74] J. A. Campbell. Symbolic computing and its relationship to particle physics. Acta Physica Austriaca, Suppl. XIII:595–647, 1974.

Caprasse:84

[Cap84] H. Caprasse. Description et utilisation d’une extension du programme REDUCE. Technical report, Physique Théorique et Mathématique, Université de Liège, October 1984.

Caprasse:86a

[Cap86a] H. Caprasse. A complete simplification package for the absolute value function in REDUCE. SIGSAM Bulletin, 20(1 and 2): 18–21, February and May 1986.

Caprasse:86

[Cap86b] H. Caprasse. Description of an extension of the matrix package of REDUCE. SIGSAM Bulletin, 20(4):7–10, December 1986.

Caprasse:89a

[Cap89] H. Caprasse. Les théories des Champs dans le monde de RE-DUCE (in french). CALSYF (to appear), 1989.

Caprasse:90

[Cap90] H. Caprasse. Renormalization group, function iterations and computer algebra. J. Symbolic Computation, 9(1):61–72, January 1990.

Carroll:73

[Car73] R. Carroll. The Anomalous Magnetic Moment of the Electron in the Mass Operator Formalism. PhD thesis, University of Michigan, Ann Arbor, MI, USA, 1973.

Carroll:75

[Car75] R. Carroll. Mass-operator calculation of the electron g-factor. Phys. Rev. D, 8:2344–2354,

Carlson:80

[Car80] P. Carlson. Coordinate free relativity. J. Math. Phys., 21:1149– 1154, 1980.

Cohen:76a

[CB76] I. Cohen and F. Bark. Perturbation calculations for the spin up problem using REDUCE. Technical Report TRITA-MEK-76-03, The Royal Institute of Technology, Department of Mechanics, 1976.

Conwell:84

[CBR84] P. R. Conwell, P. W. Barber, and C. K. Rushworth. Resonant spectra of dielectric sphere. J. Opt. Soc. Am. A, 1:62–67, 1984.

Chaffy:88

[CC88] C. Chaffy-Camus. An application of REDUCE to the approximation of f(x,y). In Proc. of IS-SAC ’88, volume 358, pages 73– 84. Springer-Verlag, 1988.

Connor:84a

[CCEL84] J. N. L. Connor, P. R. Curtis, C. J. Edge, and A. Laganà. The uniform asymptotic swallowtail approximation: Application to the collinear H + F_{2}. J. Chem. Phys., 80(3):1362–1363, February 1984.

Connor:84

[CCF84] J. N. L. Connor, P. R. Curtis, and D. Farrelly. The uniform asymptotic swallowtail approximation: Practical methods for oscillating integrals with four coalescing saddle points. J. Phys. A, 17:283–310, 1984.

Campbell:70a

[CCH70] J. A. Campbell, R. B. Clark, and D. Horn. Low-t theorems for charged-pion photoproduction. Phys. Rev. D, 2:217–224, 1970.

Caprasse:88

[CDS88] H. Caprasse, J. Demaret, and E. Schruefer. Can EX-CALC be used to investigate High-dimensional cosmological models with Non-Linear Lagrangians. In Proc. of ISSAC ’88, pages 116–124. Springer-Verlag, 1988.

Cohen:84

[CFÅ84] H. I. Cohen, I. B. Frick, and J. E.

Åman. Algebraic computing in general relativity. General Relativity and Gravitation, ed., pages 139–162, 1984.

Campbell:87

[CFW87] J. A. Campbell, P. O. Fröman, and E. Walles. Explicit series formulae for the evaluation of integrals by the method of steepest descents. Studies in Applied Mathematics, 77:151–172, 1987.

Cowan:79

[CG79] Richard M. Cowan and Martin L. Griss. Hashing – the key to rapid pattern matching. In Proc. EUROSAM 1979, Lecture Notes in Computer Science, volume 72, pages 266–278. Springer-Verlag, 1979.

Chinnick:86

[CGGK86] K. Chinnick, C. Gibson, J. F. Griffiths, and W. Kordylewski. Isothermal interpretations of oscillatory ignition during hydrogen oxidation in an open system. I. Analytical predictions and experimental measurements of periodicity. Proc. Royal Soc. Lond., A405:117–128, 1986.

Campbell:70

[CH70] J. A. Campbell and A. C. Hearn. Symbolic analysis of Feynman diagrams by computer. Journ. of Comp. Phys., 5:280–327, 1970.

Caprasse:85

[CH85] H. Caprasse and M. Hans. A new use of operators in the algebraic mode of REDUCE. SIG-SAM Bulletin, 19(3):46–52, August 1985.

Cohen:76

[CLS76] H. I. Cohen, O. Leringe, and Y. Sundblad. The use of algebraic computing in general relativity. Technical Report TRITA-MEK-76-02, The Royal Institute of Technology Department of Mechanics, 1976.

Cejchan

[CNxx] A. Cejchan and J. Nadrchal. Application of REDUCE-2 and analytic integration program in the theoretical solid state physics. Technical report, Institute of Physics, CSAV, Prague, 19xx.

Cohen:89

[Coh89] Joel S. Cohen. The effective use of computer algebra systems. In Transactions of the Sixth Army Conference on Applied Mathematics and Computing, pages 677–698, 1989.

Cohen:77

[CS77] I. Cohen and S. Yu. Slavyanov.

Smooth perturbations of the schrödinger equation with a linear potential related to the charmonium models. USIP Report 77-17, University of Stockholm Institute of Physics, 1977.

Cung:75

[Cun75] V. K. Cung. Differential cross section of e+ + e- to e+ + mu- + nubar(mu) + nubar(e). Phys. Lett., 55B:67–70, 1975.

Calmet:83

[CvH83] J. Calmet and J. A. van Hulzen.

Computer algebra applications. In B. Buchberger, G. E. Collins, R. Loos, and R. Albrecht, editors, Computer Algebra Symbolic and Algebraic Computation. Springer-Verlag, 2nd edition, 1983.

Darbaidze:86

[Dar86] Ya. Z. Darbaidze. A gluon bremsstrahlung in supersymmetry QCD. Preprint P2-86-825, JINR, 1986.

Dautcourt:79

[Dau79] G. Dautcourt. Application of REDUCE to algebraic computations in general relativity and astrophysics. In Proc. of the Workshop in Symbolic Computation, Dubna, U.S.S.R., September 1979.

Dautcourt:83

[Dau83] G. Dautcourt. The cosmological problem as an initial value problem on the observer’s past light cone: Geometry. J. Phys. A, 16: 3507–3528, 1983.

Davenport:81

[Dav81] James Harold Davenport. On the integration of algebraic functions. Lecture Notes in Computer Science, 102:1–197, 1981.

Davenport:82

[Dav82a] James H. Davenport. Fast REDUCE: The trade-off between efficiency and generality. SIGSAM Bulletin, 16(1):8–11, February 1982.

Davenport:82a

[Dav82b] James H. Davenport. What do we want from a high-level language? SIGSAM Bulletin, 16 (4):6–9, November 1982.

Davenport:88

[Dav88] J. H. Davenport. The world of computer algebra. New Scientist, 1629:71–72, September 1988. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Demaret:89

[DCM^{+}89] J. Demaret, H. Caprasse, A. Moussiaux, Ph. Tombal, and D. Papadopoulos. Tendimensional Lovelock-type Space-Times. To appear Phys. Rev. D, July 1989.

Dorfi:85

[DD85] E. A. Dorfi and L. O’C. Drury.

Simple adaptive grids for 1D initial value problems. Technical Report MPI H-1985-V21, Max-Plack-Institut fuer Kernphysik, Heidelberg, West Germany, 1985.

Della-Dora:85

[DDDD85] Jean Della-Dora, Claire Dicrescenzo, and Dominique Duval. About a new method for computing in algebraic number fields. Technical report, Université de Grenoble, Institut Fourier, France, November 1985.

Dewar:89

[Dew89] M. C. Dewar. IRENA – an integrated symbolic and numerical computation environment. In Proc. of ISSAC ’89, pages 171–179. ACM Press, New York, 1989.

Dubowsky:75

[DG75] S. Dubowsky and J. L. Grant.

Application of symbolic manipulation to time domain analysis of nonlinear dynamic systems. Journ. of Dynamic Systems, Measurement, and Control, (75-Aut-J), 1975.

Dorizzi:86

[DGH^{+}86] B. Dorizzi, B. Grammaticos, J. Hietarinta, A. Ramani, and F. Schwarz. New integrable three dimensional quartic potentials. Phys. Lett., 116A:432–436, 1986.

Dicrescenzo:85

[Dic85] Claire Dicrescenzo. Algebraic computation on algebraic numbers. Technical report, Institut Fourier, Laboratoire de Mathématiques, France, December 1985.

Dautcourt:80

[DJ80] G. Dautcourt and K. P. Jann. A program package in REDUCE 2 for algebraic computations in general relativity. Technical report, Zentralinstitut fuer Astrophysik der Akademie der Wissenschaften, 1980.

Dudley:89

[DJ89] M. L. Dudley and R. W. James. Computer-aided derivation of spherical harmonic spectral equations in astrogeophyics. J. Symbolic Computation, 8(4): 423–427, October 1989.

Dautcourt:81

[DJRR81] G. Dautcourt, K. P. Jann, E. Riemer, and M. Riemer. User’s guide to REDUCE subroutines for algebraic computations in general relativity. Astron. Nachr., 302:1–13, 1981.

Diver:86

[DL86] D. A. Diver and E. W. Laing. Proc. 8th Europhysics conference on computational physics. In Computing in Plasma Physics, 1986.

Diver:88a

[DL88a] D. A. Diver and E. W. Laing.

Alfven resonance absorption in a magnetofluid. Internal Report GUTPA 88/04-01, July 1988.

Diver:88

[DL88b] D. A. Diver and E. W. Laing.

Proc. XV European conference on controlled fusion and plasma heating. 1988.

Diver

[DLS88] D. A. Diver, E. Q. Laing, and C. C. Sellar. Waves in a cold plasma with a spatially rotating magnetic field. Report GU TPA 88/12-1, Department of Physics and Astronomy, University of Glasgow, Plasma Physics Group, 1988.

Dhar:85

[DM85] D. Dhar and J-M. Maillard. Susceptibility of the checkerboard Ising model. J. Phys. A, 18: L383–L388, 1985.

Darbaidze:86a

[DMMS86] J. Z. Darbaidze, V. A. Matveev, Z. V. Merebashvili, and L. A. Slepchenko. Gluon bremsstrahlung in supersymmetric QCD. Phys. Lett., B177, 1986.

Darbaidze:88

[DMR88] Ya. Z. Darbaidze, Z. V. Merebashvili, and V. A. Rostovtsev. Some computer realizations of the REDUCE-3 calculations for exclusive processes. Preprint P2-88-769, JINR, 1988.

Davenport:85

[DP85] James Davenport and Julian Padget. HEUGCD: How elementary upperbounds generate cheaper data. In Proc. EU-ROCAL 1985, Lecture Notes in Computer Science, volume 204, pages 18–28. Springer-Verlag, 1985.

Demichev:85

[DR85] A. P. Demichev and A. Ya. Rodionov. A REDUCE program for the calculation of geometrical characteristics of compactified multidimensional Riemannian space. Comp. Phys. Comm., 38:441–448, 1985.

Demichev:86

[DR86a] A. P. Demichev and A. Ya. Rodionov. Freund-Rubin type solutions for different compactifications of the eleven-dimensional space. Preprint 86-85, Institute for High Energy Physics, 1986.

Duncan:86

[DR86b] Anthony Duncan and Ralph Roskies. Representations of unusual mathematical structures in scientific applications of symbolic computation. J. Symbolic Computation, 2(2):201– 206, June 1986.

Darbaidze:89

[DR89] Ya. Z. Darbaidze and V. A. Rostovtsev. Analysis of the differential equations for the exclusive processes and explanation for the “mystery” of the gammadistribution. Preprint E2-89-286, JINR, 1989.

deRop:88

[dRD88] Y. de Rop and J. Demaret. Using EXCALC to study nondiagonal multidimensional spatially homogeneous cosmologies. Gen. Rel. Grav., 20:1127–1139, 1988.

dosSantos:87a

[dS87] R. P. dos Santos. O Método de Supercampos para o Cálculo de Potencial Efetivo em Modelos com Supercampos Quirais: Os Modelos de Wess e Zumino e de O’Raifeartaigh. PhD thesis, Centro Brasileiro de Pesquisas Físicas, 1987.

dosSantos:88a

[dS88] Renato P. dos Santos. Introdução ao sistema REDUCE de cálculo algébrico. Technical Report CBPF-NT-001/88, CBPF, Rio de Janeiro, Brazil, 1988.

dosSantos:87

[dS89] Renato P. dos Santos. Using REDUCE in supersymmetry. J. Symb. Comp., 7:523–525, 1989.

dosSantos:85

[dSS85] R. P. dos Santos and P. P. Srivastava. Two-loop effective potential for Wess–Zumino model using superfields. Technical Report IC/85/205, International Centre for Theoretical Physics, October 1985.

Davenport:88a

[DST88] J. H. Davenport, Y. Siret, and E. Tournier. Computer Algebra, Systems and Algorithms for Algebraic Computation. Academic Press, 1988.

Dufner:69

[DT69] A. M. Dufner and Y. S. Tsai. Phenomenological analysis of the γnn* form factors. Phys. Rev., 168:1801–1809, 1969.

Della-Dora:81

[DT81] J. Della Dora and E. Tournier. Solutions formelles D’Equations differentielles au voisinage de points singuliers reguliers. Report 239, Centre National de la Recherche Scientifique, 1981.

Della-Dora:84

[DT84] J. Della Dora and E. Tournier. Homogeneous linear difference equation (Frobenius-Boole Method). In Proc. EUROSAM 1984, Lecture Notes in Computer Science, volume 174, pages 2–12. Springer-Verlag, 1984.

Dulyan:87

[Dul87] L. S. Dulyan. The calculation of QCD triangular Feynman graphs in the external gluonic field using REDUCE-2 system. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 172– 173. Springer-Verlag, 1987.

Duval:87

[Duv87] Dominique Duval. Diverses questions relatives au Calcul Formel Avec des Nombres Algébriques. PhD thesis, L’Université Scientifique, Technologique et Médicale de Grenoble, 1987.

Earles:70

[Ear70] D. Earles. A measurement of the electron-production of muon pairs. Phys. Rev. Lett., 25:129– 133, 1970.

Eastwood:1991:OVR

[Eas91] James W. Eastwood. OR-THOVEC: version 2 of the RE-DUCE program for 3-D vector analysis in orthogonal curvilinear coordinates. Computer Physics Communications, 64(1): 121–122, April 1, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/001046559190054O.

Eastwood:1993:ACR

[Eas93] J. W. Eastwood. Book review: Algebraic computing with Reduce: M. MacCallum and F. Wright. Oxford University Press, 1991. £15. ISBN 0-19-853443-4. Computer Physics Communications, 76(1):140, June 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/001046559390127X.

Elishakoff:87a

[EC87] Isaac Elishakoff and Brian Couch. Application of symbolic algebra to the instability of a nonconservative system. J. Symbolic Computation, 4(3): 391–396, December 1987.

Edelen:81

[Ede81] Dominic G. B. Edelen. Programs for calculation of isovector fields in the REDUCE-2 environment. Technical Report TBD, Center for the Application of Mathematics, Lehigh University, August 1981.

Edelen:82

[Ede82] D. G. B. Edelen. Isovector fields for problems in the mechanics of solids and fluids. Int. Journ. Eng. Sci., 20:803–815, 1982.

Eliseev:85

[EFK85] V. P. Eliseev, R. N. Fedorova, and V. V. Kornyak. A REDUCE program for determining point and contact Lie symmetries of differential equations. Comp. Phys. Comm., 36:383–389, 1985.

Eissfeller:86

[EH86] Bernd Eißfeller and Günter W. Hein. A contribution to 3D-Operational geodesy. Technical Report Heft 17, Universitärer Studiengang Vermessungswesen and Universität der Bundeswehr München, December 1986.

Elishakoff:87

[EH87] Isaac Elishakoff and Joseph Hollkamp. Computerized symbolic solution for a nonconservative system in which instability occurs by flutter in one range of a parameter and by divergence in another. Comp. Methods in Applied Mechanics and Engineering, 62:27–46, 1987.

Eisenberger:90

[Eis90] Moshe Eisenberger. Application of symbolic algebra to the analysis of plates on variable elastic foundation. J. Symbolic Computation, 9(2):207–213, February 1990.

Eitelbach:73

[Eit73] D. L. Eitelbach. Automatic Analysis of Problems in Elementary Mechanics. PhD thesis, University of Illinois, 1973.

Edneral:89

[EKR89] Viktor F. Edneral, Aleksandr P.

Kryukov, and Anatolii Ia. Rodionov. The language of the analytic computer program REDUCE. Moscow, Izd-vo, Moskovskogo un-ta, 1989.

Eleuterio:82

[EM82] S. M. Eleutério and R. V.

Mendes. Note on equivalence and singularities: An application of computer algebra. Journ. Comp. Phys., 48:150–156, 1982.

Esteban:90

[ER90] E. P. Esteban and E. Ramos.

Algebraic computing and the Newman–Penrose formalism. Computers in Physics, pages 285–290, May/June 1990.

Leler:85

[eS85] Wm eler and Neil Soiffer. An interactive graphical interface for REDUCE. SIGSAM Bulletin, 19 (3):17–23, August 1985.

ESPRIT:1988:EPT

[ESP88] ESPRIT ’88: putting the technology to use: proceedings of the 5th Annual ESPRIT Conference, Brussels, November 14– 17, 1988. North-Holland, Amsterdam, The Netherlands, 1988. ISBN 0-444-87145-4. LCCN QA75.5.E84 1988. Two volumes. Publication no. EUR 11852 of the Commission of the European Communities, Directorate-General Telecommunications, Information Industries, and Innovation, Luxembourg.

Foster:89

[FB89] Kenneth R. Foster and Haim H. Bau. Symbolic manipulation programs for the personal computer. Science, 243:679–243, February 1989.

Flatau:86

[FBC86] Piotr J. Flatau, John P. Boyd, and William R. Cotton. Symbolic algebra in applied mathematics and geophysical fluid dynamics - REDUCE examples. Technical report, Dept. of Atmospheric and Oceanic Science, University of Michigan, and Dept. of Atmospheric Science, Colorado State University,

Cohen:79

[FC79] J. P. Fitch and H. I. Cohen. Using CAMAL for algebraic calculations in general relativity. General Relativity and Gravitation, 11:411–418, 1979.

Fazio:84

[FC84] P. M. Fazio and G. E. Copeland.

Cooper-type minima in multipole cross sections of atomic hydrogen. Phys. Rev. Lett., 53(2), July 1984.

Fuzio:85

[FC85] P. M. Fuzio and G. E. Copeland.

Partial radiative-recombination cross sections for excited states of hydrogen. Phys. Rev. A, 31 (1):187–195, 1985.

Feuillebois:84

[Feu84] F. Feuillebois. Sedimentation in a dispersion with vertical inhomogenieties. Journ. Fluid Mech., 139:145–171, 1984.

Fogelholm:82

[FF82a] Rabbe Fogelholm and Inge B.

Frick. Standard LISP for the VAX: A provisional implementation. SIGSAM Bulletin, 16(4): 10–12, November 1982.

Frick:82

[FF82b] I. G. Frick and R. Fogelholm.

An implementation of Standard Lisp built on top of Franz Lisp. Report, University of Stockholm, Institute of Physics, April 1982.

Fujimoto:84

[FG84] Y. Fujimoto and T. Garavaglia. Phase diagrams in Scalar QED. Physics Letters, 148B(1,2,3): 220–224, November 1984.

Fedorova:87

[FGGS87] R. N. Fedorova, V. P. Gerdt, N. N. Govorun, and V. P. Shirikov. Computer algebra in physical research of Joint Institute for Nuclear Research. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 1–10. Springer-Verlag, 1987.

Falck:89

[FGK89] N. K. Falck, D. Graudenz, and G. Kramer. Cross section for five-parton production in e^{+}e^{−} annihilation. Comp. Phys. Comm., 56(2):181–198, December 1989.

Fleischer:73

[FGM73] J. Fleischer, J. L. Gammel, and M. T. Menzel. Matrix Padé approximants for the 1SO- and 3PO- partial waves in nucleonnucleon scattering. Phys. Rev. D, 8:1545–1552, 1973.

Freire:89

[FGP89] E. Freire, E. Gamero, and E. Ponce. An algorithm for symbolic computation of Hopf bifurcation. In E. Kaltofen and S. M. Watt, editors, Proc. Computers and Mathematics ’89, pages 109–118. Springer-Verlag, New York, 1989.

Freire:88

[FGPF88] E. Freire, E. Gamero, E. Ponce, and L. G. Franquelo. An algorithm for symbolic computation of center manifolds. In Proc. of ISSAC ’88, volume 358, pages 218–230. Springer-Verlag, 1988.

Fox:74

[FH74] John A. Fox and Anthony C.

Hearn. Analytic computation of some integrals in fourth order quantum electrodynamics. Journ. Comp. Phys., 14:301– 317, 1974.

Fitch:87a

[FH87] J. P. Fitch and R. G. Hall. Symbolic computation and the finite element method. In Proc. EU-ROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 95–96. Springer-Verlag, 1987.

Fitch:73

[Fit73] John Fitch. Problems #3 and #4 in REDUCE and MAC-SYMA. SIGSAM Bulletin, pages 10–11, 1973.

Fitch:81

[Fit81] J. P. Fitch. User-based integration software. In Proc. 1981 ACM Symposium on Symbolic and Algebraic Computation, pages 245–248, 1981.

Fitch:83

[Fit83] J. P. Fitch. Implementing RE-DUCE on a microprocessor. In Proc. EUROCAL 1983, Lecture Notes in Computer Science, volume 162, pages 128– 136. Springer-Verlag, 1983.

Fitch:85a

[Fit85a] J. P. Fitch. Applying computer algebra. In International Conference on Computer Algebra and its Application in Theory, pages 262–275, 1985.

Fitch:85

[Fit85b] J. P. Fitch. Solving algebraic problems with REDUCE. J. of Symbolic Computation, 1(2): 211–227, June 1985.

Fitch:87

[Fit87] J. P. Fitch. Utilisation du calcul formel. In P. Chenin, editor, Calcul Formel et Automatique, pages 119–136. Editions du CNRS, 1987.

Fitch:89a

[Fit89a] J. Fitch. Compiling for parallelism. Computer Algebra and Parallelism, pages 19–31, 1989.

Fitch:89

[Fit89b] J. P. Fitch. Can REDUCE be run in parallel? In Proc. of IS-SAC ’89, pages 155–162. ACM Press, New York, 1989.

Fitch90

[Fit90] J. P. Fitch. A delivery system for REDUCE. In S. Watanabe and Morio Nagata, editors, Proceedings of the International Symposium on Symbolic and Algebraic Computation, pages 76–81. ACM, Addison-Wesley, 1990.

Feldmar:86

[FK86] E. Feldmar and K. S. Kölbig.

REDUCE procedures for the manipulation of generalized power series. Comp. Phys. Comm., 39:267–284, 1986.

Fedorova:87a

[FK87] R. N. Fedorova and V. V. Kornyak. Computer algebra application for determining local symmetries of differential equations. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 174– 175. Springer-Verlag, 1987.

Flath:86

[Fla86] Dan Flath. Remarks on tensor operators. Research Report 266, National University of Singapore, Department of Mathematics, July 1986.

Fleischer:71

[Fle71] J. Fleischer. Partial wave analysis of nucleon-nucleon Bethe-Salpeter equation on the computer. Journ. of Comp. Phys., 12:112–123, 1971.

Fox:71

[Fox71] J. A. Fox. Recalculation of the crossed graph contribution to the 4th order Lamb shift. Phys. Rev. D, 3:3228–3230, 1971.

Franceschetti:85

[FP85] G. Franceschetti and I. Pinto.

Nonlinear propagation and scattering: Analytical solution and symbolic code implementation. J. Opt. Soc. Am. A, 2:997–1006, 1985.

Fleischer:75

[FT75] J. Fleischer and J. A. Tjon. Bethe-Salpeter equation for J=0 nucleon-nucleon scattering with one-boson exchange. Nuclear Physics, B84:375–396, 1975.

Garavaglia:80

[Gar80] T. Garavaglia. A covariant formulation for polarized electron (muon) scattering on spin-zero and polarized spin-\frac{1}{2} targets. Il Nuovo Cimento, 56A:121–128, 1980.

Garavaglia:84

[Gar84] Theodore Garavaglia. Diracand Majorana-neutrino-mass effects in neutrino-electron elastic scattering. Physical Review D, 29(3):387–392, February 1984.

Garavaglia

[Garxx] Theodore Garavaglia. Polarized electron scattering on spin zero and polarized spin \frac{1}{2} targets: Deep inelastic scattering, elastic electron-muon scattering, and elastic electron-nucleon scattering. Preprint, Inst. Teich. Bhaile Atha Cliath, Eire, 19xx.

Gates:85a

[Gat85a] Barbara L. Gates. Gentran: An automatic code generation facility for REDUCE. SIGSAM Bulletin, 19(3):24–42, August 1985.

Gates:85c

[Gat85b] Barbara L. Gates. Gentran design and implementation, RE-DUCE version. Memorandum INF-85-12, Twente University of Technology, Department of Computer Science, The Netherlands, August 1985.

Gates:85b

[Gat85c] Barbara L. Gates. Gentran user’s manual - REDUCE version. Memorandum INF-85-11, Twente University of Technology, Department of Computer Science, The Netherlands, June 1985.

Gates:86

[Gat86] Barbara L. Gates. A numerical code generation facility for RE-DUCE. In Proc. SYMSAC ’86, pages 94–99, July 1986.

Gatermann:90

[Gat90a]  Karin Gatermann.  Grup-

pentheoretische konstruktion von symmetrischen kubaturformeln. Preprint TR 90-1, Konrad-Zuse-Zentrum für Informationstechnik Berlin, January 1990.

Gatermann90

[Gat90b] Karin Gatermann. Symbolic solution of polynomial equation systems with symmetry. In S. Watanabe and Morio Nagata, editors, Proceedings of the International Symposium on Symbolic and Algebraic Computation, pages 112–119. ACM, Addison-Wesley, 1990.

Gunion:72

[GBB72] J. F. Gunion, S. J. Brodsky, and R. Blankenbecler. Composite theory of large angle scattering and new tests of parton concepts. Phys. Lett., 39B:649–653, 1972.

Gunion:73

[GBB73] J. F. Gunion, S. J. Brodsky, and R. Blankenbecler. Large angle scattering and the interchange force. Report SLAC-PUB-1183, SLAC, 1973.

Grammaticos:85

[GDRH85] B. Grammaticos, B. Dorizzi, A. Ramani, and J. Hietarinta. Extending integrable Hamiltonian systems from 2 to N dimensions. Phys. Lett., 109A:81–84, 1985.

George:68

[Geo68] D. J. George. A covariant theory of the disintegration of the deuteron by pions and photons at high energy. Phys. Rev., 167: 1357–1364, 1968.

Gerdt:80

[Ger80a] V. P. Gerdt. Analytical calculations in high energy physics by computer. Comp. Phys. Comm., 20:85–90, 1980.

Gerdt:80b

[Ger80b] V. P. Gerdt. On global structure of the general solution of the Chew-Low equations. Preprint P2-80-436, J.I.N.R., Dubna, 1980.

Gaemers:78

[GG78] K. J. F. Gaemers and G. J.

Gounaris. Polarization amplitudes for e+e- → w+w- → zz. Preprint TH.2548-CERN, CERN, August 1978.

Gaemers

[GGRxx] K. J. F. Gaemers, R. Gastmans, and F. M. Renard. Neutrino counting in e+ e- collisions. Preprint, NIKHEF-H, Amsterdam, 19xx.

Griss:79

[GH79] Martin L. Griss and Anthony C.

Hearn. Portable LISP compiler. Software - Practice and Experience, 11:541–605, 1979.

Ucoluk:82

[GH82] G. Üçoluk and A. Hacinliyan. A proposal for extensions to RE-DUCE. SIGSAM Bulletin, 16 (2):4–14, May 1982.

Griss:78a

[GK78] Martin L. Griss and Robert R.

Kessler. REDUCE/1700: A micro-coded algebra system. In Proc. Micro, IEEE, volume 11, pages 130–138. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1978.

Griss:79a

[GK79] Martin L. Griss and Robert R.

Kessler. A micro-programmed implementation of Standard LISP and REDUCE on the Burroughs B1700/B1800 computer. Report, University of Utah, February 1979.

Grimm

[GK80] R. Grimm and H. Kühnelt. Using REDUCE in problems of supersymmetry and supergravity. Comp. Phys. Comm., 20:77, 1980.

Gunion:85

[GK85] J. F. Gunion and Z. Kunszt. Improved analytic techniques for tree graph calculations and the ggqq̄ll̄ subprocess. Phys. Lett., 161B:333–340, 1985.

Gerdt:89

[GK89] V. P. Gerdt and N. A. Kostov. Computer algebra in the theory of ordinary differential equations of halphen type. In E. Kaltofen and S. M. Watt, editors, Proc. Computers and Mathematics ’89, pages 279– 288. Springer-Verlag, New York, 1989.

Gerdt:87a

[GKK87] V. P. Gerdt, N. A. Kostov, and Z. T. Kostova. Computer algebra and computation of Puiseux expansions of algebraic functions. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 206– 207. Springer-Verlag, 1987.

Gerdt:85b

[GKR85] V. P. Gerdt, N. A. Kostov, and P. P. Raychev. Calculation of the matrix elements of the Hamiltonian of the interacting vector boson model using computer algebra - matrix elements of the Hamiltonian and some U(6)-Clebsch–Gordan coefficients. Technical Report E4-85-263, Institute for Nuclear Research and Nuclear Energy, Bulgarian Academy of Sciences, Sofia, Bulgaria, 1985.

Gerdt:85a

[GKRR85a] V. P. Gerdt, N. A. Kostov, P. P. Raychev, and R. P. Roussev. Calculation of the matrix elements of the Hamiltonian of the interacting vector boson model using computer algebra - basic concepts of the interacting vector boson model and matrix elements of the SU(3)- Quadrupole operator. Technical Report E4-85-262, Institute for Nuclear Research and Nuclear Energy, Bulgarian Academy of Sciences, Sofia, Bulgaria, 1985.

Gerdt:85c

[GKRR85b] V. P. Gerdt, N. A. Kostov, P. P.

Raychev, and R. P. Roussev. Calculation of the matrix elements of the Hamiltonian of the interacting vector boson model using computer algebra - matrix elements of the Hamiltonian - analytical results. Technical Report E4-85-264, Institute for Nuclear Research and Nuclear Energy, Bulgarian Academy of Sciences, Sofia, Bulgaria, 1985.

Gerdt90a

[GKZ90] Vladimar P. Gerdt, Nikolai V.

Khutornoy, and Alexey Yu. Zharkov. Solving algebraic systems which arise as necessary integrability conditions for polynomial-nonlinear evolution equations. In S. Watanabe and Morio Nagata, editors, Proceedings of the International Symposium on Symbolic and Algebraic Computation, page 299. ACM, Addison-Wesley, 1990.

Gladkih:83

[GL83] I. Gladkih and E. Lovas. On the application of computer algebra languages in the Central Research Institute for Physics. In Proceedings of the International Conference on Systems and Techniques of Analytical Computing and Their Applications in Theoretical Physics, D11-83-511, Dubna, 1983.

Ganzha:89

[GL89] V. Ganzha and R. Liska. Application of the REDUCE computer algebra system to stability analysis of difference schemes. In E. Kaltofen and S. M. Watt, editors, Proc. Computers and Mathematics ’89, pages 119– 129. Springer-Verlag, New York, 1989.

Gladd:82

[Gla82] N. T. Gladd. Computational aspects of research on the relativistic whistler instability. Technical Report J530-82-020, Jaycor, June 1982.

Good:75

[GLB75] D. Good, R. L. London, and W. W. Bledsoe. An interactive program verification system. Sigplan Notices, 10(6):482– 492, 1975.

Gould:84

[GM84] H. W. Gould and M. E. Mays.

Series expansions of means. Journ. of Mathematical Analysis and Applications, 101(2): 611–621, July 1984.

Gebauer:85

[GM85] Rüdiger Gebauer and H. Michael Möller. A fast variant of Buchberger’s algorithm. Technical report, Universität Heidelberg and Fernuniversität Hagen, October 1985.

Gebauer:88

[GM88] Rüdiger Gebauer and H. Michael Möller. On an installation of Buchberger’s algorithm. J. Symbolic Computation, 6(2 and 3): 275–286, 1988.

Gerdt:86

[GMS86] V. P. Gerdt, M. G. Meshcheryakov, and D. V. Shirkov. Computers in theoretical physics. Technical Report P2-86-848, J.I.N.R., Dubna, 1986.

Ganzha90a

[GMS90] V. G. Ganzha, S. V. Meleshko, and V. P. Shelest. Application of REDUCE system for analyzing consistency of systems of P.D.E.’s. In S. Watanabe and Morio Nagata, editors, Proceedings of the International Symposium on Symbolic and Algebraic Computation, page 301. ACM, Addison-Wesley, 1990.

Gervois:74

[GP74] A Gervois and Y. Pomeau. Logarithmic divergence in the virial expansion of transport coefficients of hard spheres. Phys. Rev. A, 9:2196–2213, 1974.

Golley

[GP82] Bruce W. Golley and Joseph Petrolito. An alternative finite strip technique for the static analysis of single-span, multispan and continuous plates. In Proc. International Conference on Finite Element Methods, 1982.

Garrad:86

[GQ86] A. D. Garrad and D. C. Quarton. Symbolic computing as a tool in wind turbine dynamics. Journ. of Sound and Vibration, 109(1):65–78, 1986.

Gragert:81

[Gra81] Peter Gragert. Symbolic Computations in Prolongation Theory. PhD thesis, Twente University of Technology, The Netherlands, 1981.

Greenland:84

[Gre84] P. T. Greenland. Comparison between phase diffusion and random telegraph signal models of laser bandwidth. Journ. Phys. B, 17:1919–1925, 1984.

Griss:74

[Gri74] M. L. Griss. The algebraic solution of large sparse systems of linear equations using REDUCE 2. In Proc. ACM 74, pages 105– 111, 1974.

Griss:75

[Gri75] Martin L. Griss. The REDUCE system for computer algebra. In Proc. ACM 75, pages 261–262, 1975.

Griss:74a

[Gri76a] M. L. Griss. The algebraic solution of sparse linear systems via minor expansion. ACM TOMS 2, pages 31–49, 1976.

Griss:76

[Gri76b] Martin L. Griss. The definition and use of data-structures in REDUCE. In Proc. SYMSAC 76, pages 53–59, 1976.

Griss:76a

[Gri76c] Martin L. Griss. An efficient sparse minor expansion algorithm. In Proc. ACM 76, pages 429–434, 1976.

Griss:77a

[Gri77a] M. L. Griss. Efficient recursive minor expansion. ACM TOMS, 1977.

Griss:77

[Gri77b] Martin L. Griss. Efficient expression evaluation in sparse minor expansion, using hashing and deferred evaluation. In Proc. 10th Hawaii International Conference on Systems Sciences, Western Periodicals, Calif., pages 169–172, 1977.

Griss:1978:UES

[Gri78] Martin L. Griss. Using an efficient sparse minor expansion algorithm to compute polynomial subresultants and the greatest common denominator. IEEE Transactions on Computers, C-27(10):945–950, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1674974.

Grozin:88

[Gro88a] A. G. Grozin. Solving physical problems with REDUCE. 1. REDUCE language 2. Classical nonlinear oscillator. Preprint 88-115, Institute of Nuclear Physics 630090, Novosibirsk, USSR, 1988.

Grozin:88a

[Gro88b] A. G. Grozin. Solving physical problems with REDUCE. 3. Nonlinear water waves 4. Calculation of the curvature tensor 5. Angular momentum addition. Preprint 88-136, Institute of Nuclear Physics 630090, Novosibirsk, USSR, 1988.

Grozin:88b

[Gro88c] A. G. Grozin. Solving physical problems with REDUCE. 6. Quantum nonlinear oscillator 7. Rotator in a weak field 8. Radiative transitions in charmonium. Preprint 88-140, Institute of Nuclear Physics 630090, Novosibirsk, USSR, 1988.

Goto:77

[GS77] E. Goto and T. Soma. MOL (moving objective lens) formulation of deflective aberration free system. Optik, 48:255–270, 1977.

Goto:78

[GS78] E. Goto and T. Soma. Electron beam lithography for advanced LSI fabrication. In Proc. 1978 National Computer Conference, AFIPS Press, New Jersey, pages 1223–1228, 1978.

Ganzha90

[GS90] Victor G. Ganzha and Michail Yu.

Shaskov. Local approximation study of difference operators by means of REDUCE system. In S. Watanabe and Morio Nagata, editors, Proceedings of the International Symposium on Symbolic and Algebraic Computation, pages 185–192. ACM, Addison-Wesley, 1990.

Gerdt:87

[GSSZ87] V. P. Gerdt, A. B. Shabat, S. I. Svinolupov, and A. Yu. Zharkov. Computer algebra application for investigating integrability of nonlinear evolution systems. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 81– 92. Springer-Verlag, 1987.

Gerdt:85

[GSZ85] V. P. Gerdt, A. B. Shvachka, and A. Yu. Zharkov. Computer algebra application for classification of integrable non-linear evolution equations. J. Symb. Comp., 1:101–107, 1985.

Gerdt:80a

[GTS80] V. P. Gerdt, O. V. Tarasov, and D. V. Shirkov. Analytical calculations on digital computers for applications in physics and mathematics. Sov. Phys. USP, 23:59–77, 1980.

Grammaticos:78

[GV78] B. Grammaticos and A. Voros. Semi-classical approximations for nuclear Hamiltonians I. Spinindependent potentials. Preprint DPh-T/78-75, CEN, Saclay, August 1978.

Garcia:86

[GV86] Arnaldo Garcia and Paulo Viana. Weierstrass points on certain non-classical curves. Arch. Math., 46:315–322, 1986.

Grammaticos

[GVxx] B. Grammaticos and A. Voros. Semi-Classical Approximations for Nuclear Hamiltonians: II. Spin-dependent Potentials. 19xx.

Gates:85

[GvH85] Barbara L. Gates and J. A. van Hulzen. Automatic generation of optimized programs. In Proc. EUROCAL ’85, April 1985.

Goldman:89

[GvH89] V. V. Goldman and J. A. van Hulzen. Automatic code vectorization of arithmetic expressions by bottom-up structure recognition. Computer Algebra and Parallelism, pages 119–132, 1989.

Gastmans:79

[GvPV79] R. Gastmans, A. van Proeyen, and P. Verbaeten. Symbolic evaluations of dimensionally regularized Feynman diagrams. Comp. Phys. Comm., 18:201– 203, 1979.

Gladkih:84

[GZ84] I. Gladkih and M. Zimanyi.

Comparison of systems for symbolic computing in use in the Central Research Institute for Physics (in Russian). In Proceedings of the International Conference on Computer-Based Scientific Research, Plovdiv, 1984.

Gerdt:90b

[GZ90a] V. P. Gerdt and A. Yu. Zharkov.

Computer classification of integrable coupled KdV-Like systems. J. Symb. Comp., 10:203– 207, 1990.

Gerdt90

[GZ90b] V. P. Gerdt and A. Yu. Zharkov.

Computer generation of necessary integrability conditions for polynomial-nonlinear evolution systems. In S. Watanabe and Morio Nagata, editors, Proceedings of the International Symposium on Symbolic and Algebraic Computation, pages 250–254. ACM, Addison-Wesley, 1990.

Handy:87

[Han87] N. C. Handy. The derivation of vibration-rotation kinetic energy operators, in internal coordinates. Mol. Phys., 61:207–223, 1987.

Harper:89a

[Hap89] David Haprer. Vector33: A RE-DUCE program for vector algebra and calculus in orthogonal curvilinear coordinates. Comp. Phys. Comm., 54(2 and 3):295– 305, June and July 1989.

Harrington:77

[Har77] Steven J. Harrington. A symbolic limit evaluation program in REDUCE. 1977.

Harrington:77a

[Har78] S. J. Harrington. REDUCE solution to problem #8. SIGSAM Bulletin, 11 and 12(4 and 1):7– 8, November and February 1977 and 1978.

Harrington:1979:NSI

[Har79a]  Steven J. Harrington.  A

new symbolic integration system in REDUCE. The Computer Journal, 22(2):127–131, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/127.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_
                                        02/tiff/128.tif;   http:/
                                        /www3.oup.co.uk/computer_
                                        journal/hdb/Volume_22/Issue_

02/tiff/129.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/130.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
02/tiff/131.tif.

Harrington:79a

[Har79b] Steven J. Harrington. A symbolic limit evaluation program in REDUCE. SIGSAM Bulletin, 13 (1):27–31, February 1979.

Harper:87

[Har87] David Harper. Dynamics of the Outer Satellites of Saturn. PhD thesis, Univ. of Liverpool, England, 1987.

Hartley:1997:ERP

[Har97] David Hartley. EDS: a RE-DUCE package for exterior differential systems. Computer Physics Communications, 100 (1–2):177–194, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0010465596001385.

Hearn:71

[Hea71a] Anthony C. Hearn. Calculation of traces of products of gamma matrices. In Proc. of the Second Colloquium on Advanced Computing Methods in Theoretical Physics, CNRS, Marseilles, pages I–30–I–44, 1971.

Hearn:71a

[Hea71b] Anthony C. Hearn. The computer solution of algebraic problems by pattern matching. In Proc. of the Second Colloquium on Advanced Computing Methods in Theoretical Physics, CNRS, Marseilles, pages I–45– I–57, 1971.

Hearn:72

[Hea72a] Anthony C. Hearn. Computer solution of symbolic problems in theoretical physics. In Computing as a Language of Physics, IAEA, Vienna, pages 567–596, 1972.

Hearn:72a

[Hea72b] Anthony C. Hearn. Improved non-modular polynomial GCD algorithm. SIGSAM Bulletin, pages 10–15, 1972.

Hearn:72b

[Hea72c] Anthony C. Hearn. A REDUCE solution of problem #2 - the Y(2n) functions. SIGSAM Bulletin, 14, 1972.

Hearn:73a

[Hea73] Anthony C. Hearn. The RE-DUCE program for computer algebra. In Proc. of the Third Colloquium on Advanced Computing Methods in Theoretical Physics, CNRS, Marseilles, 1973.

Hearn:74a

[Hea74a] Anthony C. Hearn. A mode analyzing algebraic manipulation program. In Proc. ACM 74, pages 722–724, 1974.

Hearn:74

[Hea74b] Anthony C. Hearn. Polynomial and rational function representations. In Proc. Math Software II, Purdue University, 1974.

Hearn:76a

[Hea76a] A. C. Hearn. A new REDUCE model for algebraic simplification. In Proc. SYMSAC 76, ACM, pages 46–52, 1976.

Hearn:76b

[Hea76b] A. C. Hearn. Symbolic computation. In Proc. CERN 1976 Computing School, CERN Geneva, pages 201–211, 1976.

Hearn:76

[Hea76c] Anthony C. Hearn. Scientific applications of symbolic computation. Computer Science and Scientific Comp., pages 83–108, 1976.

Hearn:77

[Hea77] A. C. Hearn. The structure of algebraic computations. In Proc. of the Fourth Colloquium on Advanced Comp. Methods in Theor. Physics. St. Maximin, France, pages 1–15, 1977.

Hearn:78

[Hea78] Anthony C. Hearn. Algebraic manipulation by computer. In Proc. Intern. Meeting on Programm. and Math. Meth. for Solving Phys. Probs., Dubna, USSR, pages 96–116, 1978.

Hearn:79

[Hea79]  Anthony C. Hearn.  Non-

modular computation of polynomial GCDs using trial division. In Proc. EUROSAM 79, volume 72, pages 227–239, 1979.

Hearn:80

[Hea80] Anthony C. Hearn. The personal algebra machine. In Information Processing 80, Proc. IFIP Congress 80, pages 621– 628, 1980.

Hearn:81a

[Hea81] Anthony C. Hearn. Symbolic computation and its application to high-energy physics. In Proc. 1980 CERN School of Computing, Geneva, pages 390–406, 1981.

Hearn:82

[Hea82] Anthony C. Hearn. REDUCE - a case study in algebra system development. In Proc. of EUROCAM ’82, Lecture Notes on Comp. Science, volume 144, pages 263–272, 1982.

Hearn:85

[Hea85] Anthony C. Hearn. Structure: The key to improved algebraic computation. In Proc. of the Second RIKEN International Symposium on Symbolic and Algebraic Computation by Computers, pages 215– 230. World Scientific, Singapore, 1985.

Hearn:86

[Hea86] Anthony C. Hearn. Optimal evaluation of algebraic expressions. In Proc. of AAECC-3, Lecture Notes on Comp. Science, volume 229, pages 392– 403, 1986.

Hearn:87

[Hea87] Anthony C. Hearn. REDUCE user’s manual, Version 3.3. Report CP 78, The RAND Corporation, July 1987.

Hermann:83

[Her83] R. Hermann. Geometric construction and properties of some families of solutions of nonlinear partial differential equations. J. Math. Phys., 24(3):510–521, 1983.

Hess:84

[HG84] P. O. Hess and W. Greiner.

The collective modes of nuclear molecules. Il Nuovo Cimento, 83A:76–177, 1984.

Hietarinta:89

[HG89] J. Hietarinta and B. Grammaticos. On the h̄^{2}-correction terms in quantum integrability. J. Phys. A: Mat. Gen., TBD:TBD, 1989.

Hearn:82a

[HGB82] Anthony C. Hearn, M. L. Griss, and E. Benson. Current status of a portable LISP compiler. In Proc. SIGPLAN ’82 Symp. on Compiler Construction, ACM, pages 276–283, 1982.

Hietarinta:84b

[HGDR84] J. Hietarinta, B. Grammaticos, B. Dorizzi, and A. Ramani. Coupling-constant metamorphosis and duality between integrable Hamiltonian systems. Phys. Rev. Lett., 53:1707–1710, 1984.

Hasenfratz:80

[HH80] Anna Hasenfratz and Peter Hasenfratz. The connection between the parameters of lattice and continuum QCD. Phys. Lett., 93B(1,2):165–169, June 1980.

Hirota:89

[HI89] Ryogo Hirota and Masaaki Ito. Introduction to REDUCE — Doing Symbolic Computation on PC. Science sha, Tokyo, June 1989.

Hietarinta:83a

[Hie83a] J. Hietarinta. Integrable families of Henon-Heiles type Hamiltonians and a new duality. Phys. Rev. A, 28:3670–3672, 1983.

Hietarinta:83

[Hie83b] J. Hietarinta. A search for integrable two-dimensional Hamiltonian systems with polynomial potential. Phys. Lett., 96A:273– 278, 1983.

Hietarinta:84

[Hie84a] J. Hietarinta. Classical versus quantum integrability. J. Math. Phys., 25:1833–1840, 1984.

Hietarinta:84a

[Hie84b] J. Hietarinta. New integrable Hamiltonians with transcendental invariants. Phys. Rev. Lett., 52:1057–1060, 1984.

Hietarinta:85

[Hie85] J. Hietarinta. How to construct integrable Fokker–Planck and electromagnetic Hamiltonians from ordinary integrable Hamiltonians. J. Math. Phys., 26:1970–1975, 1985.

Hietarinta:87

[Hie87a] J. Hietarinta. Direct methods for the search of the second invariant. Physics Reports, 147: 87–154, 1987.

Hietarinta:87a

[Hie87b] J. Hietarinta. A search of bilinear equations passing Hirota’s three- soliton condition: I. KdVtype bilinear equations. J. Math. Phys., 28:1732–1742, 1987.

Hietarinta:87b

[Hie87c] J. Hietarinta. A search of bilinear equations passing Hirota’s three- soliton condition: II. mKdV-type bilinear equations. J. Math. Phys., 28:2094– 2101, 1987.

Hietarinta:87c

[Hie87d] J. Hietarinta. A search of bilinear equations passing Hirota’s three- soliton condition: III. Sine-Gordon-type bilinear equations. J. Math. Phys., 28:2586– 2592, 1987.

Hietarinta:88

[Hie88] J. Hietarinta. A search of bilinear equations passing Hirota’s three- soliton condition: IV. complex bilinear equations. J. Math. Phys., 29:628–635, 1988.

Hearn:69

[HKY69] A. C. Hearn, P. K. Kuo, and D. R. Yennie. Radiative corrections to an electron-positron scattering experiment. Phys. Rev., 187:2088–2096, 1969.

Hearn:73

[HL73] Anthony C. Hearn and Rüdiger G. K. Loos. Extended polynomial algorithms. In Proc. ACM 73, pages 147–152, 1973.

Horowitz:75

[HM75] E. Horowitz and D. R. Musser. The synthesis and use of algebraic specifications of data structures. Preprint, University of Southern California, 1975.

Hearn:79a

[HN79] Anthony C. Hearn and Arthur C. Norman. A one-pass prettyprinter. Sigplan Notices, ACM 12, 14:50–58, 1979.

Horwitz:83

[Hor83] B. Horwitz. Unequal diameters and their effects on time varying voltages in branched neurons. BioPhys. J., 41:51–66, 1983.

Hadinger:87

[HT87] G. Hadinger and Y. S. Tergimen. Recurrence relations for the Dunham coefficients and analytic expressions of the diagonal radial matrix elements for an anharmonic oscillator. Journ. Chem. Phys., 87(4):2143–2150, 1987.

Husberg:81

[Hus81] N. Husberg. Preliminary II REDUCE-2 and Analitik-74, a comparison. Technical report, Helsinki University of Technology Computing Center, November 1981.

Hettich:77

[HvH77] R. P. Hettich and J. A. van Hulzen. Approximation with a class of rational functions. Memorandum 165, Department of Applied Mathematics, Twente University of Technology, The Netherlands, May 1977.

Hulshof:83

[HvH83] B. J. A. Hulshof and J. A. van Hulzen. Some REDUCE facilities for pretty printing subscripts and formal derivatives. SIGSAM Bulletin, 17(1):16–20, February 1983.

Hulshof:84

[HvH84] B. J. A. Hulshof and J. A. van Hulzen. Automatic error cumulation control. In Proc. EU-ROSAM 1984, Lecture Notes in Computer Science, volume 174, pages 260–271. Springer-Verlag, 1984.

Hulshof:85

[HvH85] B. J. A. Hulshof and J. A. van Hulzen. An expression compression package for REDUCE based on factorization and controlled expansion. In Proc. EU-ROCAL 1985, Lecture Notes in Computer Science, volume 204, pages 315–316. Springer-Verlag, 1985.

Hulshof:81

[HvHS81] B. J. A. Hulshof, J. A. van Hulzen, and J. Smit. Code optimization facilities applied in the Netform context. Memorandum 368, Department of Applied Mathematics, Twente University of Technology, The Netherlands, December 1981.

Hearn:81

[HW81] Anthony C. Hearn and S. Watanabe. Analytic integration by computer. Information Processing Society of Japan 22, pages 639–650, 1981.

Harper:89

[HWH89] David Harper, Chris Wooff, and David Hodgkinson. A guide to computer algebra systems. Report, Computer Laboratory, The University of Liverpool, Liverpool, England, September 1989.

Ito:90

[IC90] Nobuyasu Ito and Tetsuhiko Chikyu. Multi-spin-flip dynamics of the Ising chain. Physica A, 166:193–205, 1990.

Ito:85a

[IK85] M. Ito and F. Kako. A RE-DUCE program for finding conserved densities of partial differential equations with uniform rank. Comp. Phys. Comm., 38: 415–419, 1985.

Ilyin:87

[IK87] V. A. Ilyin and A. P. Kryukov.

DIMREG — the package for calculations in the dimensional regularization with 4-dimensional γ^{5}-matrix in quantum field theory. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 225– 232. Springer-Verlag, 1987.

Ilyin:1996:ARP

[IK96] V. A. Ilyin and A. P. Kryukov.

ATENSOR — REDUCE program for tensor simplification. Computer Physics Communications, 96(1):36–52, July 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0010465596000604.

Ilyin:89

[IKRT89] V. A. Ilyin, A. P. Kryukov, A. Ya. Rodioniov, and A. Yu. Taranov. Fast algorithm for calculation of Dirac’s Gamma-Matrices traces. SIGSAM Bulletin, 23(4):15–24, October 1989.

Inada:80

[Ina80] Nobuyuki Inada. Fortran-based LISP system for REDUCE. Technical report, Information Science Laboratory, The Institute of Physical and Chemical Research, 1980.

Ito:85

[Ito85] M. Ito. A REDUCE program for evaluating a Lax pair form. Comp. Phys. Comm., 34:325– 331, 1985.

Ito:88

[Ito88] Masaaki Ito. A REDUCE program for Hirota’s bilinear operator and Wronskian operations. Comp. Phys. Comm., 50(3):321– 330, August 1988.

Ito:90a

[Ito90] Nobuyasu Ito. Discrete-time and single-spin-flip dynamics of the Ising chain. Progress of Theoretical Physics, 83(4):682–692, April 1990.

Ito:1994:SRP

[Ito94] Masaaki Ito. SYMCD — a REDUCE package for finding symmetries and conserved densities of systems of nonlinear evolution equations. Computer Physics Communications, 79 (3):547–554, May 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0010465594901937.

Idesawa:77

[IY77] M. Idesawa and T. Yatagai. General theory of projectiontype moiré topography. Scientific Papers 71, Institute of Physical and Chemical Research, Wako-Shi, Saitama,

Jeffrey:84

[JO84] D. J. Jeffrey and Y. Onishi. The forces and couples acting on two nearly touching spheres in lowreynolds-number flow. Z. Ang. Math. Phys., 35:634–641, 1984.

Janssen:87

[JPS87] M. H. M. Janssen, D. H. Parker, and S. Stolte. Saturation in laser-induced fluorescence: Effects on alignment parameters. Chemical Phys., 113:357–382, 1987.

Jansen:86

[JW86] Paul Jansen and Peter Weidner. High-accuracy arithmetic software–some tests of the ACRITH problem-solving routines. ACM TOMS, 12(1):62–70, March 1986.

Kadlecsik:88

[Kad88] J. Kadlecsik. New approaches to the axisymmetric vacuum. Zeitschrift für Physik C. Particles and Fields, 41:265–269, 1988.

Kadlecsik:1996:RCP

[Kad96] József Kadlecsik. Ricci calculus package in REDUCE. Computer Physics Communications, 93(2–3):265–282, February 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0010465595001379.

Kagan:88

[Kag88] Y. Y. Kagan. Static sources of elastic deformation in a homogeneous half-space. J. Geophys. Res., 93(B9):10,560– 10,574, September 1988.

Kahn:69

[Kah69]  M. E. Kahn.  The near-

minimum-time control of open loop articulated kinematic chains. Report AIM-106, Stanford University, Computer Science Dept., 1969.

Kamel:69

[Kam69a] A. A. Kamel. Perturbation method in the theory of nonlinear oscillations. Report, Stanford University, Dept. of Aeronautics and Astronautics, 1969.

Kamel:69a

[Kam69b] A. A. Kamel. Perturbation theory based on Lie transforms and its application to the stability of motion near Sun-perturbed Earth-Moon triangular libration points. Report 391, Stanford University, Dept. of Aeronautics and Astronautics, 1969.

Kamel:78

[Kam78] A. A. Kamel. Synchronous satellite ephemeris due to earth’s triaxiality and luni-solar effects. In AIAA/AAS Astrodynamics Conference, Palo Alto, CA, August 1978.

Kanada:75

[Kan75] Y. Kanada. Implementation of HLISP and algebraic manipulation language REDUCE 2. Report 75-01, University of Tokyo Information Science Lab, 1975.

Karr:85

[Kar85] Michael Karr. Canonical form for rational exponential expressions. In Proc. EUROCAL 1985, Lecture Notes in Computer Science, volume 204, pages 585– 594. Springer-Verlag, 1985.

Katsura:85

[Kat85] Shigetoshi Katsura. Application of the formula manipulating system to statistical mechanics. In Proc. of the Second RIKEN International Symposium on Symbolic and Algebraic Computation by Computers, pages 155– 180. World Scientific, Singapore, 1985.

Kauffman:73

[Kau73]  S. K. Kauffman.  Ortho-

Positronium Annihilation: Steps Toward Computing the First Order Radiative Corrections. PhD thesis, California Institute of Technology, Pasadena, CA, USA, 1973.

Kazasov:87

[Kaz87] C. Kazasov. Laplace transformations in REDUCE 3. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 132–133. Springer-Verlag, 1987.

Kinoshita:72

[KC72] T. Kinoshita and P. Cvitanovic.

Sixth order radiative corrections to the electron magnetic moment. Report CLNS-197, Cornell Lab. for Nuclear Studies, October 1972.

Kinoshita:73

[KC73] T. Kinoshita and P. Cvitanovic. Feynman-Dyson rules in parametric space. Report CLNS-209, Cornell Lab. for Nuclear Studies, January 1973.

Kersten:86a

[KE86a] P. H. M. Kersten and H. M. M. Ten Eikelder. Infinite hierarchies of t-independent and t-dependent conserved functionals of the Federbush model. J. Math. Phys., 27:2140–2145, 1986.

Kersten:86b

[KE86b] P. H. M. Kersten and H. M. M. Ten Eikelder. An infinite number of infinite hierarchies of conserved quantities of the Federbush model. J. Math. Phys., 27:2791–2796, 1986.

Keady:85

[Kea85] Grant Keady. The power concavity of solutions of some semilinear elliptic Boundary-Value problems. Bull. Austral. Math. Soc., 31:181–184, 1985.

Keener:83

[Kee83] James P. Keener. Oscillatory coexistence in the chemostat: a codimension two unfolding. SIAM J. Appl. Math., 43(5): 1005–1018, 1983.

Keener:85

[Kee85]  James P. Keener.  Oscilla-

tory coexistence in a food chain model with competing predators. J. Math. Biology, 22:123– 135, 1985.

Keener:89

[Kee89] James P. Keener. Knotted scroll wave filaments in excitable media. Physica D 34, pages 378– 390, 1989.

Keener:90

[Kee90] James P. Keener. Knotted vortex filaments in an ideal fluid. J. Fluid Mech., 211:629–651, 1990.

Kendall:88

[Ken88] W. S. Kendall. Symbolic computation and the diffusion of shapes of triads. Adv. Appl. Prob., 20:775–797, 1988.

Kendall:89

[Ken89a] W. S. Kendall. The diffusion of Euclidean shape. Research Report 161, University of Warwick, Dept. of Statistics, 1989.

Kendall:89a

[Ken89b] W. S. Kendall. Probability, convexity, and harmonic maps with small image i: Uniqueness and fine existence. Research Report 162, University of Warwick, Dept. of Statistics, 1989.

Kersten:83

[Ker83] P. H. M. Kersten. Infinitesimal symmetries and conserved currents for nonlinear Dirac equation. J. Math. Phys., 24:2374– 2376, 1983.

Kersten:86

[Ker86] P. H. M. Kersten. Creating and annihilating Lie–Bäcklund transformations of the Federbush model. J. Math. Phys., 27: 1139–1144, 1986.

Kornyak:87

[KF87] V. V. Kornyak and R. N. Fedorova. A REDUCE program to calculate determining equations of Lie-Baecklund symmetries of differential equations. Technical Report P11-87-19, J.I.N.R., Dubna, 1987.

Kerner:75

[KG75] W. Kerner and R. C. Grimm. MHD spectra for Tokamaks with non-circular cross sections. In Proc. Seventh Conference on Numerical Simulation of Plasmas, Courant Institute, NYU, 1975.

Kagan:85

[KK85] Y. Y. Kagan and L. Knopoff. The first-order statistical moment of the seismic moment tensor. Geophys. J. R. Astron. Soc., 81:429–444, 1985.

Kaneko:89

[KK89] Toshiaki Kaneko and Setsuya Kawabata. A preprocessor for fortran source code produced by REDUCE. Comp. Phys. Comm., 55(2):141–147, September 1989.

Kamal:81

[KKM81] A. N. Kamal, J. Kodaira, and T. Muta. Gluon jets from heavy paraquarkonium. Technical Report SLAC-PUB-2725, University of Alberta, Canada and Stanford University, California and Fermi National Accelerator Laboratory, Illinois, April 1981.

Koh:82

[KKP^{+}82] I. G. Koh, Y. D. Kim, Y. J. Park, C. H. Kim, and Y. S. Kim. Complete set of SU(5) monopole solution. J. Math. Phys., 23:1210– 1212, 1982.

Killalea:80

[KM80] M. K. Killalea and B. J. Mc-Coy. Concentration distribution and spatial moments of moving macromolecules undergoing isomerization. Biopolymers, 19: 1875–1886, 1980.

Kersten:84

[KM84] P. Kersten and R. Martini. The harmonic map and killing fields for self-dual SU(3) Yang–Mills equations. J. Phys. A, 17:L227– L230, 1984.

Kobayashi:88

[KMH88] H. Kobayashi, S. Moritsugu, and R. W. Hogan. Solving systems of algebraic equations. In Proc. of ISSAC ’88, volume 358, pages 139–149. Springer-Verlag, 1988.

Kitatani:86

[KMS86] H. Kitatani, S. Miyashita, and M. Suzuki. Reentrant phenomena in some Ising spin systems - rigorous results and effects of an external field. J. Phys. S. Japan, 55(3):865–876, 1986.

Kruse:83

[KO83] Hans-Guenther Kruse and Karin Ohlsen. About the realization of an extended, but really interactive REDUCE by integration of a small editing and executing system. SIGSAM Bulletin, 17(1):21–25, February 1983.

Kobayashi:84

[Kob84] Hidestune Kobayashi. Weierstrass points on a curve, X_{0}^{7} + X_{1}^{7} + X_{2}^{7} = 0. Preprint 28, Research Institute of Science and Technology, Nihon University, March 1984.

Koelbig:81b

[Köl81] K. S. Kölbig. A program for computing the conical functions of the first kind P_{−1/2+iτ}^{m} (x) for m = 0 and m = 1. Comp. Phys. Comm., 23:51–61, 1981.

Koelbig:82

[Köl82] K. S. Kölbig. Closed expressions for ∫_{0}^{1} t^{−1} log^{n−1t} log^{p}(1 − t)dt. Math. Comp., 39(160):647–654, October 1982.

Koelbig:83a

[Köl83a] K. S. Kölbig. On the integral ∫_{0}^{∞} e^{−μt}t^{ν−1} log^{m}tdt. Math. Comp., 41:171–182, 1983.

Koelbig:83

[Köl83b] K. S. Kölbig. On the integral ∫_{0}^{π/2} log^{n} cos x log^{p} sin x dx.

Math. Comp., 40:565–570, April 1983.

Koelbig:84b

[Köl84] K. S. Kölbig. Some problems involving special functions arising from physics at CERN. Technical Report DD 84-14, CERN, Data Handling Division, September 1984.

Koelbig:85a

[Köl85a] K. S. Kölbig. Explicit evaluation of certain definite integrals involving powers of logarithms. J. Symbolic Computation, 1(1): 109–114, March 1985.

Koelbig:85

[Köl85b] K. S. Kölbig. On the integral ∫_{0}^{1} x^{ν−1}(1 − x)^{−λ} ln^{m}xdx. Technical Report DD/85/18, CERN, Data Handling Division, September 1985.

Koelbig:86

[Köl86] K. S. Kölbig. On the integral ∫_{0}^{∞} x^{ν−1}(1+βx)^{−λ} ln^{m}xdx. Journal of Comp. and Appl. Math., 14:319–344, 1986.

Kotorynski:86

[Kot86] W. P. Kotorynski. Steady laminar flow through a twisted pipe of elliptical cross-section. Computers and Fluids, 14:433–444, 1986.

Kaps:85

[KPB85] P. Kaps, S. W. H. Poon, and T. D. Bui. Rosenbrock methods for stiff ODEs: A comparison of Richardson extrapolation and embedding techniques. Computing, 34:17–40, 1985.

Kuppers:71

[KPT71] G. Kuppers, D. Pfirsch, and H. Tasso. M.H.D. - stability of axisymmetric plasmas. Report CN -28/F-14, Max-Planck-Institut fuer Plasmaphysik, 1971.

Koelbig:82a

[KR82] K. S. Kölbig and W. Rühl. Complex zeros of the partition function for two-dimensional U(N) lattice gauge theories. Z. Phys. C - Particles and Fields, 12:135– 143, 1982.

Kryukov:85

[KR85a] A. P. Kryukov and A. Ya. Rodionov. Dynamic-debugging system for the REDUCE programs. SIGSAM Bulletin, 19(2):34–37, May 1985.

Kryukov:85a

[KR85b] A. P. Kryukov and A. Ya. Rodionov. Interactive REDUCE. SIGSAM Bulletin, 19(3):43–45, August 1985.

Kryukov:87a

[KR87] A. P. Kryukov and A. Ya. Rodionov. CTS - algebraic debugging system for REDUCE programs. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 233– 243. Springer-Verlag, 1987.

Kryukov:88a

[KR88] A. P. Kryukov and A. Ya. Rodionov. Program “COLOR” for computing the group-theoretic weight of Feynman diagrams in Non-Abelian gauge theories. Comp. Phys. Commun., 48(2): 327–334, February 1988.

Kryukov

[KRxx] A. P. Kryukov and A. Ya. Rodionov. Usage of REDUCE for computations of grouptheoretical weight of Feynman diagrams in non-abelian gauge theories. Technical report, Institute of Nuclear Physics, Moscow, USSR, 19xx.

Kraus:73

[Kra73] J. Kraus. Delbrückstreuung und Prüfung der Quantenelektrodynamik. PhD thesis, Ludwig-Maximilians-Universität zu München, 1973.

Krack:82

[Kra82] K. Krack. Rechnerunterstützte Entwicklung der Mittelbreitenformeln und abschätzung ihrer ellipsoidischen Anteile zur lösung der zweiten geodätischen Hauptaufgabe auf dem Rotationsellipsoid. Z. Vermessungswes., 107:502–513, 1982.

Kredel:88

[Kre88] Heinz Kredel. Admissible termorderings used in computer algebra systems. SIGSAM Bulletin, 22(1):28–31, January 1988.

Kryukov:87

[KRR87] A. P. Kryukov, A. Ya. Rodionov, and V. A. Rostovtsev. Pattern compilation in RE-DUCE. Technical Report P11-87-302, J.I.N.R., Dubna, 1987.

Kryukov:88

[KRR88] A. P. Kryukov, A. Ya. Rodionov, and V. A. Rostovtsev. New programming tools for computing substitution rules in REDUCE system. Technical Report P11-88-402, J.I.N.R., Dubna, 1988.

Kryukov:84

[Kry84] A. P. Kryukov. An antitranslator of the RLISP language. SIG-SAM Bulletin, 18(3):12–15, August 1984.

Kanada:81

[KS81a] Yasumasa Kanada and Tateaki Sasaki. LISP-based big-float system is not slow. SIGSAM Bulletin, 15(2):13–19, May 1981.

Koelbig:81

[KS81b] K. S. Kölbig and F. Schwarz. On positive function series. Computing, 27:319–337, 1981.

Koelbig:84

[KS84a] K. S. Kölbig and B. Schorr. Asymptotic expansions for the Landau density and distribution function. Comp. Phys. Comm., 32:121–131, 1984.

Koelbig:84a

[KS84b] K. S. Kölbig and B. Schorr. A program package for the Landau distribution. Comp. Phys. Comm., 31:97–111, 1984.

Kryukov:88b

[KS88] A. P. Kryukov and D. A.

Slavnov. The role of the gg → c\overline{c}g process in the cross section of production of charmed particles (in Russian). Preprint 88-49/70, Moscow State University, 1988.

Kodaira:85

[KT85] Hiroshi Kodaira and Hiroshi Toshima. Gini coefficient of wealth in life cycle model. In Proc. of the Second RIKEN International Symposium on Symbolic and Algebraic Computation by Computers, pages 119– 151. World Scientific, Singapore, 1985.

Liska:87

[LD87] R. Liska and D. Drska. Evaluation of plasma fluid equations collision integrals using REDUCE. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, page 178. Springer-Verlag, 1987.

Liska90

[LD90] R. Liska and L. Drska. FIDE: A REDUCE package for automation of FInite difference method for solving pDE. In S. Watanabe and Morio Nagata, editors, Proceedings of the International Symposium on Symbolic and Algebraic Computation, pages 169–176. ACM, Addison-Wesley,

Lottati

[LExx] Itzhak Lottati and Isaac Elishakoff. Refined dynamical theories of beams, plates and shells and their applications. In Proc. Euromech-Colloquium 219, 19xx.

Levi:71

[Lev71] I. M. Levi. Symbolic algebra by computer - applications to structural mechanics. In AIAA/ASME 12th Structures, Structural Dynamics and Materials Conference, Anaheim, California, April 1971.

Levi:70

[LH70] I. Levi and N. Hoff. Nonsymmetric creep buckling of circular cylindrical shells in axial compression. In Proc. Intern. Symp. in Creep Effect in Structures, Gotenburg, Sweden, August 1970.

Liebermann:75

[Lie75] R. Liebermann. Traces of high energy processes in strong magnetic fields. J. Comp. Phys., 1975.

Liska:84

[Lis84] R. Liska. Program for stability and accuracy analysis of finite difference methods. Comp. Phys. Comm., 34:175–186, 1984.

Luegger:73

[LM73] J. Luegger and H. Melenk. Darstellung und Bearbeitung umfangreicher LISP-Programme.

Angewandte Informatik, pages 257–263, June 1973.

London:74

[LM74] R. London and D. R. Musser.

The application of a symbolic mathematical system to program verification. In Proc. ACM 74, pages 265–273, 1974.

Lee:85

[LM85] H-C Lee and M. S. Milgram.

On the axial gauge: Ward identities and the separation of infrared and ultraviolet singularities by analytical regularization. J. Math. Phys., 26:1793–1804, 1985.

Lepage:83

[LMSZ83] G. P. Lepage, P. B. Mackenzie, K. H. Streng, and P. M. Zernas. Multiphoton decays of positronium. Phys. Rev. A, 28:3090– 3091, 1983.

Loe:85

[LOG85] Kia Fock Loe, Noritaka Ohsawa, and Eiichi Goto. Circuit simulation code generation by computer algebra. In Proc. of the Second RIKEN International Symposium on Symbolic and Algebraic Computation by Computers, pages 87–103. World Scientific, Singapore, 1985.

Loos:72

[Loo72] Rüdiger Loos. Analytic treatment of three similar Fredholm integral equations. SIGSAM Bulletin, 11:32–40, 1972.

Lang:79

[LP79] C. B. Lang and W. Porod. Symmetry breaking and π K amplitudes in the unphysical region. Report UNIGRAZ-UTP 08/79, Institut für Theor. Physik, Univ. Graz, 1979.

Lloyd:90

[LP90] N. G. Lloyd and J. M. Pearson. REDUCE and the bifurcation of limit cycles. J. Symbolic Computation, 9(2):215–224, February 1990.

Lukacs

[LPSV82] B. Lukács, Z. Perjés, A. Sebestyén, and A. Valentini. Stationary vacuum fields with a conformally flat three-space, ii. proof of axial symmetry. Technical Report KFKI-1982-19, Central Research Institute for Physics, Budapest, Hungary, 1982.

Laursen:79

[LS79] M. L. Laursen and M. A. Samuel. The n-bubble diagram contribution to the g-2 of the electron - Mathematical structure of the analytical expression. Research Note 96, Oklahoma State Univ. Quantum Theoretical Research Group, 1979.

Laursen:80

[LS80] Morten L. Laursen and Mark A. Samuel. Borel transform technique and the n-Bubble diagram contribution to the lepton anomaly. Research Note 10, Oklahoma State Univ. Quantum Theoretical Research Group, August 1980.

Laursen:81

[LS81] M. L. Laursen and M. A.

Samuel. The n-bubble diagram contribution to g-2. J. Maths. Phys., 22:1114–1126, 1981.

Louw:86

[LSS86] J. A. Louw, F. Schwarz, and W. H. Steeb. First integrals and Yoshida analysis of Nahm’s equation. J. Phys. A, 19:L569– L573, 1986.

Lukaszuk:87

[LSS87] L. Lúkaszuk, D. M. Siemienczuk, and L. Szymanowski. Evaluation of helicity amplitudes. Phys. Rev. D, 35:326–329, 1987.

Lux:75

[Lux75] Augustin Lux. Etude d’un Modele Abstrait pour une Machine LISP et de son Implantation. PhD thesis, Université Scientifique et Medicale de Grenoble, March 1975.

Lambin:84

[LV84] P. Lambin and J. P. Vigneron.

Computation of crystal Green’s functions in the complex-energy plane with the use of the analytical tetrahedron method. Phys. Rev. B, 29(6):3430–3437, 1984.

Mack:73

[Mac73]  D. Mack.  Nichtnumerische

Verfahren und deren Anwendung in der Elementarteilchen-Physik. PhD thesis, University of Tuebingen, 1973.

MacCallum:86a

[Mac86a] M. A. H. MacCallum. Algebraic computing in relativity. Technical Report TAU 86-04, Queen Mary College, University of London, 1986.

MacCallum:86

[Mac86b] M. A. H. MacCallum. Dynamical Spacetimes and Numerical Relativity. Cambridge UP, 1986.

MacCallum:87

[Mac87] M. A. H. MacCallum. Symbolic computation in relativity theory. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 34–43. Springer-Verlag, 1987.

MacCallum:88

[Mac88] M. A. H. MacCallum. An ordinary differential equation solver for REDUCE. In Proc. of ISSAC ’88, volume 358, pages 196–205. Springer-Verlag, 1988.

MacCallum:89

[Mac89a] Malcolm A. H. MacCallum. Comments on the performance of algebra systems in general relativity and a recent paper by Nielsen and Pedersen. SIGSAM Bulletin, 23(2):22–25, April 1989.

Maclaren:89

[Mac89b] N. M. Maclaren. The generation of sequences of multiple independent sequences of pseudorandom numbers. Applied Statistics JRSS Series C, 38(2):351– 359, 1989.

Maguire:81

[Mag81] Gerald Quentin Maguire Jr.

Program transformation in RE-DUCE using rule sequencing. Master’s thesis, Department of Computer Science, The University of Utah, March 1981.

Malm:82

[Mal82] Bengt Malm. A program in RE-DUCE for finding explicit solutions. In Proc. EUROCAM 1982, Lecture Notes in Computer Science, volume 144, pages 289–293. Springer-Verlag, 1982.

Marti:78

[Mar78] Jed Marti. The META/REDUCE translator writing system. Sigplan Notices, 13:42–49, 1978.

Marti:85a

[Mar85] Jed B. Marti. The role of explanation in symbolic computation. In Proc. of the Second RIKEN International Symposium on Symbolic and Algebraic Computation by Computers, pages 13–34. World Scientific, Singapore, 1985.

Marti:88

[Mar88] J. Marti. A graphics interface to REDUCE. In Proc. AAECC-6 1988, Lecture Notes in Computer Science, volume 357, pages 274–296. Springer-Verlag, 1988.

Mazzarella:85

[Maz85] Giuseppe Mazzarella. Improved simplification of odd and even functions in REDUCE. SIG-SAM Bulletin, 19(2):29–30, May 1985.

McCrea:87a

[MBG87] J. D. McCrea, P. Baekler, and M. Guerses. A Kerr-like solution of the Poincaré gauge field equations. Il Nuovo Cim, 99B: 171–177, 1987.

McCrea:81

[McC81] J. D. McCrea. The Petrov type of a static vacuum spacetime near a normal-dominated singularity. J. Phys., A14:1351–1356, 1981.

McCrea:82

[McC82] J. D. McCrea. A stationary cylindrically symmetric electrovac spacetime. J. Phys., A15: 1587–1590, 1982.

McCrea:83

[McC83] J. D. McCrea. Static, vacuum, cylindrical and plane symmetric solutions of the quadratic Poincaré gauge field equations. J. Phys., A16:997–1004, 1983.

McCrea:84

[McC84a] J. D. McCrea. A NUT-like solution of the quadratic-poincaré gauge field equations. Phys. Lett., 100A:397–399, 1984.

REDUCE:84

[McC84b] J. D. McCrea. The Use of RE-DUCE in Finding Exact Solutions of the Quadratic Poincaré Gauge Field Equations, in Classical General Relativity. Cambridge University, 1984. 173–182 pp.

McCrea:87

[McC87] J. D. McCrea. Poincaré gauge theory of gravitation: Foundations, exact solutions and computer algebra. In Differential Geometric Methods in Mathematical Physics, Proc. 14th International Conference, Salamanca, 1985 (Springer Lecture Notes in Mathematics, No. 1251), page 16, 1987.

McIsaac:85

[McI85] Kevin McIsaac. Pattern matching algebraic identities. SIG-SAM Bulletin, 19(2):4–13, May 1985.

Matveev:87

[MDMS87] V. A. Matveev, Ya. Z. Darbaidze, Z. V. Merebashvili, and L. A. Slepchenko. Gluon fusion in SUSY QCD. Phys. Lett. B, 191(1 and 2):179–181, June 1987.

Marti:83

[MF83] Jed Marti and John Fitch. RE-DUCE 2 for CP/M. SIGSAM Bulletin, 17(1):26–27, February 1983.

Moritsugu:88

[MG88] S. Moritsugu and E. Goto. A proposal for improvement of facilities of REDUCE. Technical report, Department of Information Science, University of Tokyo, Japan, December 1988.

Moritsugu:89

[MG89] Shuichi Moritsugu and Eiichi Goto. A note on the preconditioning for factorization of homogeneous polynomials. SIG-SAM Bulletin, 23(1):9–12, January 1989.

Marti:85

[MH85] Jed B. Marti and Anthony C. Hearn. REDUCE as a LISP benchmark. SIGSAM Bulletin, 19(3):8–16, August 1985.

Marti:79

[MHGG79] J. B. Marti, A. C. Hearn, M. L. Griss, and C. Griss. Standard Lisp report. Sigplan Notices, ACM, 14(10):48–68, 1979.

Marti:80

[MHGG80] J. Marti, A. C. Hearn, M. L. Griss, and C. Griss. Standard Lisp report. SIGSAM Bulletin, 14(1):23–41, February 1980.

Maurer:86

[MHK86] M. Maurer, A. Hayd, and H. J. Kaeppeler. Quasi-analytical method for solving nonlinear differential equations for turbulent self-confined magnetoplasma. J. Comp. Phys., 66:151– 172, 1986.

Moritsugu:85

[MIG85] S. Moritsugu, N. Inada, and E. Goto. Symbolic Newton iteration and its application. In Proc. of the Second RIKEN International Symposium on Symbolic and Algebraic Computation by Computers, pages 105– 117. World Scientific, Singapore, 1985.

Mack:73a

[MM73] D. Mack and H. Mitter. Calculation of electron-electron-bremsstrahlung cross-sections. Phys. Lett., 44A:71–72, 1973.

Moritsugu:89a

[MM89] Shuichi Moritsugu and Makoto Matsumoto. A note on the numerical evaluation of arctangent function. SIGSAM Bulletin, 23 (3):8–12, July 1989.

McCrea:88

[MMH88] J. D. McCrea, E. W. Mielke, and F. W. Hehl. A remark on the axisymmetric Chen et al. solution of the Poincaré gauge theory. Phys. Lett., 127A:65–69, 1988.

Mueller:81

[MMK81] R. Müller and H. J. W. Müller-Kirsten. Iteration of single- and two-channel schrödinger equations. J. Math. Phys., 22:733– 749, 1981.

Melenk:88

[MMN88] H. Melenk, H. M. Möller, and W. Neun. On gröbner bases computation on a supercomputer using REDUCE. Preprint SC 88-2, Konrad-Zuse-Zentrum für Informationstechnik Berlin, January 1988.

Melenk:89

[MMN89] H. Melenk, H. M. Möller, and W. Neun. Symbolic solution of large stationary chemical kinetics problems. Impact of Computing in Science and Engineering, 1(2):138–167, June 1989.

DeMenna:87

[MMR87] L. De Menna, G. Miano, and G. Rubinacci. Volterra’s series solutions of free boundary plasma equilibria. Phys. Fluids, 30:409–416, 1987.

Melenk:89a

[MN89a] Herbert Melenk and Winfried Neun. Implementation of portable standard LISP for the SPARC processor. Preprint SC 89-6, Konrad-Zuse-Zentrum für Informationstechnik Berlin, July 1989.

Melenk:89b

[MN89b] Herbert Melenk and Winfried Neun. Parallel polynomial operations in the large Buchberger algorithm. Computer Algebra and Parallelism, pages 143–158, 1989.

Neun:89a

[MN89c] Herbert Melenk and Winfried Neun. Parallel polynomial operations in the large buchberger algorithm. In J. Della-Dora and J. Fitch, editors, Computer Algebra and Parallelism, pages 143–158. Academic Press, 1989.

Moller:89

[Möl89] H. Michael Möller. Multivariate rational interpolation reconstruction of rational functions.

Preprint SC 89-4, Konrad-Zuse-Zentrum für Informationstechnik Berlin, July 1989.

Mirie:84

[MS84] R. M. Mirie and C. H. Su. Internal solitary waves and their head-on collision: Part I. J. Fluid Mechanics, 147:213–231, 1984.

Mazepa:85

[MS85] N. E. Mazepa and S. I.

Serdyukova. The stability investigation of some difference boundary problem with the application of symbolic computation system. Technical Report E5-85-39, J.I.N.R., Dubna, 1985.

Murzin:85

[Mur85] F. A. Murzin. Syntactic properties of the REFAL language. Int. J. Computer Maths., 17: 123–139, 1985.

Noor:79

[NA79] A. K. Noor and C. M. Andersen.

Computerized symbolic manipulation in structural mechanics - progress and potential. Computers and Structures, 10:95–118, 1979.

Namba:86

[Nam86] Kenji Namba. Some improvements on Utah Standard Lisp. SIGSAM Bulletin, 20(1 and 2): 29–36, February and May 1986.

Norman:79

[ND79] A. C. Norman and J. H. Davenport. Symbolic integration - the dust settles? In Proc. EU-ROSAM 1979, Lecture Notes in Computer Science, volume 72, pages 398–407. Springer-Verlag, 1979.

Ng:89

[Ng89]  Tze Beng Ng.  Computa-

tion of the cohomology of BŜO_{n} < 16 > for 23 ≤ n ≤ 26 using REDUCE. J. Symbolic Computation, 7(1):93–99, January 1989.

Nikityuk:87

[Nik87] N. M. Nikityuk. Some questions of using coding theory and analytical calculation methods on computers. Technical Report E11-87-10, J.I.N.R., Dubna, 1987.

Niki:84

[NK84] Naoto Niki and Sadanori Konishi. Higher order asymptotic expansions for the distribution of the sample correlation coefficient. Comm. Statist.-Simula. Comp., 13(2):169–182, 1984.

Norman:77

[NM77] A. C. Norman and P. M. A. Moore. Implementing the new Risch integration algorithm. In Proc. of the Fourth Colloquium on Advanced Comp. Methods in Theor. Phys., St. Maximin, France, March 1977.

Neun:88

[NM88] Winfried Neun and Herbert Melenk. Implementation of the LISP-arbitrary precision arithmetic for a vector processor. Preprint SC 88-1, Konrad-Zuse-Zentrum für Informationstechnik, Berlin, January 1988.

Nakamura:89

[NM89a] Hideharu Nakamura and Shouichi Matsui. Symbolic Computation in Structural Mechanics using REDUCE. Gihodo Shuppan Company Ltd., 1-11-41, Akasaka, Minato-Ku, 107 Tokyo, Japan, 1989.

Neun:89

[NM89b] W. Neun and H. Melenk. Implementation of the LISP- arbitrary precision arithmetic for a Vector processor. Computer Algebra and Parallelism, pages 75– 89, 1989.

Nakashima:84a

[NMJ84a] T. T. Nakashima, R. E. D. Mc-Clung, and B. K. John. Experimental and theoretical investigation of _{2D}−_{13C} dept spectra on CD_{N}. J. Magnetic Resonance, 58:27–36, 1984.

Nakashima:84

[NMJ84b] T. T. Nakashima, R. E. D. Mc-Clung, and B. K. John. A simple method for the determination of the deuterium decoupler pulse angle. J. Magnetic Resonance, 56:262–274, 1984.

Norman:78

[Nor78] Arthur Norman. Towards a REDUCE solution to SIGSAM problem 7. SIGSAM Bulletin, 12 (4):14–18, November 1978.

Norton:80

[Nor80] Lewis M. Norton. A note about Laplace transform tables for computer use. SIGSAM Bulletin, 14(2):30–31, May 1980.

Norman90

[Nor90] A. C. Norman. A criticalpair/completion based integration algorithm. In S. Watanabe and Morio Nagata, editors, Proceedings of the International Symposium on Symbolic and Algebraic Computation, pages 201–205. ACM, Addison-Wesley, 1990.

Nagata:82

[NS82] Morio Nagata and Makoto Shibayama. COSMOS: A conversational algebraic system. Technical Report No. 8201, Department of Administration Engineering, Keio University, March 1982.

Nagata:85

[NS85] Morio Nagata and Makoto Shibayama. An interactive algebraic system for personal computing. In IEEE International Symposium on New Directions in Computing, 1985.

Neutsch:86

[NS86] W. Neutsch and E. Schrüfer. Simple integrals for solving Kepler’s equation. Astrophysics and Space Science, 125:77–83, 1986.

Neutsch:85

[NSJ85] W. Neutsch, E. Schrüfer, and A. Jessner. Note on efficient integration on the hypersphere. J. Comp. Phys., 59:167–175, 1985.

Norman:83

[NW83] Arthur C. Norman and Paul S.

Wang. A comparison of the Vaxima and REDUCE. SIGSAM Bulletin, 17(1):28–30, February 1983.

Nemeth:82

[NZ82] G. Németh and M. Zimányi.

Polynomial type Padé approximants. Math. Comp., 38:553– 565, 1982.

Nemeth:87

[NZ87] G. Németh and M. Zimányi.

Computation of generalized Padé approximants. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 450–451. Springer-Verlag, 1987.

Ogilvie:82

[Ogi82] J. F. Ogilvie. Applications of computer algebra in physical chemistry. Computers in Chemistry, 6(4):169–172, 1982.

Ogilvie:89

[Ogi89] J. F. Ogilvie. Computer algebra in modern physics. Computers in Physics, pages 66–74, January/February 1989.

Ochiai:90

[ON90] Mitsuyuki Ochiai and Kiyokazu Nagatomo. Linear Algebra using REDUCE. Kindai Kagaku sha, Tokyo, January 1990.

Ono:1979

[Ono79] Kiyoshi Ono. BFORT – a Fortran system with arbitrary precision integer and real arithmetic. Technical report, Department of Physics, University of Tokyo, January 1979.

Ozieblo

[Ozixx] A. Ozieblo. Application of REDUCE 2 in general theory of relativity. Technical report, Cyfronet - Krakow, Poland, 19xx.

Piessens:86

[PA86] R. Piessens and S. Ahmed. Note on approximation for the turning points of Bessel functions. J. Comp. Phys., 64:253–257, 1986.

Parsons:68

[Par68] R. G. Parsons. An estimate of the sixth order contribution to the anomalous magnetic moment of the electron. Phys. Rev., 168:1562–1567, 1968.

Parsons:71

[Par71] R. G. Parsons. S-channel transformation matrices for helicity and invariant amplitudes for λ+ N to O + B. Memo CPT-88, Center for Particle Theory, University of Texas, January 1971.

Padget90

[PB90] Julian Padget and Alan Barnes. Univariate power series expansions in REDUCE. In S. Watanabe and Morio Nagata, editors, Proceedings of the International Symposium on Symbolic and Algebraic Computation, pages 82– 87. ACM, Addison-Wesley, 1990.

Perrottet:78

[Per78] M. Perrottet. Signature for W boson production from jet analysis in e+e- → W+W- → hadrons. Preprint 78/P.1019, CPT 2, CNRS, Marseille, June 1978.

Perjes:84b

[Per84a] Z. Perjés. Improved characterization of the Kerr metric. Technical Report KFKI-1984-115, Hungarian Academy of Sciences, Central Research Institute for Physics, 1984.

Perjes:84

[Per84b] Z. Perjés. Stationary vacuum fields with a conformally flat three-space. III. Complete solution. General Relativity and Gravitation, 18:531–547, 1984.

Perjes:84c

[Per84c] Z. Perjés. Stationary vacuum fields with a conformally flat three-space. IV. Complete solution. Technical Report INS-REP.-487, January 1984.

Perjes:86

[Per86a] Z. Perjés. Ernst coordinates.

Preprint KFKI-1986-33/B, Hungarian Academy of Sciences, Central Research Institute for Physics, 1986.

Perjes:86a

[Per86b] Z. Perjés. Stationary vacuum fields with a conformally flat three-space. II. Proof of axial symmetry. General Relativity and Gravitation, 18(5):511–530, May 1986.

Perjes:88

[Per88] Z. Perjés. Approaches to axisymmetry by man and machine. 1988.

Pesic:73

[Pes73] P. D. Pesic. Two-photon cross section for W-pair production by colliding beams. Report SLAC-PUB-1188, Stanford University, 1973.

Pattnaik:83

[PFF83] P. C. Pattnaik, G. Fletcher, and J. L. Fry. Improved numerical stability for norm-conserving ion-Ure pseudopotentials. Phys. Rev. B, 28(6):3364–3365, 1983.

Pearce:81

[PH81] P. D. Pearce and R. J. Hicks. The application of algebraic optimisation techniques to algebraic mode programs for RE-DUCE. SIGSAM Bulletin, 15 (4):15–22, November 1981.

Pearce:83

[PH83] P. D. Pearce and R. J. Hicks. Data structures and execution times of algebraic mode programs for REDUCE. SIGSAM Bulletin, 17(1):31–37, February 1983.

Pictiaw:69

[Pic69] Chen Pictiaw. An Analytical Investigation of Infinitesimal Spatial Motion Theory and its Application to Three-Dimensional Linkages. PhD thesis, Dept. of Mech. Eng., Stanford University, March 1969.

Piessens:84

[Pie84] R. Piessens. A series expansion for the first positive zero of the Bessel function. Math. Comp., 42:195–197, 1984.

Pignataro:85

[PLR85] M. Pignataro, A. Luongo, and N. Rizzi. On the effect of the local overall interaction on the postbuckling of uniformly compressed channels. Thin-Walled Structures, 3:292–321, 1985.

Perjes:84a

[PLSV84] Z. Perjés, B. Lukács, A. Sebestyén, and A. Valentini. Solution of the stationary vacuum equations of relativity for conformally flat 3-spaces. Phys. Lett., 100A(8): 405–406, February 1984.

Pankau:73a

[PN73a] E. Pankau and W. Nakel. Eine koinzidenzmessung zum elementarprozess der elektron-elektron-bremsstrahlung bei 300 kev. Z. Physik, 264:139–153, 1973.

Pankau:73

[PN73b] E. Pankau and W. Nakel. Measurement of the absolute cross section of the elementary process of electron-electron bremsstrahlung at 300 keV. Phys. Lett., 44A:65–67, 1973.

Podgorzak:84

[PR84] E. Podgórak and I. Romanowska. Application of RE-DUCE 2 to the construction of recurrence relations. Master’s thesis, Institute of Computer Science, University of Wroclaw, Wroclaw, Poland, 1984.

Perlt:90

[PRH90] H. Perlt, J. Ranft, and J. Heinrich. Calculation of Qed graphs with the Spinor technique. Comp. Phys. Commun., 56(3): 385–390, January 1990.

Price:84

[PSA84] S. L. Price, A. J. Stone, and M. Alderton. Explicit formulae for the electrostatic energy, forces and torques between a pair of molecules of arbitrary symmetry. Molecular Phys., 52: 987–1001, 1984.

Quarton:84

[QG84] D. C. Quarton and A. D. Garrad. Symbolic computing as a tool in wind turbine dynamics. Technical report, Wind Energy Group, Taylor Woodrow Construction Ltd., 1984.

Quarton

[QGxx] D. C. Quarton and A. D. Garrad. Some comments on the stability analysis of horizontal axis wind turbines. Technical report, Wind Energy Group, Taylor Woodrow Construction Ltd., 19xx.

Rogers:89

[RA89] C. Rogers and W. F. Ames.

Nonlinear Boundary Value Problems in Science and Engineering. Academic Press, Inc., 1989.

Rao:85

[Rao85] R. H. Rao. Deformation of a fluid-filled cylindrical membrane by a slow viscous shear flow. Master’s thesis, Washington University, Dept. of Mech. Eng., Washington University, St. Louis, MO, USA, 1985.

Rayna:87

[Ray87] G. Rayna. REDUCE: A System for Computer Algebra. Springer-Verlag, 1987.

Roque:88

[RdS88] Waldir L. Roque and Renato P. dos Santos. Computação algébrica: “um assistente matemãtico”. Ciência e Cultura, 40(9):843–852, September 1988.

Renner:1992:CRP

[Ren92] Friedrich Renner. A constructive REDUCE package based upon the Painlevé analysis of nonlinear evolutions equations in Hamiltonian and / or normal form. Computer Physics Communications, 70 (2):409–416, June 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/001046559290203B.

Rink:71

[Rin71] R. A. Rink. Application of a Digital Computer to Solve Analytically Special Classes of Linear and Nonlinear Differential Equations. PhD thesis, Stanford University, Stanford, CA, USA, 1971.

Reusch:86

[RN86] M. F. Reusch and G. H. Neilson. Torodially symmetric polynomial multipole solutions of the vector Laplace equation. J. Comp. Phys., 64:416–432, 1986.

Rodionov:84

[Rod84] A. Ya. Rodionov. Work with non-commutative variables in the REDUCE-2 system for analytical calculations. SIGSAM Bulletin, 18(3):16–19, August 1984.

Rizzi:85

[RT85] N. Rizzi and A. Tatone. Symbolic manipulation in buckling and postbuckling analysis. Computers and Structures, 21:691– 700, 1985.

Rodionov:87a

[RT87a] A. Ya. Rodionov and A. Yu. Taranov. Combinatorial aspects of simplification of algebraic expressions. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 192– 201. Springer-Verlag, 1987.

Rodionov:87

[RT87b] A. Ya. Rodionov and A. Yu.

Taranov. Computation of covariant derivatives of the geodetic interval within the coincident arguments. Class. Quantum Grav., 4:1767–1775, 1987.

Rodionov:88

[RT88a] A. Ya. Rodionov and A. Yu.

Taranov. RTENSOR - Packet for work with tensoric expressions. Preprint 88-29/50, Moscow State University, Scientific Research Institute of Nuclear Physics, 1988.

Ronveaux:88

[RT88b] A. Ronveaux and G. Thiry.

Polynomial solution of recurrence relation and differential equation. SIGSAM Bulletin, 22 (4):9–19, October 1988.

Ronveaux:89

[RT89] A. Ronveaux and G. Thiry. Differential equations of some orthogonal families in REDUCE. J. Symbolic Computation, 8(5): 537–541, November 1989.

Sage:88

[Sag88] Martin L. Sage. An algebraic treatment of quantum vibrations. J. Symbolic Computation, 5(3):377–384, June 1988.

Sasaki:79

[Sas79] Tateaki Sasaki. An arbitrary precision real arithmetic package in REDUCE. In Proc. EU-ROSAM 1979, Lecture Notes in Computer Science, volume 72, pages 358–368. Springer-Verlag, 1979.

Savage:90

[Sav90] Stuart B. Savage. Symbolic computation of the flow of granular avalanches. J. Symbolic Computation, 9(4):515–530, April 1990.

Sayers:87

[Say87a] C. M. Sayers. The elastic anisotropy of polycrystalline aggregates of zirconium and its alloys. J. Nuclear Materials, 144: 211–213, 1987.

Sayers:87a

[Say87b] C. M. Sayers. Elastic wave anisotropy in the upper mantle. Geophysical J. R. Ast. Soc., 88: 417–424, 1987.

Schmuck:77

[Sch77] P. Schmuck. Verification of the transient, two phase fluid flow program kachina using computerized similarity analysis. In Second GAMM Conference on Numerical Methods in Fluid Mechanics, köln, October 1977.

Schwarz:80

[Sch80] F. Schwarz. An approximation scheme for constructing π_{0}π amplitudes from ACU requirements. Fortschritte der Physik, 28:201–235, 1980.

Schruefer:82

[Sch82a] E. Schrüfer. An implementation of the exterior calculus in REDUCE: A status report.

SIGSAM Bulletin, 16(4):27–31, November 1982.

Schwarz:82a

[Sch82b] F. Schwarz. A REDUCE package for determining Lie symmetries of ordinary and partial differential equations. Computer Physics Communications, 27:179–186, 1982.

Schwarz:82

[Sch82c] F. Schwarz. Symmetries of the two dimensional Korteweg– de Vries equation. J. Phys. S. Japan, 51(8):2387–2388, 1982.

Schwarz:83a

[Sch83a] Fritz Schwarz. Automatically determining symmetries of ordinary differential equations. In Proc. EUROCAL 1983, Lecture Notes in Computer Science, volume 162, pages 45–54. Springer-Verlag, 1983.

Schwarz:83

[Sch83b] Fritz Schwarz. A REDUCE package for series analysis by Hadamard’s theorem and QD schemes. SIGSAM Bulletin, 17 (1):38–44, February 1983.

Schwarz:84

[Sch84] F. Schwarz. The Riquier-Janet theory and its application to nonlinear evolution equations. Physica, 11D:243–251, 1984.

Schwarz:85

[Sch85a] F. Schwarz. Automatically determining symmetries of partial differential equations. Computing, 34:91–106, 1985.

Schwarz:85a

[Sch85b] Fritz Schwarz. An algorithm for determining polynomial first integrals of autonomous systems of ordinary differential equations. J. Symbolic Computation, 1(2):229–233, June 1985.

Schwarz:86

[Sch86] F. Schwarz. A REDUCE package for determining first integrals of autonomous systems of ordinary differential equations. Computer Physics Communications, 39:285–296, 1986.

Schwarz:87

[Sch87] F. Schwarz. Symmetries and involution systems: Some experiments in computer algebra. In Topics in Soliton Theory and Exactly Solvable Nonlinear Equations. World Science Press, Singapore, August 1987.

Schruefer:88

[Sch88a] E. Schrüfer. A note on Einstein metrics. SIGSAM Bulletin, 22 (3):22–26, July 1988.

Schwarz:88

[Sch88b] F. Schwarz. Symmetries of differential equations: From Sophus Lie to computer algebra. Siam Review, 30:450–481, 1988.

Seiler:1991:SRP

[Sei91] Werner M. Seiler. SUPERCALC — a REDUCE package for commutator calculations. Computer Physics Communications, 66(2– 3):363–376, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/001046559190082V.

Smit:87

[SGM87] J. Smit, S. H. Gerez, and R. Mulder. Application of a structured LISP system to computer algebra. In Proc. EURO-CAL ’87, Lecture Notes in Computer Science, volume 378, pages 149–160. Springer-Verlag, 1987.

Stuart:90

[SGT90] Robin G. Stuart and A. Góngora-T. Algebraic reduction of one-loop Feynman diagrams to scalar integrals ii. Comp. Phys. Commun., 56(3):337–350, January 1990.

Soderstrand:72

[SH72] M. A. Soderstrand and D. C.

Huey. Sensitivities of fourthorder filters obtained by a lowpass to band-pass transformation. Report, University of California, Davis, 1972.

Stroscio:74

[SH74] M. A. Stroscio and J. M.

Holt. Radiative corrections to the decay rate of orthopositronium. Phys. Rev. A, 10:749–755, September 1974.

Schruefer:81

[SH81] E. Schrüfer and H. Heintzmann. Lorentz-covariant eikonal method in magnetohydrodynamics II - the determination of the wave amplitude. Phys. Lett., 81A(9):501–506, February 1981.

Shablygin:87

[Sha87] E. Shablygin. Integral equation with hidden Eigenparameter Solver: REDUCE and FOR-TRAN in tandem. In Proc. EU-ROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 186–191. Springer-Verlag, 1987.

Schruefer:87

[SHM87] E. Schrüfer, F. W. Hehl, and J. D. McCrea. Exterior calculus on the computer: The REDUCE-package EXCALC applied to general relativity and to the Poincaré gauge theory. General Relativity and Gravitation, 19(2):197–218, February 1987.

Shtokhamer:75

[Sht75] R. Shtokhamer. Canonical form of polynomials in the presence of side relations. Technical Report Technion-PH-76-25, Technion, 1975.

Shtokhamer:77

[Sht77] R. Shtokhamer. The use of “LET” statements in producing short comprehended outputs. Technical Report Technion-PH-77-36, Department of Physics, Technion-Israel Institute of Technology, Haifa, Israel, 1977.

Shmueli:83a

[SK83] U. Shmueli and U. Kaldor.

Moments of the trigonometric structure factor. Acta Cryst., A39:615–621, 1983.

Steuerwald

[SKxx] J. Steuerwald and W. Kerner. A contribution to the efficient solution of extensive symbolic computations. Comp. Phys. Comm., 19xx.

Soderstrand:74

[SL74] M. A. Soderstrand and J. F.

Lathrop. Two computer programs for the sensitivity analysis of higher order filters. Report SLL-73-0225, Sandia Laboratories, January 1974.

Soderstrand:72a

[SM72] M. A. Soderstrand and S. K.

Mitra. Computer-aided sensitivity analysis of higher filters. In Proc. Second Symposium on Network Theory, Herzegnovia, Yugoslavia, July 1972.

Saez:83

[SM83] A. E. Saez and B. J. McCoy.

Transient analysis of packed-bed thermal storage systems. Int. J. Heat Mass Transfer, 26(1):49– 54, 1983.

Smit:79

[Smi79] J. Smit. New recursive minor expansion algorithms, a presentation in a comparative context. In Proc. EUROSAM 1979, Lecture Notes in Computer Science, volume 72, pages 74–87. Springer-Verlag, 1979.

Soma:77

[Som77] T. Soma. Relativistic aberration formulas for combined electricmagnetic focusing-deflection system. Optik, 49:255–262, 1977.

Soma:85

[Som85] Takashi Soma. Recent applications of REDUCE in RIKEN. In Proc. of the Second RIKEN International Symposium on Symbolic and Algebraic Computation by Computers, pages 181– 182. World Scientific, Singapore, 1985.

Spiridonova:87

[Spi87] M. Spiridonova. Some extensions and applications of RE-DUCE system. In Proc. EURO-CAL ’87, Lecture Notes in Computer Science, volume 378, pages 136–137. Springer-Verlag, 1987.

Squire

[Squxx] W. Squire. Some applications of symbolic matrix inversion. Technical report, Dept. of Mechanical and Aerospace Engineering, West Virginia University, 19xx.

Schwarz:84a

[SS84] F. Schwarz and W. H. Steeb. Symmetries and first integrals for dissipative systems. J. Phys. A: Math. Gen., 17:L819–L823,

Suppes:89

[ST89a] Patrick Suppes and Shuzo Takahashi. An interactive calculus theorem-prover for continuity properties. J. Symbolic Computation, 7(6):573–590, June 1989.

Surguladze:89

[ST89b] L. R. Surguladze and F. V. Tkachov. LOOPS: Procedures for multiloop calculations in quantum field theory for the RE-DUCE system. Comp. Phys. Comm., 55(2):205–215, September 1989.

Steinberg:82

[Ste82] Stanly Steinberg. Mathematics and symbol manipulation. SIG-SAM Bulletin, 16(3):11–15, August 1982.

Stoutemyer:75

[Sto75] David R. Stoutemyer. Symbolic computer solution of an equation in finite terms. Report UCP-33, Dept. of Comp. Science, Univ. of Utah, 1975.

Stoutemyer:74

[Sto77a] D. Stoutemyer. Automatic error analysis using the computer symbolic manipulation language. TOMS 3, 3(1):26–43, March 1977.

Stoutemyer:77

[Sto77b] David R. Stoutemyer. Analytically solving integral equations by using computer algebra. TOMS, 3(2):128–146, June

Stuart:88

[Stu88] Robin G. Stuart. Algebraic reduction of one-loop Feynman diagrams to scalar integrals. Comp. Phys. Commun., 48(3): 367–389, March 1988.

Smit:82

[SvH82] J. Smit and J. A. van Hulzen. Symbolic numeric methods in microwave technology. In Proc. EUROCAM 1982, Lecture Notes in Computer Science, volume 144, pages 281–288. Springer-Verlag, 1982.

Smit:81

[SvHH81] J. Smit, J. A. van Hulzen, and B. J. A. Hulshof. NET-FORM and code optimizer manual. SIGSAM Bulletin, 15(4): 23–32, November 1981.

Shmueli:83

[SW83] U. Shmueli and A. J. C. Wilson. Generalized intensity studies: The subcentric distribution and effects of dispersion. Acta Cryst., A39:225–233, 1983.

Tallents:84

[Tal84] G. J. Tallents. The relative intensities of hydrogen-like fine structure. J. Phys. B, 17:3677– 3691, 1984.

Tao90

[Tao90]  Qingsheng Tao.  Symbolic

and algebraic manipulation for formulae of interpolation and quadrature. In S. Watanabe and Morio Nagata, editors, Proceedings of the International Symposium on Symbolic and Algebraic Computation, page 306. ACM, Addison-Wesley, 1990.

Todd:88

[TC88] P. H. Todd and G. W. Cherry.

Symbolic analysis of planar drawings. In Proc. of ISSAC ’88, volume 358, pages 344–355. Springer-Verlag, 1988.

Tsai:65

[TH65] Y. S. Tsai and A. C. Hearn. Differential cross-section for e+ + e- → W+ + W- → e- + \overline{ν}_{e} +μ+ ν_{μ}. Phys. Rev., 140:B721–B729, 1965.

Thas:89

[Tha89a] C. Thas. A collection of RE-DUCE and MACSYMA programs about college geometry. part 1. Technical Report 5, State University of Gent, September 1989.

Thas:89a

[Tha89b] C. Thas. A collection of RE-DUCE and MACSYMA programs about college geometry. part 2. Technical Report 5, State University of Gent, September 1989.

Tournier:79

[Tou79] Evelyne Tournier. An algebraic form of a solution of a system of linear differential equations with constant coefficients. In Proc. EUROSAM 1979, Lecture Notes in Computer Science, volume 72, pages 153–163. Springer-Verlag, 1979.

Tournier:87

[Tou87] Evelyne Tournier. Solutions Formelles D’Equations Differentielles, le Logiciel de Calcul Formel: DESIR Etude Theorique et Realisation. PhD thesis, L’Université Scientifique, Technologique et Medicale de Grenoble, April 1987.

Trotter:89

[Tro89] H. F. Trotter. Use of symbolic methods in analyzing an integral operator. In E. Kaltofen and S. M. Watt, editors, Proc. of Computers and Mathematics ’89, pages 82–90. Springer-Verlag, New York, 1989.

Tasso:76

[TS76] H. Tasso and J. Steuerwald. Subroutine for series solutions of linear differential equations. Technical Report IPP 6/143, Max Planck Institut for Plasmaphysik, 1976.

Tsai:74

[Tsa74] Y. S. Tsai. Pair production and bremsstrahlung of charged leptons. Rev. Mod. Phys., 46:815– 851, 1974.

Toth:86

[TSM86] K. Tóth, K. Szegő, and A. Margaritis. Radiative corrections for semileptonic decays of Hyperons: ‘Model-Independent’ Part. Physical Review D, 33(11):3306– 3315, June 1986.

Umeno:89

[UYSA89] Takaji Umeno, Syuichi Yamashita, Osami Saito, and Kenichi Abe. Symbolic computation application for the design of linear multivariable control systems. J. Symbolic Computation, 8(6):581–588, December 1989.

VanProeyen:1976:QGC

[Van76] A. Van Proeyen. Quantum gravity corrections on the anomalous magnetic and quadrupole moments of a spin-1 particle. Technical report, Instituut voor Theor. Fys., Leuven, October 1976.

VanProeyan:1979:GDE

[Van79] A. Van Proeyan. Gravitational divergences of the electromagnetic interactions of massive vectorparticles. Preprint KUL-TF-79/032, Universiteit Leuven, October 1979.

vandenHeuvel:86

[vdH86a] Pim van den Heuvel. Adding statements to REDUCE. SIG-SAM Bulletin, 20(1 and 2):8–14, February and May 1986.

vandenHeuvel:86a

[vdH86b] Pim van den Heuvel. Some experiments in REDUCE related to the calculation of Groebner bases. Technical report, Department of Computer Science, Twente University of Technology, The Netherlands, June

vandenHeuvel:87a

[vdHHvH87] P. van den Heuvel, B. J. A. Hulshof, and J. A. van Hulzen. Some simple Pretty-Print facilities for REDUCE. SIGSAM Bulletin, 21 (1):14–17, February 1987.

vandenHeuvel:87

[vdHvHG87] P. van den Heuvel, J. A. van Hulzen, and V. V. Goldman. Automatic generation of FORTRAN-coded Jacobians and Hessians. In Proc. EURO-CAL ’87, Lecture Notes in Computer Science, volume 378, pages 120–131. Springer-Verlag, 1987.

vanHulzen:80

[vH80] J. A. van Hulzen. Computational problems in producing Taylor coefficients for the rotating disk problem. SIGSAM Bulletin, 14(2):36–49, May 1980.

vanHulzen:81

[vH81] J. A. van Hulzen. Breuer’s grow factor algorithm in computer algebra. Memorandum 332, Department of Applied Mathematics, Twente University of Technology, The Netherlands, April 1981.

vanHulzen:82a

[vH82] J. A. van Hulzen. Computer algebra systems viewed by a notorious user. In Proc. EUROCAM 1982, Lecture Notes in Computer Science, volume 144, pages 166–180, 1982.

vanHulzen:83a

[vH83] J. A. van Hulzen. Code optimization of multivariate polynomial schemes: A pragmatic approach. In Proc. EUROCAL 1983, Lecture Notes in Computer Science, volume 162, pages 286–300. Springer-Verlag, 1983.

vanHulzen:87

[vH87] J. A. van Hulzen. Program generation aspects of the symbolicnumeric interface. In Proc. Third Intern. Conf. on Computer Algebra and its applications in Theor. Phys, 1985, pages 104–113. J.I.N.R., Dubna, USSR, 1987.

vanHulzen:88

[vH88] J. A. van Hulzen. Formule manipulatie m.b.v. REDUCE (in Dutch). Technical report, Department of Computer Science, Twente University of Technology, The Netherlands, October 1988.

vanHulzen:89a

[vH89] J. A. van Hulzen. Computer algebra and numerical mathematics: The odd couple? Technical Report Informatica 89-40, Department of Computer Science, Twente University of Technology, The Netherlands, June 1989.

vanHulzen:83

[vHC83] J. A. van Hulzen and J. Calmet. Computer algebra systems. In B. Buchberger, G. E. Collins, R. Loos, and R. Albrecht, editors, Computer Algebra and Symbolic and Algebraic Computation. Springer-Verlag, 2nd edition, 1983.

vanHulzen:82

[vHH82] J. A. van Hulzen and B. J. A. Hulshof. An expression analysis package for REDUCE. SIGSAM Bulletin, 16(4):32–44, November 1982.

vanHulzen:89

[vHHGV89] J. A. van Hulzen, B. J. A. Hulshof, B. L. Gates, and M. C. Van Heerwaarden. A code optimization package for REDUCE. In Proc. of ISSAC ’89, pages 163–170. ACM Press, New York, 1989.

vanHeerwaarden

[vHvH88] M. C. van Heerwaarden and J. A. van Hulzen. Pretty print facilities for REDUCE. Memorandum INF-88-36, Department of Computer Science, University of Twente, The Netherlands, August 1988.

Voros:77

[Vor77]  A. Voros.  Asymptotic k-

expansions of stationary quantum states. Ann. Inst. H. Poincare, 26A, 1977.

DeVos:89

[Vos89] Alexis De Vos. The use of Reduce in solar energy conversion theory. Reports of the CAGe Project 4, State University of Gent, CAGe Computer Algebra Group, August 1989.

Vinitsky:87

[VR87] S. I. Vinitsky and V. A. Rostovtsev. A use of REDUCE system in problems of hydrogen atom in an electric field. Preprint P11-87-303, J.I.N.R., Dubna, 1987.

Wanas:85

[Wan85] M. I. Wanas. Manipulation of parameters indicating the physical significance of any absolute parallelism space using RE-DUCE 2. In Tenth International Congress for Statistics, Computer Science, Social and Demographic Research, 1985.

Wanas

[Wanxx] M. I. Wanas. The third face of computer–computer solution of symbolic problems. Technical Report CAP-3 837, Military Technical College, Cairo, Egypt, 19xx.

Watanabe:76

[Wat76] Shunro Watanabe. Formula manipulations solving linear ordinary differential equations II. Publications of the Research Institute for Mathematical Sciences, Kyoto University, 11(2): 297–337, 1976.

Watanabe:79

[Wat79] Shunro Watanabe. A verification for non-existence of movable branch points of six painlevé transcendents by formula manipulations. Tokyo Journal of Mathematics, 2(2):285–291,

Watanabe:85

[Wat85] Yoichi Watanabe. Symbolic manipulation of structure functions in availability analysis. Technical Report UWFDM-658, Fusion Technology Institute, University of Wisconsin, Madison, Wisconsin, November 1985.

Wang:84

[WCvH84] Paul S. Wang, T. Y. P. Chang, and J. A. van Hulzen. Code generation and optimization for finite element analysis. In Proc. EUROSAM 1984, Lecture Notes in Computer Science, volume 174, pages 237–247. Springer-Verlag, 1984.

Witham:77

[WD77] C. R. Witham and S. Dubowsky. An improved symbolic manipulation technique for the simulation of nonlinear dynamic systems with mixed time-varying and constant terms. Journal of Dynamic Systems, Measurement, and Control, pages 157– 165, September 1977.

Wright:84

[WD84] F. J. Wright and G. Dangelmayr. Explicit iterative algorithms to reduce a univariate catastrophe to normal form. Technical report, Universität Tübingen, 1984.

Wulkow:90

[WD90] Michael Wulkow and Peter Deuflhard. Towards an efficient computational treatment of heterogeneous polymer reactions.

Preprint SC 90-1, Konrad-Zuse-Zentrum für Informationstechnik Berlin, January 1990.

Winkelmann:89

[WH89] Volker Winkelmann and Friedrich W. Hehl. REDUCE for beginners. six lectures on the application of computer algebra. In D. Stauffer, F. W. Hehl, V. Winkelmann, and J. G. Zabolitzky, editors, Computer Simulation and Computer Algebra. Lectures for Beginners, chapter 3. Springer-Verlag, 2nd edition, 1989.

Winkler:88

[WKL88] F. Winkler, B. Kutzler, and F. Lichtenberger. Computeralgebrasysteme (in German). Report 88-10, RISC - LINZ, Austria, 1988.

Wood:89

[Woo89] John C. Wood. Harmonic two spheres in the unitary group. Proc. London Math. Soc., 3(58): 608–624, 1989.

Weber:79

[WR79] Lawrence A. Weber and Gerhard Rayna. Problem #11 solved in REDUCE: A case study in program translation. SIGSAM Bulletin, 13(4):21–24, November 1979.

Wassam:87

[WTV87] W. A. Wassam, Jr. and Go.

Torres-Vega. Dual Lanczos transformation theory: Closed set of algebraic equations connecting Lanczos parameters with moments in moment expansions of time-dependent quantities. Chemical Phys. Lett., 134 (4):355–360, 1987.

Wassam:87a

[WTVNF87] W. A. Wassam, Jr., Go. Torres-Vega, and J. Neito-Frausto. Dual Lanczos transformation theory: Exact continued fraction expression for resonant γray absorption spectrum of a harmonically bound atom executing classical motion described by Smoluchowski dynamics. Chemical Phys. Lett., 136(1):26–30, 1987.

Wehner:1986:PHS

[WW86] M. F. Wehner and W. G. Wolfer. The pressure of a hard sphere fluid on a curved surface. Journal of Statistical Physics, 42 (3–4):509–521, February 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/
10.1007/BF01127724.

Yamamoto:87

[YA87] T. Yamamoto and Y. Aoki. RE-DUCE 3.2 on iAPX 86/286based personal computers. In Proc. EUROCAL ’87, Lecture Notes in Computer Science, volume 378, pages 134–135. Springer-Verlag, 1987.

Yannouleas:88

[YP88] C. Yannouleas and J. M.

Pacheco. An algebraic program for the states associated with the U(5) ⊃ O(5) ⊃ O(3) chain of groups. Comp. Phys. Comm., 52(1):85–92, December 1988.

Yannouleas:89

[YP89] C. Yannouleas and J. M.

Pacheco. Algebraic manipulation of the states associated with the U(5) ⊃ O(5) ⊃ O(3) chain of groups: Orthonormalization and matrix elements. Comp. Phys. Comm., 54(2 and 3):315– 328, June and July 1989.

Zhidkova:78

[ZNR78] I. E. Zhidkova, I. P. Nedyalkov, and V. A. Rostovtsev. On applicability limits of the experimental method for investigating strong gravitational fields. Technical Report P2 - 11589, J.I.N.R., Dubna, 1978.

Zahalak:87

[ZRS87] G. I. Zahalak, P. R. Rao, and S. P. Sutera. Large deformations of a cylindrical liquid-filled membrane by a viscous shear flow. J. Fluid Mech., 179:283– 305, 1987.

Zacrep:75

[ZY75] Douglas Zacrep and Bing-Lin Young. Trace and Ward-Takahashi identity anomalies in an SU(3) current model with energy-momentum tensor. Phys. Rev. D, 12:513–522, 1975.

Zeng:84

[zZlH84] Wan zhen Zeng and Bail lin Hao. Scaling property of periodn-tupling sequences in onedimensional mappings. Commun. in Theor. Phys., Beijing, China, 3(3):283–295, 1984.